A Bibliography of Publications about Common Lisp

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
             WWW URL: http://www.math.utah.edu/~beebe/

19 August 2025 Version 2.52

Title word cross-reference

∀∃ [CM98].

0 [Ano89a]. 0-12-774861-X [Ano89a]. 002R [BDG^{+}88c, BDG^{+}88d].

1.3 [Ano90e, App90]. 102 [AC94]. 10th [Fre90]. 11th [IEE89c, IEE95b, JU91]. 12th [LS97]. 13th [IEE86b]. 15-16 [Gil86]. 16th [IEE89f]. 18th [Ano87a]. 1986 [ACM86, HJS86, SA86]. 1987 [TW87]. 1990s [IEE88f]. 1991 [MCH91]. 1997 [AHWN97]. 199x [AC94]. 1st [IEE85a, IEE89b, SA86].

2 [Lan86]. 2.0 [Ano91b, Ano92, Ano93, Vas97]. 20-22 [Kna89]. 2002 [Bän03]. 20th [ACM87, IEE94b]. 25th [IEE88b]. 2A [Sym88]. 2b [Sym86]. 2nd [IEE89b].

3.0 [Sun88]. 3D [WNA90].

4.0 [Pur90a].

5th [MHG90].

68th [Ano87a].

7th [IEE86e, Nie93].

’85 [Ano85a, Cav85]. ’86 [ACM86, Ano86a]. ’87 [IEE87e]. 88 [BDG^{+}88c, BDG^{+}88d]. 88-002R [BDG^{+}88c, BDG^{+}88d]. ’89 [IEE89d, IEE89g].

’90 [IEE90e, N^{+}90]. ’91 [IEE91d, Wat91c]. ’93 [ACM93, IEE93b, Nie93]. 93-102 [AC94]. ’94 [IEE94a, IEE94b]. ’95 [ACM95, K^{+}95]. ’97 [ACM97, Küc97].

A.I. [SCP^{+}85]. Århus [Ano94a]. Abstract [GP97, PE88, CK87a]. Abstracts [Ano87a].

Academic [Ano89a]. ACL2 [KM96, You97]. ACM [ACM93, IEE88b]. ACM/IEEE [IEE88b]. Acoustics [IEE97a]. Acquisition [PRD88]. across [VS90]. Action [Lan89a, Rat12]. active [Cha95]. Ada [IEE86c, RW86, War86]. adaptive [HK91, HBG96]. add [LZ94]. Adding [Car84, LZ97]. Addison [Ano89b, Tou90a]. Addison-Wesley [Ano89b, Tou90a]. addressing [SIG89]. Advanced [AM99, Gra94, Hal90, Tei88]. Advances [DH98]. Advancing [IEE86a]. Advisor [Mu 95, MuT95, Mun95b]. Advisory [Sig85]. Aerospace [IEE85b, IEE91c, IEE97b, NJB89]. agent [DL96]. agent-based [DL96]. agents [And97, Gat97, Mai91, Pog96]. AI [Ano87c, Bir85, Cag86, Fal88, HG91, Kai86, Pog96, PM93, SGBL86, Vau85]. AI-based [HG91]. AIAA [IEE86e]. Aid [Nar88a, PM93, SK90a]. aided [Jel89, Sal90, SK90b]. aiding [IEE91a]. AIE [ACM90a]. aircrew [SB89]. Alberta [ACM88]. Albuquerque [ACM93]. Algebra [Cav85, Pav85, Cla92, Coo95a, Coo95b, Fat91, McD89b]. Algebraic [ACM94a, Küc97, Lev95, OK89, Wat91c]. algebras [CT89]. Algorithm [Bur86, GK90]. algorithms [Car90, KH88]. Allegro [Ano90e, App90, Fra88a, Fra94, Fra96, KLF^{+}91, Chl91, CF88, Kem91, Lan88b, Lan88a]. allocation [LC95]. alternative [CR89]. American [Ano87a, IEE87f, IEE89e, MCH91, Ame92]. amulet [Mye95]. Anaheim [IEE88b, RCG^{+}89]. analog [SGW89]. Analysis [ABC^{+}88, LK87, Nar88a, Nar88b, Ehn89b, Hur91, Kv89, Lan92, MTI95, ND95, SK96, Shi88, Sly93]. Analysis-Based [Nar88b]. analyzer [Bas93]. Angeles [E^{+}93, IEE95b]. animate [Fut90]. animation [MSBH93]. Annual [Ano85a, Ano86a, Ano87a, HSKS88, HMS^{+}90, IEE88h, IEE89f, IEE93a, Kna89, PO90, ACM87, HMS^{+}89, IEE84a, IEE86b, Kin89, Tei88]. ANSI [Gra96, Gra97, Gra99, Mar04]. Antonio [IEE86a]. Apollo [Gre84]. Appendix [Ano90a]. Apple [Ano92]. Application [GM86, KPB^{+}88, K^{+}90, SA86, Ste90b]. Applications [ACM86, ACM90a, Gil86, IEE85c, IEE86c, IEE88e, IEE93a, IEE93c, IEE95b, IEE97b, KUJP97, Kin89, Kna89, SCP^{+}85, TBE^{+}94, AP88, BR93, BP97, Fat91, FGGB89, Hen92b, LLC89, NJB89, PK91, PM93, RBLS98, Spe92a, SMK^{+}88, VS90]. Applied [Ren86]. Approach [CH86, CH87, FBHJ87, Her93, OOT87, BHW91, BHA89, CR89, DL93, HK90, LC95, MC90, RW90, Sha86, Sha92, Ste89]. Approaches [Vau85]. April [Ano86b, Cav85, Gil86, IEE86c, IEE86d, IEE88g, IEE93d, IEE97a, Jan90, N^{+}90, SA86]. Architectural [Gre84]. Architecture [IEE86b, IEE89f, Lim87, MSD^{+}85, THL^{+}86, Yua92, BGHW88, Bla86, HH93, Kv89, Lam90, RMM94, VB95]. Architectures [Cha95, GL94]. Arithmetic [Yan91]. Arizona [IEE93a, NKC91]. Array [K^{+}90]. art [KdR91]. Artificial [AAA93, ACM90a, ACM95, Ano94b, Fol86, FBHJ87, Had87, IEE85c, IEE88e, IEE91d, IEE93b, IEE93c, IEE95b, IEE96, Nae84, Noy92, PG92, SA86, Gil86, Har90b, HD89, LLC89, Nor91, Nor92, Nor04, Tan90, Tan95, Tei88, Wat91b]. ASCE [MCH91]. ASME [BPR^{+}94, GA^{+}86, RCG^{+}89, TPB^{+}88]. Aspect [IO85]. Assemblies [AIST88]. Assembly [CB86, DL93, MC90]. assess [SB89]. assistant [WT91]. Assisting [BB89]. Atlanta [AHWN97]. Atlantic [Fre90]. atmospheric [Jan90]. Attitude [KBS^{+}87]. Attribute [NP88]. Attributes [Flo99]. August [ACM84, Ano94b, HJS86, JU91, RCG^{+}89, TPB^{+}88]. Austin [ACM84]. Australia [IEE95a, MDM97]. Austria [Cav85, Gil86]. AutoCAD [Sal90, Sly93].

AUTOLISP [AP88]. Automated [BY96, CM98, Nis94]. Automatic [Bal85, Cor86, CB86, IEE86a, JU91, SYM92, Van98, Wat91a, Lin91]. Automation [IEE86d, IEE88b, IEE88g]. Autonomous [Gli86, BCG90, Gat97]. Autotestcon [IEE86a]. Avalon [Cla89, CLNW90, HW87]. Avalon/Common [Cla89, CLNW90]. Avionics [IEE86e]. Aviv [IEE88i]. avoidance [MC90]. AWS [Ano87a]. Axis [Dol90].

B [IEE86a]. balancing [CH88]. Baltimore [IEE88c]. Banff [ACM88]. Base [CK87b, BWM89, HK91]. Based [BBD88, Cor86, CB86, DD90, Nar88b, OSD86, Pin85, Que76, Sig85, BB90, BHA89, Bur91, CH93, DZB97, DL96, DL93, DKL95, DL98, GKS89, HR89, HK90, HG91, IEE85c, KM97, Kv89, KM90, Lam90, Lar96, LC95, LS92a, LS92b, Lin91, Lub91, Luc92b, MC90, MTI95, PAR93, PM93, Rat93, RBLS98, Sal90, SH96, Sly93, SB89, SST89, Tib93, UJ92, WT91, WGS96, Kar89, WNA90]. Bayer [Lan89b]. BC [IEE89b]. Beach [IEE85c, IEE86c, IEE88a, Tei88]. bed [SMK^{+}88]. behavior [ND95, Sha91]. belief [Har90a]. Benjamin [Whi91]. Benjamin/Cummings [Whi91]. Berlin [Bän03]. better [Hol92]. beyond [PO90]. Biltmore [E^{+}93, NKC91]. binding [Coo95a, Coo95b]. bioinformatics [SK96]. Biology [PO90]. Biomedical [PO90, BWM89]. Biosphere [Ste91a]. BIST [SK90b]. blackboard [HH93, KK93, Lam90]. Block [KH88]. board [LMM89a, LMM89b]. boiling [Lin91]. Bologna [IEE94b]. Bombay [IEE89g]. Bond [KL95]. Bonn [Wat91c]. Book [Ano89a, Ano89b, Ano90b, Tou90a, Whi91, Mar04]. Boston [IEE93b]. Boulder [MCH91]. Brazing [Ano87a, NTS90]. BRICAL [HI94]. bridge [HI94]. Broaching [GM86]. Brook [IEE87e]. browser [CF88]. Build [Hol14]. Builder [HKLH93, CFTZ89]. Building [BR93, Her93, JR90, FGGB89]. Bureau [MCH91].

C [Ecc88, Hol14, Nym87, USE88, Vos94]. CAD [IO85]. CADD [Boo84]. California [ACM92, E^{+}93, IEE85a, IEE86d, IEE87b, IEE87c, IEE88e, IEE89d, IEE91d, IEE93d, IEE95b, RCG^{+}89, Tei88, TPB^{+}88, IEE84b]. Call [Hal92]. calls [Sto94]. Cambridge [IEE89a, Ste90b]. Canada [ACM88, IEE89b, Lev95]. Canonical [Fat91]. Capri [IEE92]. Care [Kin89, PO90]. Carolina [ACM90a]. case [DZB97, Nor91, Nor92, Nor04]. case-based [DZB97]. Castle [IEE88a]. Casualty [Sig85]. Cathedral [IEE87b, IEE87c]. CCRC [Lou89]. cell [MC90]. Center [Ano85a, Ano86a, IEE85b, IEE86a, IEE88b, IEE90e, IEE91c, Tei88]. Centre [IEE90d]. Chains [BWZ94]. Challenges [Bän03, Bän03]. changing [IEE90c]. chaos [Nis94, Wat91b]. Character [YY94]. Charleston [ACM90a]. Charlottesville [IEE91a]. Chart [Mu 95, MuT95, Mun95b]. Chatter [EWB87]. checked [Moo99]. checker [HG91]. checking [Ake94, SIG89]. CHI [K^{+}95]. Chicago [Ano86b, Ano87a, GA^{+}86]. Chinese [Gui88]. chips [BGHW88]. Churchill [Ste90b]. CIMVIEW’ [Ano90a]. CIMVIEW’-prototype [Ano90a]. Circuit [MP87, LMM89a, LMM89b, LM89]. circuit-board [LMM89b]. Circuits [IEE87a, IEE88d, IEE90a, IEE91b, Wuo89, Got91, KL89, SGW89, SK90b]. CISC [Bän03]. City [Fre90, IEE87a, LS97]. Civil [MCH91, PK91]. CL [Fra94, Fra96, KLSW94, Kem91, KLF^{+}91, LW95]. CL/TB [Kem91, KLF^{+}91]. Class [BK91]. classes [Lou89]. Classification [Pup87, BHA89]. CLIB [HKLH93]. CLIM [McK91, Vas97]. CLING [CR90]. CLING/SQL [CR90].

Clisp [Tei76]. CLM [Sch94]. Clojure [ECG12, FH11, Hal09, Rat12, VS10]. CLOS [Ano87b, GWB91, HP96, KG89, LM91, LZ94, LZ97, Pae93, PAR93, SO91, Swa90a, Tau90, Tau91, Wat90, Ano89b]. closed [BWZ94]. closed-form [BWZ94]. CLtL2 [NSI^{+}92]. CLUE [Kim90]. CMOS [LM89]. CMU [Mac91, Mac92a, Mac92b, McD87a, MFW87, MFW88, McD89a]. CO [ACM87, USE88]. Code [Bro84, Ano90a, BY96, KVJ90, Lan86, Nym87, WW92]. Codecs [MP87]. CODER [NJB89]. Cognitive [Pir86, RGG90]. Collaborative [DL96]. Collection [Bro84, FN81, Yua92, Ehn89a, Ehn89b, Zor90]. College [ACM95, Ste90b]. collision [MC90]. Colorado [ACM87, K^{+}95, MCH91]. colored [YYS95]. Combination [OK89]. Combining [WL95]. Commercial [JR90]. Commercial-Quality [JR90]. COMMON [Chl91, Lan88b, Pur90a, Pur90b, Whi91, WS85, Ame92, Ano89b, Ano90b, Ano90e, Ano91a, Ano91b, Ano92, Ano93, App90, Bro87, CK87a, CF88, Flu91, Giu87, Gra96, Gra97, Gra99, HP96, HSW85, HSW87, IAC94, Jac94, Kem91, Kim90, Lan88a, Luc92b, Mac91, Mac92a, Mac92b, Mar04, McD87a, MFW87, MFW88, McD89a, NSI^{+}92, Neu03, Rei96, SWH85, Ste86, Ste89, Sun88, Sym86, Sym88, Tou90a, WS84, YY94, Yua90, KLF^{+}91, Wer89, AP88, AM86, Bas93, Bla86, CT89, CT90, Fra88a, Fra88b, Hen89, LM91, Tat87, YH87, AC94, AI96, Abd90, Acu89, Acu90, Ake94, AH91, AM99, Ano87b, Ano90a, ABFM90, Bak92, BKK^{+}86, Bob87, BDG^{+}88a, BDG^{+}88b, BDG^{+}88c, BDG^{+}88d, BR93, BCG90, Bra86, BK91, BGPS91, BG84, Bro85, Bro87, CL88, CR90, Chi89]. Common [Chr93, Cla89, CLNW90, Cla92, CHD09, DB97, Dem88, DD90, Dol90, Ecc88, Esa90, FL91, FBWR95, Flu91, Gad88, Giu87, GK90, Gra94, Har90b, HD89, Hen92b, Her93, HKLH93, Hum94, Ida90, JR90, JS94, JS97, KUJP97, KT88, Kan91, KM97, KG89, KA95, Kv89, Koo87, Kos90, Lam05, Lei08, Lou89, Luc92a, Ma89, MK90, MN91, McD87b, MK86, McK91, MBPLRR10, Mil88, Mu 95, MuT95, Mun95b, Mur90, NC88, Nar88b, Neu02, Nor91, Nor92, Nor04, Noy92, Nym87, OK89, PKSA89, PAR93, Ric90, RMM93, RMM94, Ril91, Ril92, Sch94, Sei05, Sev91, Sha92, Sim87, Sla98, Sny86a, Sny86b, Spe92a, Spe92b, Ste84, Ste90a, Ste91b, Tan90, Tan95, Tau90, Tau91, Thr88, Tou90b, Van98, Ver19, VS90, Vog98, Wat89, Wat91b, Wat90, WF84, WS87, Wil86]. Common [Yan91, YYK^{+}90, You92, Yua88, Yua92, YYN90, Ano89a]. Common-LISP [CL88]. CommonLoops [KHDS87]. CommonObjects [Sny86a, Sny86b]. Communication [IEE90d, CR89]. Communications [IEE84a, IEE89b, IEE89g, Jan90]. companies [KL95]. companion [Kos90]. Company [Tou90a, Whi91]. Comparing [Vau85, You97, Zor90]. Comparison [SO91]. Compcon [IEE84b, IEE87b, IEE87c]. compilation [KH88]. Compiler [KT88, Sam77, YYK^{+}90, Chi89, Mac92b]. Complex [Flo99, GL94, IEE91a, LQ96, Sch92, Tib93]. Components [CH86, CH87]. Composers [SH96]. Composite [MHMR96]. Composition [Tau90, Tau91]. Compsac [Kna89, IEE93a]. Computation [IEE95a, Küc97, Lev95, ME95, Wat91c, ACM97, GHL^{+}90, KK93, Tou90b, Wan97, ACM94a, Whi91]. Computational [KL89, Sha91]. computations [EL89]. Computer [Cav85, IEE84b, IEE85a, IEE85c, IEE86b, IEE86c, IEE87b, IEE87c, IEE87e, IEE88a, IEE88i, IEE89f, IEE90b, IEE90d, IEE93a, IEE94a, IEE94c, Kin89, Kna89, Kru85, Pav85, Pir86, Que76, Ren86, SK90b, Ano94a, Bur91, Cla92, Fat91, GHL^{+}90, Jel89, LC95, Pur90a, Pur90b, Sal90, Tei88].

Computer-aided [SK90b, Jel89, Sal90]. Computers [GA^{+}86, IEE84a, IEE87e, IEE89b, RCG^{+}89, TPB^{+}88, IEE89g, LS97]. Computing [Ano90c, Bän03, BKR^{+}91, IEE87d, IEE88f, Neu03, Car90, Cla89, CLNW90, K^{+}95, Lou89, Neu02, PG92, Bän03]. Conception [Flu91]. Concepts [BDG^{+}88a, IEE86a, Sym88]. Conceptual [IO85, KL89, SB89]. Concurrent [DLK92, SST89]. Conference [AAA93, ACM84, ACM86, ACM90b, ACM90a, ACM92, ACM93, ACM94b, ACM95, AHWN97, Ano87a, Ano94b, Bän03, Cav85, E^{+}93, Fre90, GA^{+}86, HSKS88, HMS^{+}89, HMS^{+}90, HJS86, IEE84b, IEE85a, IEE85b, IEE85c, IEE86a, IEE86b, IEE86c, IEE86e, IEE86d, IEE87b, IEE87f, IEE87e, IEE88b, IEE88a, IEE88e, IEE88h, IEE88g, IEE88i, IEE89a, IEE89b, IEE89e, IEE89c, IEE89g, IEE90c, IEE90b, IEE91a, IEE91c, IEE91d, IEE92, IEE93a, IEE93b, IEE93c, IEE93d, IEE94a, IEE94b, IEE94c, IEE95a, IEE95b, IEE96, IEE97b, IEE97a, Kna89, K^{+}90, MDM97, NKC91, N^{+}90, PO90, RCG^{+}89, SA86, Ste90b, TPB^{+}88, USE88, Wuo89, Ano86b, Ano94a, IEE84a, IEE87c, IEE88c, IEE90d, K^{+}95, LS97, Nie93, Tei88]. conference/exhibition [Tei88]. Configuration [MSD^{+}85, ZS93]. configurations [GKS89]. Congress [JU91, MCH91]. Connection [Ste86, WS87]. connectionist [AD88]. consistency [You97]. constraint [SM93]. constraints [Ric94]. Construction [Her93, MSBH93]. Construction-Kit [Her93]. content [CR89]. Context [CHD09]. Context-oriented [CHD09]. Contribution [Ehn89a]. Contributions [Cav85]. Control [IEE87f, IEE89e, IEE94b, KBS^{+}87, MHG90, Mu 95, MuT95, Shi88, DL96, JU91, Lam90, Lan89b, Lin91, Mun95b, NRK89, NRK91, Ruo88, Sit93]. Convention [Ano87a, IEE85b, IEE86a, IEE88b, IEE90d, IEE91c, Tei88]. conventional [Fal88]. Conversational [Tei76]. Converting [Lan86]. cooperation [MCH91]. COoperative [ZB91]. coordination [KM90]. copy [Zor90]. cost [WT91]. counseling [Har90a]. Coupled [KBS^{+}87]. Coupling [BCG^{+}88]. CRAY [WW92, AGK^{+}87]. Creativity [K^{+}95]. Critique [BG84]. critiquing [MHMR96]. cross [Sim87]. cross-referenced [Sim87]. Cummings [Whi91]. cure [Hol92]. CVPR [IEE89d, IEE94a]. Cybernetics [IEE89a, IEE91a].

DAC [IEE88b]. Dagwood [FIKY88]. Data [Bro84, Car90, BPR^{+}94, HK91, WS87]. Data-parallel [Car90]. Database [BPR^{+}94, BCG^{+}88, NSS94, XRS96]. databases [Cha95]. dataflow [Chi89]. DataWorks [Ano90d]. David [Whi91]. Dayton [IEE85b, IEE91c]. DBMS [FBB93]. DC [AAA93, IEE90c, Kin89]. DeBuMA [FGGB89]. December [ACM87, AHWN97, E^{+}93, IEE85c, IEE95a, NKC91]. Decision [Bak92, NRK89, NRK91, BB89, GW98, IEE91a]. Defining [JS97]. definition [BB89]. Delivering [Hen92b]. demo [Spe92a]. Denmark [Ano94a]. Denver [K^{+}95, USE88]. dependencies [Lub91]. Description [Sam77, FGGB89, GKLM90, Mal91]. Design [ACM93, CH86, CH87, GM86, Hol92, IEE87e, IEE88b, IO85, KSv88, SPM91, TSL^{+}88, WF84, WGS96, Yua90, Chr93, DZB97, DLK92, GH94, HR89, HG91, Jel89, KL89, Lub91, MFW87, MFW88, MHMR96, NJB89, RW90, Sal90, SB89, SK90b, UJ92, ZS93]. designer [Jel89]. Designing [GL94, TS90, LQ96]. detection [BHA89]. Develop [Hal90, Pog96]. Developing [Jel89, BT94]. Development [JLHB92, KBS^{+}87, KM94, Nar88a, NTS90, RBLS98, SGBL86, BB90, Bur91, CN91, Mye95, VS90]. Device [MP87]. DEVS [PAR93].

DEVS-based [PAR93]. Diagnosis [Lar96]. Diagnostic [Lan89b, Pin85]. Diagnostics [War86]. diagram [Fut90, KH88]. diagrams [SYM92]. dialect [WS87]. dictionary [Sym86]. Diego [Ano89a, IEE88e, IEE89d]. differential [ND90, ND95]. Digest [IEE84b, IEE87b, IEE87c, Wuo89, IEE87d]. digestion [Lan89b]. Digital [IEE86e, Jan90]. DIMACS [DH98]. dimensional [ND90]. Directed [RW86]. discrete [Sev91]. Display [Bar84, Boo84, FL91]. Display-Oriented [Bar84]. Distributed [Ama91, HW87, IEE88f, Lam90, ZB91, BP97, CH88, Cla89, CLNW90, GL94, LC95, Pog96, Wan97]. distributed-memory [CH88]. Division [MCH91]. Document [BDG^{+}88c, BDG^{+}88d]. documentation [BK91, Ver19]. DOE [HC85]. DOE-MACSYMA [HC85]. Domain [Gre84]. Done [Ida90]. Draft [Ame92, AC94, IAC94, Ano91a, Ano91b]. drill [Yua88, Ano89a]. Driving [IEE84b]. DSP [KH88, KVJ90, RK97]. Dunfey [IEE87a]. During [Ida90]. Dynamic [Bur86, Hum94, Mal91, CH88, Dav88, ME95]. dynamical [NRK89, NRK91]. Dypas [Dav88].

earth [Jan90, ME95]. earths [XXL92]. EAVE [BCG90]. EC [IEE89g]. ECN [Pur90a, Pur90b]. ECOOP [Nie93]. editing [KH88]. edition [McD87a, McD89a]. Editor [Bar84]. educational [HI94]. efficiency [Ehn89a]. Efficient [Bur86, Wat87, Flu91, Hen92b, HH93, McD87b, NSI^{+}92]. effizienten [Flu91]. Eiffel [SO91]. Eighth [IEE96, IEE84b]. Electrodes [FBHJ87]. Electronics [CB86, IEE85b, IEE91c, IEE94b]. electrons [IEE89g]. elements [HR89, Tan90, Tan95]. Eleventh [AAA93]. Elis [KT88]. Emacs [MS20]. Embedded [KA95, BP97, Gat97]. Emulator [CK87a]. end [JLHB92, Lar96].

Endpaper [Acu90]. Energy [PG92, IEE89g]. Engineering [ACM90a, BPR^{+}94, IEE88i, IEE89c, IEE92, MCH91, PG92, SA86, GA^{+}86, HI94, IEE85c, NSS94, PK91, PO90, Pur90a, Pur90b, RCG^{+}89, TPB^{+}88]. Engineers [MCH91]. engines [Ruo88]. Enhancing [Kla04]. enter [Fal88]. enterprise [BPR^{+}94]. Environment [Gli86, Had87, PAR93, Ric90, CN91, Cze90, DL96, Gad88, Gui88, Kar89, KVJ90, KM94, Luc92b, Mun95a, RBLS98, RGG90, SM88, SH96, SMK^{+}88, WGS96]. Environments [BSS84, IEE86c, Mye95]. Equations [Bur86, ND90, ND95]. Equipment [Cor86]. equivalence [Fut90]. era [Wat91b]. Ergo [NP88]. ESL [Gat97]. Espoo [IEE88d]. Estimation [OSD86, YYS95]. Estonia [JU91]. EUROCAL [Cav85]. Europa [IEE92]. Europal [Ste90b]. European [Cav85, Nie93, Ste90b]. EusLisp [Mat90]. Eval [MSD^{+}85]. Evaluation [MSD^{+}85, THL^{+}86, BWZ94, Cha95]. event [Sev91]. EVES [PKSA89]. Evolution [MS20, Nar88b, SG93]. Evolutionary [BB90, IEE95a]. exams [Van98]. executable [Nym87]. Execution [MK86, GP97, Gat97]. Exhibition [GA^{+}86, IEE90d, TPB^{+}88, Tei88]. existential [Mai91]. expedite [BWZ94]. Experience [KUJP97, KHDS87, XRS96]. Experiences [Wat90]. Experimental [Wan86]. experiments [ACW^{+}94]. Expert [ACM90a, Ano86a, Boo84, GM86, GH94, Had86, KBS^{+}87, LK87, LMM89a, LMM89b, Nae84, PK91, PG92, Pup87, SPM91, War86, ZB91, Ano85a, BCG^{+}88, CFTZ89, Cla92, Hal90, HR89, Jel89, KL95, Lan89b, Lin91, NTS90, UJ92, XXL92, YJM91]. expert-seisd [UJ92]. expert-system [CFTZ89]. Explainable [SPM91]. Explanations [SPM91]. explicit [Lar96]. exploration [Nis94]. exploratory [HBG96]. explores [ND90]. Exposition [RCG^{+}89].

Expressions [Wat91a, Wat87]. Extended [Boo84, Yua92, CL88, NC88]. Extending [Sev91]. Extension [KA95, Chl91, Sny86a, Swa90a, Thr88]. extensions [ZHL^{+}89a, ZHL^{+}89b]. Extracting [LM89]. extraction [XXL92].

f2cl [BW96]. Facility [Sam77]. factors [K^{+}95]. factory [Sly93]. Fans [Flo99]. Fast [FBWR95, MSD^{+}85, OOT87, Rei96]. Fault [IEE87d, PRD88, YYS95, Hol92]. Fault-Tolerant [IEE87d]. features [Moo12]. February [IEE84b, IEE87b, IEE87c, IEE95b]. Federation [JU91]. Fifth [IEE93b, ACM95]. final [PKSA89]. Finland [IEE88d]. firing [SV97]. Firmware [OOT87]. First [Ano85a, HSKS88, IEE88h, N^{+}90, Ste90b]. Flavor [GM86]. Flavors [Ecc88, MN91, Moo86]. Flexible [DL93, GM86, Tib93, BP97, MC90]. Floating [FBWR95, Rei96, Vog98]. Floating-Point [Rei96]. Florida [ACM94b, IEE85c, IEE86c, IEE88a, IEE93c, Jan90, Kna89]. flow [Shi88, Sly93]. fly [Ehn89a, Ehn89b]. Fontainebleau [IEE85c]. Form [Flo99, BWZ94]. formal [MSBH93]. format [BHW91, WNA90]. formations [BHA89]. Formula [Que76]. Fort [IEE86e]. Forth [CK87b]. FORTRAN [WW92, BW96]. foundations [BB89]. Fourth [IEE88e, IEE92, IEE89g, Tei88]. frame [Rat93]. frame-based [Rat93]. Framework [Nar88b, LQ96, NSS94, ZB91]. France [ACM90b, IEE89c, IEE94c, IEE96, PG92]. France-Telecom [PG92]. Francisco [ACM92, IEE84b, IEE86d, IEE87b, IEE87c, IEE93d, TPB^{+}88]. Franklin [IEE88g]. Franz [KK22, Tou90a]. friendly [RGG90]. front [JLHB92]. FTCS [IEE87d]. fuel [GKS89]. Fully [TSL^{+}88]. Function [Yan91]. Functional [ACM84, ACM90b, ACM92, ACM94b, Ake94, GWB91, Web95].

Functions

[BDG^{+}88b, Car84, FL91, BWZ94]. fundamentals [Noy92]. Future [IEE88f, MCd^{+}94, Mur90, Ano90f]. futures [NC88]. fuzziness [ZS93]. Fuzzy [IEE93d].

Garbage [Bro84, FN81, Yua92, Ehn89a, Ehn89b, Zor90]. Garden [IEE88f]. Garnet [Mye95]. gate [Got91]. gene [Ste91b]. General [Giu87, SH96, WNA90]. Generalized [Fut90]. Generating [Pin85]. Generation [Bur86, Got91, KVJ90, Lin91]. generator [Bas93, GKLM90, WW92]. genetic [GK90, Har90a, WL95]. genome [XRS96]. Gentle [Whi91, Tou90b]. Gentler [Whi91]. Georgia [AHWN97]. Geotechnical [MCH91, MCH91]. German [Bro87, Flu91]. Germany [IEE97a, Nie93, Wat91c]. Getting [Ano91a, Swa90a]. GINA [Spe92a, Spe92b]. Given [FIKY88]. GKS [Cze90]. Global [Hen92a]. GLORP [Lei08]. Golden [HSW85, HSW87, SWH85, Ste89, WS84, WS85]. Gonzalez [IEE86a]. Graham [AM99]. grammar [Web95]. grammars [WL95]. graph [CF88, KLSW94, RW90]. graph-parsing [RW90]. Graphic [SMK^{+}88]. Graphical [ABFM90, Hur91, KH88]. Graphics [TSL^{+}88, Bir85]. Grasper [KLSW94]. Grasper-CL [KLSW94]. Grid [TW87]. Group [Ren86, TBE^{+}94]. Guide [Ano89b, Luc92b, Fra88a, Fra94, Fra96, Kan91, KG89, Luc92a, Tat87, Tat10].

Hall [Ano90b]. Handling [Sit93, YY94]. hands [Ste89]. hands-on [Ste89]. Harbor [IEE88e]. Hardware [Bro84, MSD^{+}85]. Hawaii [HSKS88, HMS^{+}89, HMS^{+}90, IEE88h, Küc97]. Headquarters [N^{+}90]. health [PO90]. HECODES [ZB91]. held [IEE85b, IEE91c, Tei88]. help [TAH89]. Henry [IEE86a]. heterogeneous [LC95, SK96, VS90, XRS96, ZB91].

Heuristic [CB86]. Hewlett [Cag86]. hexapod [KM90]. HF [CN91]. HI [ACM97, IEE88h]. High [Giu87, KB87, PG92, YYN90, CN91, Kar89]. High-Level [KB87, Kar89]. High-Performance [YYN90]. Higher [Mur90, PE88]. Higher-order [PE88]. HiLISP [YYK^{+}90, YYN90]. Hill [IEE87b, IEE87c]. Hilton [IEE85c, IEE86d, IEE87e, IEE88i, IEE89e, IEE90e]. History [KK22]. homogeneous [KVJ90]. Hong [IEE88f, IEE90d]. HORNE [AM86]. Hospital [Had87]. hosted [MCH91]. Hotel [AHWN97, E^{+}93, IEE84b, IEE86c, IEE86e, IEE87b, IEE87c, IEE87a, IEE87f, IEE88a, IEE88e, IEE88g, IEE88f, IEE88i, IEE89e, IEE90c, IEE90a, IEE91a, IEE92, Kna89, N^{+}90]. House [OSD86]. HP [HWK95]. HTTP [Van98]. Human [K^{+}95, Sch92]. Hyatt [IEE87f, IEE89a, Kna89]. Hybrid [CH86, CH87, CH88, HR89, SGBL86, BP97, CTB91, Lan89a]. hydraulic [KL89]. Hypadapter [HBG96]. Hypermedia [DB97]. hypertext [DL96, HBG96].

I/O [Car84]. IBM [McD87a, McD87b, MFW87, MFW88, McD89a]. ICAIL [ACM95]. ICCD [IEE87e]. ICMC [Ano94a]. ICSI [N^{+}90]. IEA [ACM90a]. IEA/AIE [ACM90a]. IECON [IEE94b]. IEE [IEE85c]. IEEE [IEE84b, MHG90, PO90, Wuo89]. IEEE/AIAA [IEE86e]. II [IEE88h, MSD^{+}85, PG92]. Illinois [Ano86b, GA^{+}86]. Image [Mun95a, BHA89, Car90, Jan90, KM94]. Immediate [FN81]. Implementation [ACM93, DD90, Giu87, GM86, KSv88, KT88, WAF90, CL88, Chr93, GK90, Kan91, Mat90, NSI^{+}92, Yua90]. Implementing [AGK^{+}87, Dol90, JS97, SV97]. Implications [War86]. improve [YJM91]. Improved [FN81]. increase [Ehn89a]. independent [Vas97]. index [Sim87]. India [IEE89g]. inductive [WL95]. Industrial [ACM90a, BBD88, IEE94b, KM97, KM96, KL95, Moo12, Sly93]. industrial-strength [Moo12]. Inference [Ma89, MK90, Hen92a]. Information [AI96, IEE89g, LM89, VB95]. infrared [BHA89]. INGRES [CR90]. INGRES/SQL [CR90]. inhibiting [Got91]. Innsbruck [Gil86]. inputs [Got91]. Insertion [CB86]. Insight [UJ92]. Instruction [AIST88, WF84, Kla04, MCd^{+}94]. Instrumentation [IEE94b]. Integrated [ABC^{+}88, CFTZ89, NSS94, TSL^{+}88, GHL^{+}90]. Integrating [LQ94, Pog96, Sha91, BPR^{+}94, GWB91, SIG89]. Integration [N^{+}90, Mul91]. intellectual [IEE87b, IEE84b]. Intelligence [AAA93, ACM90a, ACM95, Ano94b, Fol86, FBHJ87, Had87, IEE85c, IEE88e, IEE91d, IEE93b, IEE93c, IEE95b, IEE96, Nae84, PG92, SA86, Gil86, Har90b, HD89, LLC89, Nor91, Nor92, Nor04, Noy92, Tan90, Tan95, Tei88, Wat91b]. Intelligent [BS85, Cla92, IO85, MHG90, BB89, GL94, HI94, JLHB92, NSS94]. Interaction [LK87, HWK95]. Interactive [BSS84, CR89, CH86, CH87, PRD88, RK97, You97, Coo95a, Coo95b, Sha86, Sha92]. interchange [BHW91]. interests [IEE88c]. Interface [ACM88, BDG^{+}88a, BDG^{+}88b, Chr93, HKLH93, Hum94, RMM93, CR90, Cla92, DZB97, HK91, Kem91, KLF^{+}91, Luc92a, McK91, Mye95, RMM94]. Interfacing [ABFM90]. Interlisp [Bar84]. Internal [MFW87, MFW88]. International [ACM90a, ACM94a, ACM95, Ano87a, Ano94a, Fre90, GA^{+}86, HSKS88, HMS^{+}89, HMS^{+}90, HJS86, IEE84a, IEE84b, IEE85a, IEE86a, IEE86b, IEE86c, IEE86d, IEE87b, IEE87d, IEE87a, IEE87e, IEE88a, IEE88h, IEE88d, IEE88g, IEE89a, IEE89f, IEE89c, IEE89g, IEE90c, IEE90b, IEE90a, IEE91a, IEE91b, IEE91d, IEE92, IEE93a, IEE93b, IEE93d, IEE94b, IEE94c, IEE95a, IEE96, IEE97a, JU91, Kna89, Küc97, K^{+}90, Lev95, MHG90, N^{+}90, PO90, PG92, RCG^{+}89, SA86, TPB^{+}88, Wat91c, Wuo89, ACM97, IEE87c, LS97]. interoperation [XRS96]. interpretation [Wat87]. Interpreter [OOT87]. intertool [BHW91]. Interval [Flo99]. interview [MTI95]. Intranet [DB97]. Introduction [Whi91, YH87, Cag86, SH96, Tou90b, Wer89]. IRCAM [Hum94]. ISBN [Ano89a]. Island [IEE88e]. Isn’t [EWB87]. Isolation [PRD88]. Israel [IEE88i]. ISSAC [Küc97, Wat91c]. ISSAC’94. [ACM94a]. ISSAC’95 [Lev95]. issue [Lub91]. issue-based [Lub91]. Issues [Lim87]. Italian [AD88]. Italy [IEE92, IEE94b]. iterative [BB90]. IV [Gil86].

January [HMS^{+}89, PG92]. Japan [IEE86b, Ida90]. Japanese [YY94]. Jersey [Fre90, K^{+}90, N^{+}90]. Jerusalem [IEE89f]. Job [BBD88]. Job-Shop [BBD88]. Johnson [Ano85a, Ano86a]. Jose [IEE85a, IEE91d]. joy [FH11]. July [AAA93, ACM90a, ACM97, Ano94b, DH98, GA^{+}86, IEE87d, Küc97, Lev95, Nie93, RCG^{+}89, TPB^{+}88, Wat91c]. June [ACM90b, ACM92, ACM94b, Ano85a, Ano86a, Fre90, IEE86b, IEE87f, IEE88b, IEE88d, IEE88i, IEE89b, IEE89e, IEE89d, IEE91b, IEE92, IEE94a, MCH91, ACM93].

Kailua [HSKS88, HMS^{+}89, IEE88h]. Kailua-Kona [IEE88h]. Kaiserslautern [Nie93]. KB [Esa90]. Keene [Ano89b]. Kernel [MA91]. Keys [Pin85]. Kinder [Whi91]. Kit [Her93, Pur90b]. kit/programmer [Pur90b]. KNOWBOT [HK91]. Knowledge [BBD88, BHA89, CH93, Cor86, DKL95, GKS89, IEE92, LC95, OSD86, PRD88, Sig85, BWM89, BB90, Bur91, CTB91, DL93, DL98, Esa90, HR89, HK90, IEE85c, Rat93, RBLS98, SB89, SPM91, UJ92, WT91, WGS96]. Knowledge-Based [Cor86, OSD86, Sig85, BHA89, CH93, DKL95, GKS89, LC95, BB90, Bur91, DL93, DL98, HR89, HK90, IEE85c, RBLS98, SB89, WGS96]. KnowledgeWorks [Ano90d]. Kona [HSKS88, HMS^{+}89, IEE88h]. Kong [IEE88f, IEE90d]. Konover [IEE86c]. Konzeption [Flu91]. Krayola [SH96]. Kyoto [Giu87, YY94, Yua90].

L [Ano90b, Bla86]. L’Agelonde [PG92]. Lake [LS97]. Land [Gli86]. Language [ACM93, Ame92, Fat90, Giu87, HW87, HH93, IAC94, Kai86, KPB^{+}88, KA95, KB87, Que76, Ste84, Ste90a, Tau90, Tau91, AC94, AI96, DZB97, Gat97, Hol14, LQ94, Sim87, Sym86, Sym88, TAH89]. Languages [ACM86, IEE88a, IEE90b, IEE94c, MDM97, Fal88, Tat10]. Law [ACM95]. Layout [TW87, HK90, Sly93, SYM92]. learn [Hol14]. Learning [BS85, HBG96, Tat10]. lecons [HP96]. legal [VB95]. Level [Giu87, KPB^{+}88, KB87, MP87, Kar89, Moo12]. Leverage [IEE84b, IEE87b, Mul91]. lexical [Bas93, SST89]. Libaform [Que76]. Library [BKR^{+}91, HC85, KSv88, Coo95a, Coo95b, Vas97]. Lie [CT89]. like [LM89]. Line [Fol86, FIKY88, IEE87a, Tib93, VB95]. Linear [Bak93, McD89b, ND90, ND95]. link [Jel89, Ste91b]. linkage [Ste91b]. Links [BKR^{+}91]. Linz [Cav85]. LISP [AC94, Ano89b, Ano92, BDG^{+}88c, LM91, Pur90a, RMM93, Sev91, Whi91, Ano87b, Ano89a, Ano90b, Ano90e, Ano91b, Ano93, App90, Bob87, BDG^{+}88a, BDG^{+}88b, BDG^{+}88d, Chr93, Dem88, Flu91, HP96, Her93, HKLH93, Jac94, Kim90, Lou89, OK89, Ste90b, Sun88, Tou90a, Van98, You92, Hol14, ACM84, ACM90b, ACM92, ACM94b, Ame92, AI96, Acu89, Ake94, AP88, AM86, Bas93, Bla86, BGPS91, Bro85, Bro87, CL88, CT89, CT90, CR90, Chl91, DS86, Fra88a, Fra88b, GK90, Gra99, Gui88, HD89, Hen89, HSW85, HSW87, IAC94, JLHB92, KG89, KK93, MN91, ME95, NC88, Nor91, Nor92, Pav85, Pur90b, Que96, RMM94, Ril92, Sha86, Sha92, SM88, Sla98, SWH85, Ste86, Ste89, Ste91b, SST89, Tan90, Tat87, Thr88, Tou90b, WAF90, Wat91b, Wer89]. LISP [WS84, WS85, YH87, Abd90, Acu90, AH91, AM99, Ama91, AGK^{+}87, Ano90a, Ano91a, ABFM90, Bak92, BKK^{+}86, BR93, BCG90, Bra86, BK91, BG84, BW96, CK87a, CK87b, Chi89, Cla89, CLNW90, Cla92, CHD09, CF88, CB86, DB97, DD90, Dol90, Ecc88, Esa90, Fat90, Fat91, FL91, FBWR95, Flu91, Gad88, Giu87, GG89a, GG89b, Gra94, Gra96, Gra97, Har90b, Hen92b, Hum94, Ida90, JR90, JS94, JS97, Kai86, KUJP97, KT88, Kan91, Kar89, KM97, Kem91, KLF^{+}91, KA95, Kv89, Kla04, Koo87, Kos90, Kru85, KB87, KK22, Lam05, Lan88a, Lan89a, Lan86, Lei08, Lim87, Luc92a, Luc92b, Ma89, MK90, Mac91, Mac92a, Mac92b, MSD^{+}85, Mar04, McD87a, McD87b, MFW87, MFW88, McD89a, MK86, McK91, MBPLRR10, Mil88, MS20, Mu 95, MuT95, Mun95b]. Lisp [Mur90, NSI^{+}92, Nar88b, Neu03, Neu02, Nor04, Noy92, Nym87, OOT87, PKSA89, Pin85, PAR93, Que76, Rei96, Ric90, Ril91, SIG89, Sch94, Sei05, Sim87, SM93, SH96, Sny86a, Sny86b, Spe92a, Spe92b, Ste84, Ste90a, SG93, Swa90a, Sym86, Sym88, Tan95, Tau90, Tau91, THL^{+}86, Tei76, Ver19, VS90, Vog98, Wan86, War86, Wat89, Wat90, WNA90, WF84, WS87, Yan91, YY94, YYK^{+}90, Yua88, Yua90, Yua92, YYN90, ZHL^{+}89a, ZHL^{+}89b]. Lisp-based [Kar89, WNA90, CB86, Pin85, Que76, SH96]. Lisp-Language [Fat90]. LISP-Machine [Pav85]. Lisp-Systems [Flu91]. Lisp/CLOS [PAR93]. LISP/VMS [DS86]. LISPcraft [Wil86]. LispView [Mul91]. LispWorks [Ano90d, Ric90]. List [MSD^{+}85, Res92]. List-Evaluation [MSD^{+}85]. Lists [LLC89, LM89]. load [CH88]. Logic [Bak93, GM86, GKLM90, KM97, LQ94, LQ96, SM93, Thr88, WL95]. Logical [CM98]. Londe [PG92]. Londe-les-Maures [PG92]. Long [Tei88]. loop [Got91]. Loops [Wat91a, BKK^{+}86]. low [WT91].

m [PKSA89]. m-EVES [PKSA89]. Mach [McD89a]. Mach/IBM [McD89a]. Machine [HK90, MSD^{+}85, Sam77, SCP^{+}85, Sch94, Ste86, Wan86, WS87, CR89, EL89, Kv89, KM90, CK87a, Lim87, Pav85]. Machines [Giu87, VS90]. Macintosh [Ano90e, Ano91a, Ano91b, Ano93, App90, Lan89a, Wat90]. Macintosh/Lisp [Lan89a]. MACSYMA [HC85, Pav85]. main [Ruo88]. Mainframe [YYK^{+}90]. Mainstream [Ano87c]. maintainability [YJM91]. Maintaining [Mai91]. Maintenance [Had86, Nar88a]. Man [IEE89a, IEE91a, CR89]. man-machine [CR89]. Management [Boo84, Yua92, BPR^{+}94, FGGB89, GP97, GW98, KLSW94, Van98]. manager [Luc92a, McK91, RMM94, Vas97, HKLH93, RMM93]. managers [BB89]. Manipulating [FIKY88]. Manipulation [Que76]. manipulations [CT89, CT90]. Manipulator [Bur86]. mankind [JU91]. MANTRA [CTB91]. Manual [AIST88, Mu 95, MuT95, Mun95b, Lou89, Mac91, Mac92a, McD87a, McD89a, Pur90b, Spe92b, Sun88]. manufacturing [HG91, MHMR96, Ruo88, Tib93]. March [IEE84b, IEE88e, IEE90b, IEE93c, IEE93d]. marine [IEE88c]. mark [Zor90]. mark-and-sweep [Zor90]. Maryland [IEE88c]. Mash [McD87a]. Mash/IBM [McD87a]. Massachusetts [IEE89a, IEE93b, Ano89b]. Massively [Yua92, EL89]. Matching [Mu 95, MuT95, Mun95b]. material [Sly93]. Mathematica [Fat90]. Mathematica-to-Lisp [Fat90]. mathematical [Wan97]. mathematics [Fat97]. Maui [ACM97, Küc97]. Maures [PG92]. May [ACM95, IEE85b, IEE87a, IEE90a, IEE91c, IEE94c, K^{+}95, LS97, Tei88].

McESE [Chl91]. MCS [BK91]. MD [ACM95]. Meaningless [EWB87]. means [Lar96]. means-end [Lar96]. Mechanical [AIST88, CH86, CH87, IO85, DZB97]. Mechanically [Moo99]. mechanism [Sal90]. Med2 [Pup87]. Medical [Kin89, RBLS98]. Medicine [PO90]. Meeting [Ano87a]. meets [Sch94]. Melbourne [MDM97]. Memory [Nae84, Yua92, CH88, CHD09, SIG89]. mental [Mai91, Sch92]. Mentor [Had86]. merging [BKK^{+}86]. Meridien [IEE84b]. messy [GK90]. Meta [Moo12, BK91]. Meta-level [Moo12]. Metaobject [LZ97, KdR91]. metaprogramming [LZ94, LZ95]. Method [FN81, OK89, BWZ94, Got91]. methods [BGHW88, MN91, MSBH93]. Mexico [ACM93]. mGA1.0 [GK90]. Miami [IEE85c, IEE86c, IEE88a]. Micro [ACM87]. Micro-20 [ACM87]. microcode [Chi89, TS90]. Microcomputer [AP88, Nae84]. MicroExplorer [Lan89a]. microprocessor [BY96]. Microprogrammed [Kru85]. Microprogramming [ACM87]. MicroScope [ABC^{+}88]. MIDI [SH96]. Milner [Ano90b]. Minneapolis [BPR^{+}94, IEE87f]. Minnesota [BPR^{+}94]. misconceptions [Mai91]. missions [Dav88]. Mixed [MP87]. Mixed-Level [MP87]. MN [IEE87f]. mode [GH94]. Model [Pir86, AD88, KL89, Sha91]. modeling [BT94, Sev91]. Modelling [Vau85, RGG90, VB95]. models [Lar96, Mai91, Sch92, WNA90, You92]. Modems [CN91]. modular [FBB93]. Module [KSv88, SH96]. modules [Wat91b]. monitoring [Cha95, NRK89, NRK91]. monotonic [BT94]. Monsoon [Chi89]. Montréal [Lev95]. morphology [AD88]. Morristown [N^{+}90]. Mosaic [K^{+}95]. motion [KM90]. Move [Ano87c]. MPI [Coo95a, Coo95b]. Mulisp [Car84]. Multi [BBD88]. Multi-Knowledge-Based [BBD88]. multicomputers [CH88]. multiformalism [PAR93]. multimedia [HI94]. multiple [And97]. Multiprocessing [ZHL^{+}89b, ZHL^{+}89a]. multiprocessor [GHL^{+}90, Moo99]. multiprocessors [KVJ90]. multiprogramming [Bra86]. Multirobot [Fol86]. Multisensor [OSD86]. Multivariable [LK87]. Munich [IEE97a]. Music [Tau90, Tau91, Ano94a, Sch94, Tau90, Tau91].

NAECON [IEE85b, IEE91c]. NAG [BKR^{+}91]. NAND [Got91]. NASA [Ano85a, Ano86a]. NASA/Johnson [Ano85a, Ano86a]. National [AAA93, Ame92, Ano94b, IEE85b, IEE91c]. natural [TAH89]. Near [CH86, CH87]. nerve [Ste91a]. net [LM89]. net-lists [LM89]. nets [YYS95]. Network [Fat97, Pur90a, Pur90b]. networks [DH98, Har90a, Wat91b]. neural [SV97, Wat91b]. Nice [ACM90b, IEE89c]. Ninth [IEE93c]. NLH [TAH89]. NLH/E [TAH89]. non [BT94, KVJ90]. non-homogeneous [KVJ90]. non-monotonic [BT94]. Nondeterministic [SM93]. notes [Kan91]. Nov [MDM97]. November [IEE85a, IEE88c, IEE89a, IEE89g, IEE90e, IEE91d, IEE93a, IEE93b, IEE95a, IEE96, Kin89, PO90]. Nqthm [KM96]. Nuclear [PG92]. Numeric [BKR^{+}91, CH86, CH87, GHL^{+}90, Noy92].

O [Car84]. Oberoi [IEE89g]. OBJECT [LM91, ACM86, Ano87b, Ano89b, Bob87, BDG^{+}88a, BDG^{+}88b, BDG^{+}88c, BDG^{+}88d, Dem88, HP96, Her93, Hur91, IO85, KG89, KPB^{+}88, Mat90, MP87, McD89b, MDM97, Moo86, OK89, Pae93, Rat93, Sch92, Sev91, Sla98, BWM89, BCG^{+}88, BKK^{+}86, BHW91, BY96, BK91, Cla92, Cze90, DCSS88, FBB93, Fut90, GWB91, KK93, LQ94, LQ96, Nie93, Sny86b, You92]. Object-Orientated [ACM86]. Object-Oriented [Ano89b, KPB^{+}88, MP87, MDM97, Hur91, KG89, Mat90, McD89b, Moo86, Pae93, Rat93, Sch92, Sla98, BWM89, BCG^{+}88, BKK^{+}86, BHW91, Cze90, DCSS88, FBB93, GWB91, KK93, LQ94, LQ96, Nie93, Sny86b, You92]. Objectif [HP96]. objects [Pog96]. objet [Abd90]. Oblogis [GM86]. ocean [SM88]. Oceans [IEE88c]. October [ACM86, ACM88, Bän03, IEE86e, IEE87e, IEE88a, IEE88c, IEE91a, USE88]. ODYS [Mal91]. Omni [IEE91a]. On-line [VB95]. on-the-fly [Ehn89a, Ehn89b]. Online [GW98]. onto [DB97]. ontology [MTI95]. OODB [FBB93, XRS96]. OODCE [RZ96]. OOPSLA [ACM86]. Opal [BBD88]. Open [FBB93]. operating [Lam90]. operations [DKL95]. OPLA [LQ94, LQ96]. OPS5 [PM93]. OPS5-based [PM93]. Optimal [CH86, CH87]. Optimization [Web95, ZS93, GKS89, Hen92a, HK90]. optimizer [HG91]. Optimizing [AGK^{+}87, YYK^{+}90, YYN90]. order [PE88]. Orientated [ACM86]. Oriented [Ano89b, Bar84, KPB^{+}88, MP87, MDM97, SCP^{+}85, War86, BWM89, BCG^{+}88, BKK^{+}86, BHW91, Cla92, CHD09, Cze90, DCSS88, FBB93, GWB91, Hur91, KG89, KK93, LQ94, LQ96, Mat90, McD89b, Moo86, Nie93, Pae93, Rat93, Sch92, Sla98, Sny86b, You92]. Orlando [ACM94b, IEE93c, Jan90, Kna89]. Orleans [IEE90b, IEE90a]. ORM [Lei08]. Overview [Dem88, SH96, Sny86a]. own [Hol14].

P [Jac94, JS97]. P3D [WNA90]. PA [IEE87a, IEE89e]. Pacific [IEE89b]. Package [FL91, Esa90, LW95, SH96]. Packard [Cag86]. pages [Ano89a, Ano90b, Tou90a]. paint [SH96]. Palace [IEE92]. Pan [Lan88b, Lan88a]. paperback [Ano89a]. Papers [Ano87a, IEE84b, ACM84, ACM90b, ACM92, ACM94b, IEE87b, IEE87c, IEE87d, Wuo89]. Paradigms [Nor91, Nor92, Nor04, Swa90a, Swa90b, LQ94]. Parallel [DCSS88, GG89b, HJS86, JR90, MK86, Yua92, ACM97, Car90, Coo95a, Coo95b, EL89, GG89a, Wan97, WS87]. parallel/distributed [Wan97]. Parallelism [Mur90]. PaRAmeters [Sto94]. Park [ACM95]. parsing [RW90]. part [SB89, Tei88]. partnership [IEE88c]. parts [DLK92]. PASCO [ACM97]. passive [ME95]. Patching [DB97, Hum94]. path [Hal92, HG91]. Pattern [Fre90, IEE94a, Mu 95, MuT95, RW86, Lin91, Mun95b]. Pattern-Directed [RW86]. Paul [AM99]. payload [Dav88]. PC [McD87a, McD87b, MFW87, MFW88]. PCB [HG91]. PE [HWK95]. PE/SolidDesigner [HWK95]. Pennsylvania [IEE87d, IEE88g, MHG90, PO90]. Performance [Acu89, Acu90, Ehn89b, KB87, Mur90, Vog98, YYN90, BR93, CN91]. permutations [Res92]. Persistence [KPB^{+}88, LZ97, LZ94]. Persistent [JS94, JS97, Jac94]. Perspective [Bal85, GL94, Pae93]. perspectives [PO90]. Perth [IEE95a]. Petri [YYS95]. phase [ND90]. Philadelphia [IEE87a, IEE88g, MHG90, PO90]. philosophy [IEE90c]. Phoenix [IEE84a, IEE93a, NKC91]. Physics [BS85, PG92]. pieces [Que96]. piecewise [ND90, ND95]. piping [Jel89]. Pittsburgh [IEE87d, IEE89e]. place [Res92]. plan [GP97, Gat97, Sly93]. Planning [Ric94, ACW^{+}94, DLK92, Sly93]. plant [NSS94]. Platform [KPB^{+}88, Vas97]. Plaza [IEE88g, IEE91b, N^{+}90]. pLisp [Lan88b]. Point [FBWR95, Rei96, Vog98]. Polar [Flo99]. Polymorphic [Vos94]. polymorphism [YJM91]. Polynomials [Bak93, FL91, FIKY88]. Portable [Kan91, WW92, Bra86, BK91, NSI^{+}92]. Porting [Ecc88, Lan88b, Lan88a, PKSA89].

Portland [ACM86]. portraits [ND90]. Position [OSD86]. Potentials [MK86]. Power [Pav85, GH94, YYS95]. Practical [Sei05, Ste90b, VS10]. practices [Ano90c]. pragmatic [Tat10]. precomposition [SH96]. Predicate [Bak92]. Prentice [Ano90b]. preparation [HG91]. Preprocessor [DD90]. Present [Pav85]. Presented [Ano87a, ACM84, ACM90b, ACM92, ACM94b]. Press [Ano89a]. pressure [ME95]. pressured [And97]. pretty [Wat89]. Prevention [Hol92]. Preventive [Had86]. Primitive [Car84]. Princeton [K^{+}90]. printing [Wat89]. Probabilistic [Har90a]. Problem [KBS^{+}87, KK93]. Problems [SA86, ZS93]. Procedure [Bak92, Sto94]. Procedures [Sig85]. Proceeding [Ano94b]. Proceedings [AAA93, ACM86, ACM88, ACM90b, ACM90a, ACM92, ACM93, ACM94a, ACM94b, ACM97, AHWN97, Bän03, Fre90, HSKS88, HMS^{+}89, HMS^{+}90, HJS86, IEE85a, IEE85b, IEE86e, IEE86d, IEE87f, IEE87e, IEE88e, IEE88h, IEE88d, IEE88g, IEE88f, IEE89e, IEE89f, IEE89d, IEE89c, IEE90e, IEE91c, IEE93a, IEE94c, IEE95b, IEE96, Kna89, Küc97, K^{+}90, LS97, MHG90, MDM97, SA86, Ste90b, Tei88, ACM87, ACM95, Ano85a, Ano86b, Ano86a, BPR^{+}94, E^{+}93, GA^{+}86, IEE84a, IEE86a, IEE86b, IEE88b, IEE88a, IEE88c, IEE88i, IEE89a, IEE90c, IEE90d, IEE92, IEE93b, IEE93c, IEE94a, JU91, K^{+}95, Lev95, MCH91, NKC91, Nie93, PO90, PG92, RCG^{+}89, TPB^{+}88, USE88, Wat91c, Cav85, Kin89]. Process [Fol86, DLK92]. processes [NRK89, NRK91]. Processing [FBWR95, GG89b, Hum94, HJS86, IEE89b, IEE97a, RW86, Rei96, GG89a, Jan90, LLC89, Noy92]. Processor [KB87, MSD^{+}85, BGHW88, Kv89]. Processors [IEE87e, K^{+}90, Flu91]. Product [BGPS91]. production [HG91]. profile [BHW91]. profiling [Hal92]. Program [ABC^{+}88, CN91, Mu 95, MuT95, Nae84, Nar88a, Nar88b, Mun95a, Mun95b, ND90, RW90]. programmation [Abd90, Wer89]. Programmer [Ano89b, BDG^{+}88a, BDG^{+}88b, KG89, Pur90b, Tat87]. Programmieren [Bro87]. Programming [ACM84, ACM86, ACM90b, ACM92, ACM93, ACM94b, Ame92, Ano89b, Bal85, BSS84, Bro85, Hal09, IAC94, Kai86, KUJP97, KPB^{+}88, Kim90, Pir86, Ric90, Swa90a, Swa90b, Wat90, AC94, AI96, AH91, Ama91, BKK^{+}86, DCSS88, ECG12, GWB91, HD89, HBG96, Hol14, Kar89, KVJ90, KG89, Kem91, KLF^{+}91, KK93, LQ94, LQ96, Luc92b, McD89b, MCd^{+}94, Moo86, Nie93, Nor91, Nor92, Nor04, Pae93, Rat93, SM93, Sny86b, Tat10, Thr88, WS87, WL95, You92, Bro87]. Programs [Acu90, CM98, FIKY88, MK86, Ren86, Acu89, Kv89, Web95]. progress [Ruo88]. project [BB89, Bra86, Lan88a]. Prolog [CL88, GM86, MN91]. proof [Moo99]. proofs [BY96]. proposed [Ame92, AC94, IAC94]. Protocol [LZ97, KdR91]. Prototype [HKLH93, Mu 95, MuT95, Ano90a, Mun95b]. Prototyping [TBE^{+}94, BB90, Gui88]. prover [KM97, Moo12]. Prozessoren [Flu91]. pseudoexhaustive [SK90b]. PSX [ND90]. Publishing [Tou90a, Whi91]. Purpose [Giu87]. Putting [CK87b]. PVM [LW95]. pyramids [Fut90]. Python [Mac92b].

Qlisp [GG89a, GG89b]. Qualitative [Nis94, ND95, Sch92]. Quality [JR90]. Quickref [Ver19]. Quicksig [KH88].

radar [KK93]. radiation [LS97].

Radiotherapy [KUJP97]. rail [Got91]. Random [Res92]. Range [SK90a, SIG89]. rapid [Gui88]. rare [XXL92]. rating [KL95]. ratio [Lan89b]. Rational [FL91, Yan91]. RCXLisp [Kla04]. reactors [Lin91]. ready [Ano90f]. Real [AH91, Bro84, Ruo88, SCP^{+}85, Lam90, NRK89, NRK91, SM88]. Real-Time [Bro84, AH91, Ruo88, Lam90, NRK89, NRK91, SM88]. Realization [Pup87]. realm [Fal88]. reasonable [BR93]. reasoning [AM86, Koo87, Lou89, Nis94]. Reclamation [MCH91]. Recognition [Fre90, IEE94a]. Recognizing [RW90]. record [ACM84]. recurrences [BWZ94]. Recursion [Pir86]. Recursive [CM98]. Reduced [Bro84]. reference [Ano90e, Ano91b, App90, Fra88b, Lou89, Mar04, Ver19, Tou90a]. referenced [Sim87]. Reflections [LZ95]. Regency [IEE87f, IEE89a]. Region [IEE89g, IEE90d]. Register [CK87a]. relations [Thr88]. Release [Fra88a]. Reliable [Cla89, CLNW90, HW87]. reload [GKS89]. Remark [Rei96]. remote [Sto94]. Renaissance [AHWN97]. Repair [PRD88, LMM89a, LMM89b]. report [PKSA89]. Representation [Flo99, CTB91, Esa90, MA91, Rat93, Ste91b]. representations [Fat91]. Representing [Lub91, WNA90]. Requirements [Pup87, SB89]. Research [Cav85, Bra86, Lan88a, PG92]. Resort [IEE88a]. resource [GP97]. Restricted [Nae84]. result [Moo99]. reverse [Chi89]. Review [Ano89a, Ano89b, Ano90b, Gre84, Tou90a, Whi91]. Rim [IEE89b]. RNA [SK96]. roam [CF88]. ROBEXS [Ano85a, Ano86a]. Robot [OSD86]. robotic [LMM89b, MC90]. Robotics [Ano85a, Ano86a, IEE86d, IEE88g, Kla04, SMK^{+}88]. Robots [Ano86b]. robust [Gat97]. Rockefeller [IEE90e]. rod [Lin91]. rostering [CH93]. rotational [DLK92]. Roundup [BGPS91]. Royal [IEE88f]. RPC [Sto94]. RT [McD87a, McD87b, MFW87, MFW88, McD89a]. Rule [KM90, Lin91, MC90, HG91, Tib93]. Rule-based [KM90, Lin91, MC90, Tib93].

Run [SIG89]. Run-time [SIG89]. Rye [IEE87e].

S [Whi91]. SAC [Lan86]. SAC-2 [Lan86]. Salt [LS97]. San [ACM92, Ano89a, IEE84b, IEE85a, IEE86a, IEE86d, IEE87b, IEE87c, IEE88e, IEE89d, IEE91d, IEE93d, TPB^{+}88]. Sather [SO91]. scene [Mal91]. Scheduler [CB86, Sha91, Bra86, Dav88, LMM89b]. Scheduling [BBD88, DL93, DKL95, DL98, Sha91, SK90a]. schematic [LM89]. schematic-like [LM89]. Scheme [Shi88]. Science [IEE88h]. Sciences [HSKS88, HMS^{+}89, HMS^{+}90, Jan90]. Scientific [Bän03, Neu03, Neu02]. scope [SST89]. SDL [LS92a, LS92b]. SE [IEE92]. Seattle [Ano94b, IEE94a]. Second [Ano86a, IEE85c, IEE86c, IEE87b, IEE93d, PG92, ACM97, HMS^{+}89, IEE87c]. Section [MCH91, YYS95]. seisd [UJ92]. Selection [FBHJ87]. Semantics [JS94]. semiconductor [DKL95, DL98, Tib93]. semiregular [SK90b]. SENEX [BWM89]. sentential [CR89]. Sept [IEE94b]. September [ACM86, Ano94a, BPR^{+}94, IEE86a, IEE88f, IEE90c, IEE90d, Kna89, K^{+}90, MHG90]. sequences [BHA89]. Series [Wat91a, Wat87]. Server [Van98]. servers [Fat97]. service [JU91]. Services [Ano85b]. Set [WF84, YY94, Sun88]. Seven [Tat10]. Seventeenth [IEE87d, IEE93a]. Several [Acu90, Acu89]. Share [HC85]. Shell [BT94, Pup87, BCG^{+}88, CTB91, Lou89]. Sheraton [IEE87d, IEE88e, IEE90c, IEE90a]. Shop [BBD88]. Shuttle [Dav88, Ruo88]. SIGGRAPH [ACM88]. Signal [Hum94, IEE89b, IEE97a]. SIGPLAN [ACM93]. SIMD [Yua92]. Simple [FL91, IO85]. Simplification [Yan91]. Simulating [NC88]. Simulation [And97, AHWN97, E^{+}93, LS92a, LS92b, NKC91, Vau85, Bir85, PAR93, Sch92, Sev91, SMK^{+}88, Tib93, WGS96]. Simulation-based [LS92a, LS92b]. Simulator [Gli86, MP87, Bla86, SGW89]. Singapore [IEE91b]. single [Got91]. Singular [EL89]. Slisp [BP97]. small [Que96]. Smalltalk [EWB87]. SMART [BGHW88]. Society [Ano87a, IEE84b, IEE85c, IEE86c, IEE87b, IEE87c, IEE88c, IEE94a, MCH91, PO90]. Softcover [Ano90b, Tou90a]. Software [ACM88, Ano85b, IEE88h, IEE88i, IEE89c, IEE92, IEE93a, Kna89, PG92, BP97, CHD09, SYM92, WT91]. Solid [Wuo89]. Solid-State [Wuo89]. SolidDesigner [HWK95]. Solving [KK93]. Some [BKR^{+}91, XRS96]. Sonya [Ano89b]. South [ACM90a]. Southampton [SA86]. Space [Ano85a, Ano86a, Bro84, Ruo88]. Spacecraft [KBS^{+}87]. Sparse [Bak93]. spatial [Fut90]. spatial/object [Fut90]. SPE [Gad88]. Specific [K^{+}90]. Specification [BDG^{+}88a, BDG^{+}88b, CM98, OK89, BDG^{+}88c, BDG^{+}88d, DZB97, MSBH93, SYM92]. Speech [IEE97a]. sponsored [IEE85c, IEE88c, MCH91]. Spring [IEE84b, IEE87b, IEE87c]. Springs [ACM87]. Spur [THL^{+}86, ZHL^{+}89b]. SQL [CR90]. Square [IEE87d]. stages [Got91]. Stamford [IEE91b]. Standard [Ame92, AI96, IAC94, AC94, Ano87b, Bob87]. STAR [Coo95a, Coo95b]. STAR/MPI [Coo95a, Coo95b]. started [Ano91a]. State [Wuo89, ND95]. Static [Nar88a, Nar88b, Ake94, Lan92]. Station [IEE87d]. statistical [Kv89, Lou89]. Stock [Bro84]. stop [Zor90]. stop-and-copy [Zor90]. Straight [FIKY88]. Straight-Line [FIKY88]. Strategies [Fut90]. strength [KM96, KM97, Moo12]. stress [Ver19]. Strong [Ake94]. Structural [KSv88, HR89, UJ92]. structure [CR89, SK96]. STRUCTURELAB [SK96]. structuring [BB90]. Studies [XXL92, Nor91, Nor92, Nor04]. style [Lub91]. SUbprogram [Sto94]. substrate [SM93]. SUBTYPEP [Bak92]. Successful [Lam05]. Sun [Pur90a, Pur90b]. superalgebra [CT90]. Supercomputing [IEE90e]. supplies [GH94]. Support [HW87, KSv88, LZ97, Nar88b, BB89, GW98, LS92a, LS92b, NRK89, NRK91]. supporting [Gat97]. SUPRA [Sto94]. SUPRA-RPC [Sto94]. SVEPOA [PM93]. sweep [Zor90]. switched [GH94]. switching [DH98]. Symbolic [ACM94a, BKR^{+}91, CT89, CT90, CH86, CH87, Küc97, Lev95, Que76, Ren86, SGW89, TW87, Wat91c, Whi91, ACM97, Coo95a, Coo95b, Fat97, GHL^{+}90, Noy92, Tou90b]. Symbolic-Numeric [CH86, CH87]. SYMPAL [DCSS88]. Symposium [ACM84, ACM88, ACM94a, BPR^{+}94, IEE86b, IEE87d, IEE87a, IEE88d, IEE89f, IEE90a, IEE91b, Kin89, Küc97, Lev95, MHG90, Wat91c, ACM97]. synchronizable [Wat87]. Synchronous [TBE^{+}94]. Syntax [JS94, PE88]. Synthesis [CM98, KSv88, BGHW88, KL89, TS90]. System [BBD88, Boo84, CH86, CH87, FIKY88, GM86, HP96, Had86, HSKS88, HMS^{+}89, HMS^{+}90, IEE88h, JR90, KBS^{+}87, LK87, Ma89, MK90, Pin85, Pup87, Sig85, TSL^{+}88, Yan91, Yua92, AM86, BCG^{+}88, BK91, CL88, CFTZ89, CH93, Cla92, DL98, Ehn89b, Flu91, GKS89, GP97, GHL^{+}90, GH94, HI94, HR89, HBG96, Jel89, KH88, KLSW94, KL95, Koo87, Lam90, Lan89b, LS92a, LS92b, Lin91, LMM89a, LMM89b, LQ96, Mal91, MA91, MHMR96, NRK89, NRK91, NTS90, PK91, Sal90, SK96, SH96, SB89, Ste91a, Tib93, TAH89, TS90, UJ92, WAF90, Wat89, XXL92, YYS95, YJM91, LM91, BDG^{+}88a, BDG^{+}88b, BDG^{+}88c, BDG^{+}88d, BK91, Dem88, Her93, NP88, OK89, Sev91, You92]. Systems [ACM86, ACM90a, Acu90, Ano86a, Flu91, HW87, IEE86e, IEE87a, IEE88d, IEE88f, IEE88i, IEE89a, IEE90a, IEE90d, IEE91a, IEE91b, IEE93d, LK87, MDM97, N^{+}90, Pav85, PG92, SGBL86, War86, Acu89, Ano85a, BB90, BT94, Coo95a, Coo95b, Fat91, GL94, Hal90, HH93, Hol92, IEE85c, K^{+}95, LC95, LQ96, NSS94, ND95, Pog96, Sch92, SM88, SPM91, XRS96, You97, ZB91].

T [Ano89a]. Tachyon [NSI^{+}92]. tagging [Hen92a]. TAI [IEE91d, IEE93b]. Tallinn [JU91]. target [BHA89]. Task [MTI95, LC95, LMM89b]. TB [Kem91, KLF^{+}91]. teach [MSBH93]. technical [Wuo89]. Techniques [AM99, BB90, Mu 95, MuT95, Cha95, Fal88, Gra94, Mun95b, PG92, VS90]. Technologies [IEE84b, IEE89g, Jan90, PO90]. Technology [IEE88c, MDM97, Cag86, Lan89a, SV97, Tei88, AI96]. Tel [IEE88i]. Tel-Aviv [IEE88i]. Telecom [PG92]. temporal [Koo87]. TENCON [IEE89g]. TENCON’90 [IEE90d]. Test [Cor86, IEE86a, IEE90c, SMK^{+}88, Ver19]. Testing [IEE86a, Sam77, DKL95, DL98]. Texas [ACM84, IEE86a, IEE86e]. Texinfo [Ver19]. Their [Pup87]. theorem [KM97, Moo12]. theories [Sha91]. Theory [Ren86, Wat91b]. therapy [LS97]. thinning [Web95]. Third [ACM90a, HMS^{+}90, IEE88i, IEE91d, IEE84a]. Thirteenth [Kna89, Kin89]. Thirty [IEE87b, IEE87c]. Thirty-Second [IEE87b, IEE87c]. three [Got91, Kv89]. three-processor [Kv89]. TICL [Ma89, MK90]. TIG [Mu 95, MuT95, Mun95b]. Time [Bro84, SCP^{+}85, AH91, And97, Lam90, NRK89, NRK91, Ruo88, SIG89, SM88]. time-pressured [And97]. timed [YYS95]. TIMELOGIC [Koo87]. TMS320C30 [Kar89]. Tokyo [IEE86b]. Tolerant [IEE87d, Hol92]. tongues [Sch94]. Tool [SGBL86, YYN90, Bur91, MSBH93, NJB89, PM93, Pur90b]. toolbox [SH96, Wat90]. toolkit [BP97, Chr93, Kim90]. Toolkits [ABFM90]. Tools [ACW^{+}94, BGHW88, IEE91d, IEE93b, IEE96, VS90, Wan97, MDM97]. Toolset [ABC^{+}88]. Toulouse [IEE94c, IEE96]. Touretzky [Whi91]. Tower [IEE86d]. Towers [IEE89e, IEE89g]. Town [IEE87e]. Tracers [Her93]. Track [IEE88h]. tracking [BHA89, KK93]. Trading [Bro84]. training [SB89]. transactional [CHD09]. TransBase [Kem91, KLF^{+}91]. Transformation [Wat91a]. transformer [GH94]. translation [BW96]. Translator [Fat90]. Transliteration [Nym87]. trauma [GW98]. TraumaTIQ [GW98]. Trends [IEE88f]. Triennial [JU91]. Tutor [Pir86, Gui88]. Tutorial [KK93, Mil88, RK97, Ano90b]. tutoring [GL94]. Twelfth [Ano94b, PO90]. Twenty [HSKS88, HMS^{+}89, HMS^{+}90, IEE84b, IEE88h]. Twenty-eighth [IEE84b]. Twenty-First [HSKS88, IEE88h]. Twenty-second [HMS^{+}89]. Twenty-Third [HMS^{+}90]. Two [Acu90, You92, Acu89, ND90, ND95]. two-dimensional [ND90]. TX [IEE86a]. Type [Ma89, MK90, Ake94, Hen92a].

U.K [SA86]. U.S. [MCH91]. U.S.A [AHWN97]. UCL [Jac94, JS97, Lei08]. UCL-GLORP [Lei08]. unconstrained [GP97]. Undecidability [Lan92]. Understand [Lam05]. Understanding [AIST88, LM91, Fut90, KM94, Mun95a]. underwater [BCG90]. uniprocessor [Moo99]. universal [Flu91, Flu91]. Universal-Prozessoren [Flu91]. Universita [IEE94b]. University [IEE91a, IEE95a, MCH91, SA86]. Unix [Hen92b, Ric90]. US$19.95 [Ano89a]. USA [ACM86, Ano87a, Fre90, IEE87d, IEE88h, IEE90b, N^{+}90, PO90]. Use [Lam05, LS97, MSBH93, SB89]. used [BY96]. User [ACM88, DZB97, HWK95, Lou89, Luc92b, BT94, Cla92, Fra88a, Fra94, Fra96, IEE86a, Kan91, Luc92a, Mac91, Mac92a, McD87a, McD89a, Mye95, RGG90, Spe92b]. Using [BCG90, Fol86, LZ94, Mu 95, MuT95, Neu03, RZ96, YJM91, BW96, CT89, CT90, Chr93, Cla92, DL93, Got91, JLHB92, ME95, Mun95b, Neu02, Tan90, Tan95, Tib93, WL95, XRS96, YYS95]. USSR [JU91]. UT [LS97]. Utah [CK87a]. utilities [Kan91].

v [Lou89, Sch94]. vacuum [NTS90]. validation [PM93]. value [EL89]. variables [ND95]. VAX [DS86]. Vectors [Flo99]. Vehicle [Gli86, BCG90]. verification [PM93, You97]. verified [MBPLRR10]. Versatile [Bur91]. version [BK91, DL96, KM96, Spe92a, Spe92b, Lou89]. via [LZ97, Moo99]. Victoria [IEE89b]. view [Moo99]. Views [Vau85]. Virginia [IEE91a]. Virtual [CK87a, TW87]. Vision [IEE94a, Bur91, Fut90]. visual [Jan90]. VLSI [BGHW88, IEE87e, Kru85]. VMS [DS86]. Volume [Cav85, IEE88h]. vs [War86].

walking [KM90]. Washington [AAA93, Ano94b, IEE90c, IEE94a, Kin89]. water [Lin91]. Waverly [AHWN97]. Way [Ano87c]. WCL [Hen92b]. Web [DB97, RK97, RBLS98]. weeks [Tat10]. Welding [Ano87a, FBHJ87, Mu 95, MuT95, Mun95b]. Wendy [Ano90b]. Wesley [Ano89b, Tou90a]. Western [IEE95a]. Westin [IEE91b]. who [Mai91]. Wide [RBLS98]. widely [BY96]. window [Vas97, Pur90b]. Winter [AHWN97, E^{+}93, NKC91]. within [GL94]. Work [Ruo88]. workbench [Ama91]. workbook [Ril91, Ril92]. workcell [LMM89b]. Workshop [ACM87, Ano85a, Ano86a, IEE88f, PG92, DH98]. Workstation [Hum94, Cag86, SMK^{+}88].

Workstations [IEE85a, Ric90, Gad88]. World [JU91, RBLS98]. World-Wide [RBLS98]. Worth [IEE86e]. Worthington [IEE86e].

X [Ano89a, Chr93, Kim90, Luc92b].

X-based [Luc92b]. X3.226 [AC94]. X3J13 [AC94, BDG^{+}88c, BDG^{+}88d]. X3J13/93 [AC94]. X3J13/93-102 [AC94]. XLT [Luc92b]. XP [Wat89]. XVII [Sch94].

Year [Bal85]. York [IEE87e, IEE90e]. you’re [Ano90f]. Yuasa [Ano89a].

Z [Sha91]. Z-Scheduler [Sha91]. zoom [CF88].

References

AAAI:1993:PEN

[AAA93] AAAI, editor. Proceedings of the Eleventh National Conference on Artificial Intelligence, July 11–15, 1993, Washington, DC. MIT Press, Cambridge, MA, USA, 1993. CODEN PNAIEE. ISBN 0-262-51071-5. LCCN Q334.N36 1993.

Ambras:1988:MIP

[ABC^{+}88] James P. Ambras, Lucy M. Berlin, Mark L. Chiarelli, Alan L. Foster, Vicki O’Day, and Raldolph N. Splitter. MicroScope: an integrated program analysis toolset. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 39(4):71–82, August 1988. CO-DEN HPJOAX. ISSN 0018-

Abdulrab:1990:CLP

[Abd90] Habib Abdulrab. De Common Lisp à la programmation objet. Traite des Nouvelles Technologies, Serie Informatique. Hermés, Paris, France, 1990. ISBN 2-86601-244-5. 225 pp. LCCN ????

Attardi:1990:ICL

[ABFM90] G. Attardi, C. Bonini, T. Flagella, and F. F. Monai. Interfacing Common Lisp to graphical toolkits. In Steels [Ste90b], pages 253–264. ISBN ???? LCCN ????

ANSI:1994:DPA

[AC94] American National Standards Institute. Committee X3, Information Processing Systems and Computer and Business Equipment Manufacturers Association. Draft proposed American national standard programming language Common LISP. X3.226:199x. Draft 14.10, X3J13/93-102. Global Engineering Documents, Washington, DC, USA, January 10, 1994. ISBN ???? LCCN ????

ACM:1984:CRA

[ACM84]  ACM, editor.  Conference

record of the 1984 ACM Symposium on LISP and Functional Programming: papers presented at the symposium, Austin, Texas, August 6– 8, 1984. ACM Press, New York, NY 10036, USA, 1984. ISBN 0-89791-142-3. LCCN QA76.73.L23 A26 1984.

ACM:1986:OOO

[ACM86] ACM, editor. OOPSLA ’86. Object-Orientated Programming Systems, Languages and Applications. Conference Proceedings. Portland, OR, USA, 29 September–2 October 1986, volume 21(11) of ACM SIG-PLAN Notices. ACM Press, New York, NY 10036, USA, November 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

ACM:1987:MPA

[ACM87]  ACM, editor.  Micro-20:

proceedings of the 20th annual Workshop on Microprogramming, December 1–4, 1987, Colorado Springs, CO. ACM Press, New York, NY 10036, USA, 1987. CODEN MAWMDB. ISBN 0-89791-250-0. ISSN 0361-2163. LCCN ???? ACM order no. 520870. IEEE catalog no. 87TH0208-9. IEEE Computer Society order no. 831.

ACM:1988:PAS

[ACM88] ACM, editor. Proceedings of the ACM SIGGRAPH Symposium on User Interface Software, Banff, Alberta, Canada, October 17–19, 1988. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-283-7. LCCN QA76.9.U83A26 1988.

ACM:1990:PTI

[ACM90a] ACM, editor. Proceedings / the Third International Conference on Industrial and Engineering Applications of Artificial Intelligence and Expert Systems, IEA/AIE 90, July 15–18, 1990 … Charleston, South Carolina. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-372-8. LCCN Q334.I556 1990. Two volumes.

ACM:1990:PAC

[ACM90b] ACM, editor. Proceedings of the 1990 ACM Conference on LISP and Functional Programming: papers presented at the conference, Nice, France, June 27–29, 1990. ACM Press, New York, NY 10036, USA, 1990. ISBN 0-89791-368-X. LCCN QA 76.73 L23 A24 1990. ACM order no. 552900.

ACM:1992:PAC

[ACM92]  ACM, editor.  Proceedings

of the 1992 ACM Conference on LISP and Functional Programming: papers presented at the conference, San Francisco, California, June 22–24, 1992. ACM Press, New York, NY 10036, USA, 1992. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

ACM:1993:PAS

[ACM93] Proceedings of the ACM SIG-PLAN ’93 Conference on Programming Language Design and Implementation: Albuquerque, New Mexico, June 23– 25, 1993, SIGPLAN notices; ISSN: 0362-1340; v. 28, no. 6 (June 1993). ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-598-4. LCCN QA76.7.A152 v. 28, no. 6.

ACM:1994:IPI

[ACM94a] ACM, editor. ISSAC’94. Proceedings of the International Symposium on Symbolic and Algebraic Computation. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-638-7. LCCN QA76.95.I59 1994.

ACM:1994:PAC

[ACM94b] ACM, editor. Proceedings of the 1994 ACM Conference on LISP and Functional Programming: papers presented at the conference, Orlando, Florida, June 27–29, volume 7(3) of LISP Pointers. ACM Press, New York, NY 10036, USA, 1994. ISBN 0-89791-643-3. LCCN QA76.73.L23A26 1994. ACM order no. 522940.

ACM:1995:IPF

[ACM95] ACM, editor. ICAIL ’95 proceedings of the fifth International Conference on Artificial Intelligence and Law, May 21– 24 1995, College Park, MD. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-758-8. LCCN ???? URL

                                        http://www.acm.org/pubs/
                                        contents/proceedings/ai/
                                        222092/index.html.

ACM:1997:PPS

[ACM97] ACM, editor. PASCO ’97. Proceedings of the second international symposium on parallel symbolic computation, July 20–22, 1997, Maui, HI. ACM Press, New York, NY 10036, USA, 1997. ISBN ???? LCCN ????

Acuff:1989:PTC

[Acu89] Richard Acuff. Performance of two Common LISP programs on several systems. Report STAN KSL-89-02, Stanford University, Knowledge Systems Laboratory, Stanford, CA, USA, January 1989. 28 pp.

Acuff:1990:EPT

[Acu90] Richard Acuff. Endpaper: Performance of two Common Lisp programs on several systems. Lisp and Symbolic Computation, 3(2):171–204, April 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anderson:1994:TEP

[ACW^{+}94] Scott D. Anderson, Adam Carlson, David L. Westbrook, David M. Hart, and Paul R. Cohen. Tools for experiments in planning. Proceedings of the International Conference on Tools with Artificial Intelligence, pages 615–623, 1994. ISSN 1063-6730.

Arbitrio:1988:TCM

[AD88] Alessandro Arbitrio and Cristina Delogu. Towards a connectionist model of Italian morphology. Neural Networks, 1 (1 SUPPL):287–??, 1988. CO-DEN NNETEB. ISSN 0893-6080.

Anderson:1987:IOL

[AGK^{+}87] J. Wayne Anderson, William F. Galway, Robert R. Kessler, Herbert Melenk, and Winfried Neun. Implementing and optimizing Lisp for the Cray. IEEE Software, 4(4): 74–83, July 1987. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Allard:1991:RTP

[AH91] James R. Allard and Lowell B. Hawkinson. Realtime programming in Common Lisp. Communications of the ACM, 34(9):64–69, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/114679.html.

Andradottir:1997:PWS

[AHWN97] Sigrun Andradottir, Kevin J. Healy, David H. Withers, and Barry L. Nelson, editors. Proceedings of the 1997 Winter Simulation Conference: Renaissance Waverly Hotel, Atlanta, Georgia, U.S.A., 7–10 December 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN WSCPDK. ISBN 0-7803-4278-X (casebound), 0-7803-4279-8 (microfiche), 0-7803-4280-1 (CD-ROM). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1997. IEEE catalog number 97CB36141.

ANSI:1996:ANS

[AI96] American National Standards Institute and Information Technology Industry Council. American National Standard for Information Technology: programming language — Common LISP. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. various pp. Approved December 8, 1994.

Abe:1988:TUI

[AIST88] Norihiro Abe, Tomohiro Ishikawa, Sinji Sako, and Saburo Tsuji. Toward understanding of an instruction manual in mechanical assemblies. In IEEE [IEE88g], pages 1722–1727. ISBN ???? LCCN ???? Three volumes. IEEE catalog number 88CH2555-1. Computer Society order number 852.

Akers:1994:SST

[Ake94] Robert Lawrence Akers. Strong static type checking for functional Common LISP. Thesis (ph.d.), University of Texas at Austin, Austin, TX, USA, 1994. xii + 614 pp. Two volumes.

Allen:1986:HRS

[AM86] James F. Allen and Bradford W. Miller. The HORNE reasoning system in COMMON LISP. Technical report TR 126, Dept. of Computer Science, University of Rochester, Rochester, NY, USA, August 1986. 66 pp.

Alneami:1999:LAT

[AM99] Ahmed Alneami and Paul Mc Kevitt. On Lisp: Advanced techniques for Common Lisp. Paul Graham. The Artificial intelligence review, 13(3): 239–??, ???? 1999. CODEN AIRVE6. ISSN 0269-2821.

Amagai:1991:DLP

[Ama91] Yoshiji Amagai. Distributed Lisp programming workbench. NTT R&D, 40(5):679–686, 1991. CODEN NTTDEC. ISSN 0915-2326.

ANSI:1992:DPA

[Ame92] American National Standards Institute and Computer and Business Equipment Manufacturers Association, Washington, DC, USA. Draft proposed American National Standard Programming Language Common LISP, July 1, 1992. ???? pp.

Anderson:1997:SMT

[And97] Scott D. Anderson. Simulation of multiple time-pressured agents. In Andradottir et al. [AHWN97], pages 397–404. CODEN WSCPDK. ISBN 0-7803-4278-X (casebound), 0-7803-4279-8 (microfiche), 0-7803-4280-1 (CD-ROM). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1997. IEEE catalog number 97CB36141.

Anonymous:1985:RES

[Ano85a] Anonymous, editor. Robotics and expert systems, 1985; proceedings of ROBEXS ’85, the First Annual Workshop on Robotics and Expert Systems, NASA/Johnson Space Center, June 27–28, 1985. Instrument Society of America, Research Triangle Park, NC, USA, 1985. ISBN 0-87664-872-3. LCCN ????

Anonymous:1985:SS

[Ano85b] Anonymous. Software and services. ComputerWorld, 19(38): 50, September 23 1985. CO-DEN CMPWAB. ISSN 0010-4841.

Anonymous:1986:RES

[Ano86a] Anonymous, editor. Robotics and Expert Systems—1986: proceedings of ROBEXS ’86, the Second Annual Workshop on Robotics and Expert Systems, NASA/Johnson Space Center, June 4–6, 1986. Instrument Society of America, Research Triangle Park, NC, USA, 1986. ISBN 0-87664-966-5. LCCN TJ210.3.W66 1986.

Anonymous:1986:RCP

[Ano86b] Anonymous, editor. Robots 10:

conference proceedings, April 20–24, 1986, Chicago, Illinois. Robotics International of SME, Dearborn, MI, USA, 1986. ISBN 0-87263-223-7. LCCN TS191.8.R637 1986.

Anonymous:1987:APP

[Ano87a] Anonymous, editor. Abstracts of Papers Presented at 1987 AWS Convention: 68th American Welding Society Annual Meeting and 18th International AWS Brazing Conference. 1987 Chicago, IL, USA. American Welding Society, Miami, FL, USA, 1987. ISBN 0-87171-277-6. LCCN ????

Anonymous:1987:CLO

[Ano87b] Anonymous. The Common Lisp Object Standard (CLOS), October 27, 1987. 1 videocassette (VHS) (53 min.).

Anonymous:1987:NWM

[Ano87c] Anonymous. New way to move AI into the mainstream. Electronics, 60(2):90–92, January 1987. CODEN ELECEH. ISSN 0883-4989.

Anonymous:1989:BRC

[Ano89a] Anonymous. Book review: Common Lisp drill: T. Yuasa: Academic Press, San Diego, 1988, 232 pages (paperback), US$19.95, ISBN 0-12-774861-X. Mathematics and Computers in Simulation, 31(1–2):143, February 1989. CODEN MC-SIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL

                                        http://www.sciencedirect.
                                        com/science/article/pii/
                                        0378475489900748.

Anonymous:1989:BRO

[Ano89b] Anonymous. Book review: Object-Oriented Programming in Common LISP: A Programmer’s Guide to CLOS, Sonya E. Keene, Massachusetts: Addison-Wesley, 1989. Lisp and Symbolic Computation, 2 (2):163–165, June 1989. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1990:ACL

[Ano90a]  Anonymous.  Appendix E.

the Common Lisp code of ‘CIMVIEW’-prototype. Acta Polytechnica Scandinavica, Mathematics and Computer Science Series, 55:109–??, 1990. CO-DEN ASMSD7. ISSN 0355-2713.

Anonymous:1990:BRC

[Ano90b] Anonymous. Book review:

Common Lisp: A Tutorial, by Wendy L. Milner. Prentice Hall, 1988. softcover; 519 pages. Lisp and Symbolic Computation, 3(3):321– 325, September 1990. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Anonymous:1990:CP

[Ano90c] Anonymous. Computing practices. Communications of the ACM, 33(11):87–97, November 1990. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Anonymous:1990:KDL

[Ano90d] Anonymous. KnowledgeWorks, DataWorks, & LispWorks 2.0.

The AI Magazine, pages 31–32, 1990. CODEN AIMAEK. ISSN 0738-4602.

Anonymous:1990:MAC

[Ano90e] Anonymous. Macintosh Allegro Common Lisp 1.3 reference. Apple Computer, Cupertino, CA, USA, version 1.3.2 edition, 1990. ISBN ???? xxi + 413 pp. LCCN ????

Anonymous:1990:WYR

[Ano90f]  Anonymous.  When you’re

ready for the future. The AI Magazine, pages 47–49, 1990. CODEN AIMAEK. ISSN 0738-4602.

Anonymous:1991:GSM

[Ano91a] Anonymous. Getting started with Macintosh Common Lisp: draft. Developer Technical Publications, Cupertino, CA, USA, version 2.0. edition, 1991. ISBN ???? 8 + 4 + xiii + 95 pp. LCCN ????

Anonymous:1991:MCL

[Ano91b] Anonymous. Macintosh Common Lisp 2.0 reference: draft. Developer Technical Publications, Cupertino, CA, USA, 1991. ISBN ???? xxii + 535 pp. LCCN ????

Anonymous:1992:ACL

[Ano92] Anonymous. Apple Common LISP 2.0. PC Week, 9(38): 99–??, September 1992. ISSN 0740-1604.

Anonymous:1993:MCL

[Ano93] Anonymous. Macintosh Common Lisp 2.0. Macworld, 10 (10):65–??, October 1993. CO-DEN MACWEA. ISSN 0741-8647.

Anonymous:1994:IIC

[Ano94a] Anonymous, editor. ICMC 94:

International computer music conference: September 1994, Århus, Denmark. DIEM (Danish Institute of Electroacoustic Music), Århus, Denmark, 1994. ISBN ???? LCCN ???? Also published as Stanford University, Center for Computer Research in Music and Acoustics, Report no. STAN-M-89.

Anonymous:1994:PTN

[Ano94b] Anonymous, editor. Proceeding of the Twelfth National Conference on Artificial Intelligence: July 31–August 4, 1994, Seattle, Washington. MIT Press, Cambridge, MA, USA, 1994. CODEN PNAIEE. ISBN 0-262-61102-3. LCCN Q334.N36 1994. Two volumes.

Alberts:1988:MAA

[AP88] L. R. Alberts and Y. C.

Pao. Microcomputer applications of AUTOLISP and COMMON LISP. In Tipnis et al. [TPB^{+}88], pages 447– 453. LCCN TA 345 A86 1988. Three volumes.

Apple:1990:MAC

[App90] Apple Computer, Cupertino, CA, USA. Macintosh Allegro Common Lisp 1.3 reference, version 1.3.2. edition, 1990. xxi + 413 pp.

Baker:1992:DPC

[Bak92] Henry G. Baker. A decision procedure for Common Lisp’s SUBTYPEP predicate. Lisp and Symbolic Computation, 5 (3):157–190, September 1992. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Baker:1993:SPL

[Bak93]  Henry G. Baker.  Sparse

polynomials and linear logic. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 27(4):10–14, December 1993. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Balzer:1985:YPA

[Bal85] Robert Balzer. 15 year perspective on automatic programming. IEEE Transactions on Software Engineering, SE-11 (11):1257–1268, October 1985. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Bansch:2003:CSC

[Bän03] Eberhard Bänsch, editor. Challenges in Scientific Computing — CISC 2002: Proceedings of the Conference Challenges in Scientific Computing, Berlin, October 2–5, 2002, volume 35 of Lecture Notes in Computational Science and Engineering. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNCSA6. ISBN 3-642-62406-5 (print), 3-642-19014-6 (e-book). ISSN 1439-7358. LCCN ???? URL http:

//link.springer.com/book/
10.1007/978-3-642-19014-
8;   http://www.springerlink.
com/content/978-3-642-19014-
8.

Barstow:1984:DEI

[Bar84] David R. Barstow. Displayoriented editor for Interlisp. In Interactive Programming Environments [BSS84], pages 288– 299. ISBN 0-07-003885-6. LCCN QA76.6 .I5251 1984. US$34.95.

Basu:1993:LAG

[Bas93] Anand Basu. A lexical analyzer generator in COMMON LISP. Thesis (m.s.c.e.), Wright State University, Dayton, OH, USA, 1993. vii + 56 pp.

Bimson:1989:AMP

[BB89] Kent D. Bimson and Linda Boehm Burris. Assisting managers in project definition: foundations for intelligent decision support. IEEE expert: intelligent systems and their applications, 4 (2):66–76, Summer 1989. CO-DEN IEEXE7. ISSN 0885-9000.

Bimson:1990:EPT

[BB90] Kent D. Bimson and Linda Boehm Burris. Evolutionary prototyping: Techniques for structuring the iterative development of knowledge-based systems. In Hoevel et al. [HMS^{+}90], pages 211–219. CO-DEN PHISD7. ISBN 0-8186-2008-0 (paperback: vol. 1) 0-8186-2009-9 (paperback: vol.

2) 0-8186-2010-2 (paperback: vol. 3) 0-8186-2011-0 (paperback: vol. 4) 0-8186-9008-9 (hard: vol. 1) 0-8186-9009-7 (hard: vol. 2) 0-8186-9010-0 (hard: vol. 3) 0-8186-9011-9 (hard: vol. 4) 0-8186-6008-2 (microfiche: vol. 1) 0-8186-6009-0 (microfiche: vol. 2) 0-8186-6010-4 (microfiche: vol. 3) 0-8186-6011-2 (microfiche: vol. 4). ISSN 0073-1129. LCCN QA76.9.S88H38 1990. Four volumes. Computer Society order no.: 2008 (vol. 1), 2009 (vol. 2), 2010 (vol. 3), 2011 (vol. 4); IEEE catalog no.: 90TH0293-1 (vol. 1), 90TH0294-9 (vol. 2), 90TH0295- 6 (vol. 3), 90TH0296-4 (vol. 4).

Bensana:1988:OMK

[BBD88] E. Bensana, G. Bel, and D. Dubois. Opal: a multiknowledge-based system for industrial job-shop scheduling. International Journal of Production Research, 26(5):795– 819, May 1988. CODEN IJPRB8. ISSN 0020-7543.

Ballou:1988:CES

[BCG^{+}88] Nat Ballou, Hong-Tai Chou, Jorge F. Garza, Won Kim, Charles Petrie, David Russinoff, Donald Steiner, and Darrell Woelk. Coupling an expert system shell with an object-oriented database system. Journal of Object Oriented Programming, 1(2):12– 21, June–July 1988. CODEN JOOPEC. ISSN 0896-8438.

Bowen:1990:UCL

[BCG90] Paul S. Bowen, Steven G.

Chappell, and Roger Gonzalez. Using Common Lisp in the EAVE autonomous underwater vehicle. IEEE Journal of Oceanic Engineering, 15(3): 221–227, July 1990. CODEN IJOEDY. ISSN 0364-9059.

Bobrow:1988:CLOa

[BDG^{+}88a] Daniel G. Bobrow, Linda G.

Demichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common Lisp Object System specification 1. programmer interface concepts. Lisp and Symbolic Computation, 1(2):245– 298, September 1988. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bobrow:1988:CLOb

[BDG^{+}88b] Daniel G. Bobrow, Linda G.

Demichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common Lisp Object System specification 2. functions in the programmer interface. Lisp and Symbolic Computation, 1 (2):299–394, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Bobrow:1988:CLOc

[BDG^{+}88c] Daniel G. Bobrow, Linda G. DeMichiel, Richard P. Gabriel, Sonya E. Keene, Gregor Kiczales, and David A. Moon. Common LISP Object System specification X3J13 Document 88-002R. ACM SIGPLAN Notices, 23(SI):1–142, September 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Bobrow:1988:CLOd

[BDG^{+}88d] Daniel Gureasko Bobrow, L. G. DeMichiel, R. P. Gabriel, S. E. Keene, G. Kicsales, and D. A. Moon. Common Lisp Object System specification: X3J13 document 88-002R, volume 23 of ACM SIGPLAN Notices. ACM Press, New York, NY 10036, USA, September 1988. ISBN 0-89791-289-6. 48 + 94 pp. LCCN QA76.73.C28 C65 1988 Sci-Eng. ACM order number 548883.

Brooks:1984:CCL

[BG84] Rodney A. Brooks and Richard P. Gabriel. Critique of Common Lisp. In ACM [ACM84], pages 1–8. ISBN 0-89791-142-3. LCCN QA76.73.L23 A26 1984.

Bergstraesser:1988:STM

[BGHW88] T. Bergstraesser, J. Gessner, K. Hafner, and S. Wallstab. SMART: Tools and methods for synthesis of VLSI chips with processor architecture. In IEEE [IEE88b], pages 654– 657. CODEN PDAWDJ. ISBN 0-8186-8864-5. ISSN 0146-7123. LCCN TA 174 D46p 1988. IEEE catalog number 88CH2540-3.

Brinsmead:1991:CLP

[BGPS91] Mark Brinsmead, Ken Gamble, Michael Pazzani, and Glenn Silverstein. Common LISP product roundup. AI Expert, 6(6):48–??, June 1991. ISSN 0888-3785.

Bronskill:1989:KAD

[BHA89] John F. Bronskill, John S. A.

Hepburn, and Wing K. Au. Knowledge-based approach to the detection, tracking and classification of target formations in infrared image sequences. In IEEE [IEE89d], pages 153–158. ISBN 0-8186-1952-X (paperback), 0-8186-8952-8 (hardcover), 0-8186-5952-1 (microfiche). LCCN TA1632 .I36 1989.

Boning:1991:IPI

[BHW91] Duane S. Boning, Michael L.

Heytens, and Alexander S. Wong. The intertool profile interchange format: An objectoriented approach. IEEE Transactions on Computer-Aided Design of Integrated Circuits and Systems, 10(9):1150– 1156, September 1991. CO-DEN ITCSDI. ISSN 0278-0070 (print), 1937-4151 (electronic).

Birtwistle:1985:AGS

[Bir85] Graham Birtwistle, editor. AI, graphics, and simulation. Society for Computer Simulation, La Jolla, CA, USA, 1985. ISBN ???? LCCN Q334 .A47 1985.

Bretthauer:1991:MCS

[BK91] Harry Bretthauer and Jürgen Kopp. The Meta Class System, MCS: a portable object system for Common Lisp version 1.3: documentation. Arbeitspapiere der GMD 554, Gesellschaft für Mathematik und Datenverarbeitung mbH, Sankt Augustin, Germany, July 1991. 33 pp.

Bobrow:1986:CLM

[BKK^{+}86] D. G. Bobrow, K. Kahn, G. Kiczales, L. Masinter, Stefik, M., and F. Zdybel. Common Loops, merging Lisp and object-oriented programming. ACM SIGPLAN Notices, 21 (11):17–29, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Broughan:1991:SSC

[BKR^{+}91] K. A. Broughan, G. Keady, T. D. Robb, M. G. Richardson, and M. C. Dewar. Some symbolic computing links to the NAG numeric library. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(3): 28–37, July 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Blair:1986:SAC

[Bla86] Michael R. Blair. A simulator for the L architecture in COM-MON LISP. Thesis (b.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1986. 1 + 49 pp. Supervised by Stephen A. Ward.

Bobrow:1987:CLO

[Bob87] Daniel Gureasko Bobrow. The Common Lisp Object Standard, 1987. 1 videocassette (53 min.).

Boose:1984:EES

[Boo84] John H. Boose. Extended expert system for display management in a CADD system. In IEEE [IEE84a], pages 41– 44. CODEN CPACE3. ISBN 0-8186-0555-3. LCCN TK 7885 A1 I567 1984.

Brinkley:1997:SFS

[BP97] James F. Brinkley and Jeffrey S. Prothero. Slisp: a flexible software toolkit for hybrid, embedded and distributed applications. Software—Practice and Experience, 27(1):33– 48, January 1997. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Bocks:1994:EDM

[BPR^{+}94] P. Bocks, B. Prasad, R. Rangan, et al., editors. Engineering data management: integrating the engineering enterprise: proceedings of the 1994 ASME Database Symposium, September 11–14, 1994 Minneapolis, Minnesota. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1994. ISBN 0-7918-1281-2. LCCN TA345 .A85 1994.

Boreczky:1993:BCL

[BR93] John Boreczky and Lawrence A. Rowe. Building Common Lisp applications with reasonable performance. Report UCB/CSD 93/763, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, June 1993. 10 pp.

Brand:1986:PMS

[Bra86] Russell L. Brand. A portable multiprogramming scheduler for Common Lisp: research project. Master of science, plan ii, University of California, Berkeley, Department of Computer Science, Berkeley, CA, USA, 1986. 29 pp.

Brooks:1984:TDS

[Bro84] Rodney A. Brooks. Trading data space for reduced time and code space in real-time garbage collection on stock hardware. In ACM [ACM84], pages 256– 262. ISBN 0-89791-142-3. LCCN QA76.73.L23 A26 1984.

Brooks:1985:PCL

[Bro85] Rodney Allen Brooks. Programming in Common LISP. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1985. ISBN 0-471-81888-7. xv + 303 pp. LCCN QA76.73.L23B76 1985. US$18.95 (est.).

Brooks:1987:PCL

[Bro87]  Rodney A. Brooks.  Pro-

grammieren in Common LISP. (German) [Programming in Common LISP]. Oldenbourg-Verlag, München, Wien, Oldenbourg, 1987. ISBN ???? xiii + 318 pp. LCCN ???? Translated by Dieter Bielski.

Brown:1985:ILP

[BS85] Stephen F. Brown and Kathleen M. Swigger. Intelligent learning of physics. In Anonymous [Ano85a], pages 243–248. ISBN 0-87664-872-3. LCCN ????

Barstow:1984:IPE

[BSS84] David R. Barstow, Howard E.

Shrobe, and Erik Sandewall. Interactive Programming Environments. McGraw-Hill, New York, NY, USA, 1984. ISBN 0-07-003885-6. xii + 609 pp. LCCN QA76.6 .I5251 1984. US$34.95.

Brajnik:1994:SDN

[BT94] Giorgio Brajnik and Carlo Tasso. Shell for developing nonmonotonic user modeling systems. International Journal of Human Computer Studies, 40 (1):31–62, January 1994. ISSN 1071-5819.

Burdick:1986:AGE

[Bur86] Joel W. Burdick. Algorithm for generation of efficient manipulator dynamic equations. In IEEE [IEE86d], pages 212–218. ISBN 0-8186-0695-9. LCCN TJ210.3 .I44 1986. Three volumes. IEEE Computer Society order number 695.

Burger:1991:VDT

[Bur91] W. Burger. Versatile development tool for knowledge-based computer vision. Cybernetics and Systems, 22(3):313– 329, May–June 1991. CODEN CYSYDH. ISSN 0196-9722.

Broughan:1996:FLT

[BW96] K. A. Broughan and D. M. K. Willcock. Fortran to Lisp translation using f2cl. Software —Practice and Experience, 26 (10):1127–1139, October 1996. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Ball:1989:SOB

[BWM89] Sheldon Ball, Lawrence Wright, and Perry Miller. SENEX, an object-oriented biomedical knowledge base. In Kingsland III [Kin89], pages 85–89. CO-DEN PCMCDC. ISBN 0-8186-1985-6. ISSN 0195-4210. IEEE Catalog Number 89TH0286-5.

IEEE Computer Society order number 1985.

Bachmann:1994:CRM

[BWZ94] Olaf Bachmann, Paul S. Wang, and Eugene V. Zima. Chains of recurrences — a method to expedite the evaluation of closed-form functions. In ACM [ACM94a], pages 242–

249. ISBN 0-89791-638-

7. LCCN QA76.95.I59 1994.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/190347/p242-bachmann/

.

Boyer:1996:APO

[BY96] Robert S. Boyer and Yuan Yu. Automated proofs of object code for a widely used microprocessor. Journal of the ACM, 43(1):166–192, January 1996. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Cagan:1986:IHP

[Cag86] Martin R. Cagan. An introduction to Hewlett–Packard’s AI workstation technology. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 37(3):4–14, March 1986. CODEN HPJOAX. ISSN 0018-1153.

Carter:1984:APF

[Car84] Michael Carter. Adding primitive I/O functions to Mulisp. Dr. Dobb’s Journal of Software Tools, 9(11):18–24, November 1984. CODEN DDJSDM. ISSN 1044-789X.

Carlotto:1990:DAI

[Car90] Mark J. Carlotto. Data-parallel algorithms for image computing. In Janota [Jan90], pages 136–149. CODEN PSISDG. ISBN 0-8194-0352-0. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QC871 .D54 1990.

Caviness:1985:PEE

[Cav85] Bob F. Caviness, editor. Proceedings: EUROCAL ’85, European Conference on Computer Algebra, Linz, Austria, April 1–3, 1985. Volume 2. Research Contributions, volume 204 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. CODEN LNCSD9. ISBN 0-387-15984-

3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0204.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=204.

Cunningham:1986:LBH

[CB86] P. Cunningham and J. Browne. Lisp-based heuristic scheduler for automatic insertion in electronics assembly. International Journal of Production Research, 24(6):1395– 1408, November–December 1986. CODEN IJPRB8. ISSN 0020-7543.

Cousins:1988:GBZ

[CF88] Steven B. Cousins and J. Andrew Fingerhut. A graph browser with zoom and roam for Allegro Common Lisp. Technical report WUCS-88-19, Washington University, Dept. of Computer Science, Saint Louis, MO, USA, 1988. 18 pp.

Cherubini:1989:IES

[CFTZ89] M. Alessandra Cherubini, Laura Fanti, Piero Torrigiani, and Massimo Zallocco. Integrated expert-system builder. IEEE Software, 6(6):44–52, November 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Chieng:1986:IHS

[CH86] W.-H. Chieng and D. A.

Hoeltzel. Interactive hybrid (symbolic-numeric) system approach to near optimal design of mechanical components. In Gupta et al. [GA^{+}86], pages 149–159. CODEN COENEF. ISBN ???? LCCN TA 345 A86 1986. Three volumes.

Chieng:1987:IHS

[CH87] Wei-Hua Chieng and David A.

Hoeltzel. Interactive hybrid (symbolic-numeric) system approach to near optimal design of mechanical components. Engineering with Computers, 2 (2):111–123, 1987. CODEN ENCMEH. ISSN 0177-0667 (print), 1435-5663 (electronic).

Chowkwanyun:1988:HDL

[CH88] Raymond Chowkwanyun and Kai Hwang. Hybrid dynamic load balancing for distributedmemory multicomputers. In IEEE [IEE88f], pages 391–399. ISBN 0-8186-0897-8. LCCN QA76.9.D5W67 1988. IEEE catalog number 88TH0228-7.

Chow:1993:KBS

[CH93] K. P. Chow and C. K. Hui. Knowledge-based system for rostering. Expert Systems with Applications, 6(3):361– 375, July–September 1993. CODEN ESAPEH. ISSN 0957-4174. Chakravarthy:1995:AMT

[Cha95] S. Chakravarthy. Architectures and monitoring techniques for active databases: an evaluation. Data and Knowledge Engineering, 16(1):1–26, July 1995. CODEN DKENEW. ISSN 0169-023X (print), 1872-6933 (electronic).

Costanza:2009:COS

[CHD09] Pascal Costanza, Charlotte Herzeel, and Theo D’Hondt. Context-oriented software transactional memory in Common Lisp. ACM SIGPLAN Notices, 44(12):59–68, December 2009. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Chiou:1989:RCM

[Chi89] Derek Chiou. A reverse compiler: Monsoon dataflow microcode to Common Lisp. Thesis (b.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1989. 43 pp. Supervised by Gregory Michael Papadopoulos.

Chlobowski:1991:MEA

[Chl91] Andrzej Chlobowski. McESE extension of Allegro COM-MON LISP. M.Sc. Project, McMaster University, Hamilton, ON, Canada, 1991. x + 109 pp.

Christenson:1993:DIT

[Chr93] Dee Ann Christenson. The design and implementation of a toolkit using the Common Lisp X interface. Thesis (m.a., computer science), University of South Dakota, Vermillion, SD, USA, 1993. 91 pp.

Carr:1987:EUC

[CK87a] Harold Carr and Robert R.

Kessler. Emulator for Utah Common Lisp’s Abstract Virtual Register Machine. Journal of FORTH Application and Research, 5(1), 1987. CODEN JFAREL. ISSN 0738-2022.

Carr:1987:PLF

[CK87b] Harold Carr and Robert R.

Kessler. Putting Lisp on Forth base. Journal of FORTH Application and Research, 5(1):??, 1987. CODEN JFAREL. ISSN 0738-2022.

Cattaneo:1988:CLI

[CL88] Giuseppe Cattaneo and Vincenzo Loia. A Common-LISP implementation of an extended Prolog system. ACM SIG-PLAN Notices, 23(4):87–102, April 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Clamen:1989:RDC

[Cla89]  S. M. Clamen.  Reliable

distributed computing with Avalon/Common Lisp. Research paper CMU-CS-89-186, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, September 21, 1989. 18 pp.

Clarkson:1992:IUI

[Cla92] M. E. Clarkson. Intelligent user interface for computer algebra using a Common Lisp object oriented expert system. In Perret-Gallix [PG92], pages 727–?? ISBN 981-02-1122-8. LCCN QC793.47.E4 I58 1992; Q334.I595 1992.

Clamen:1990:RDC

[CLNW90] Stewart M. Clamen, Linda D. Leibengood, Scott M. Nettles, and Jeannette M. Wing. Reliable distributed computing with Avalon/Common Lisp. In IEEE [IEE90b], pages 169–179. ISBN 0-8186-2036-6. LCCN QA76.7 .I576 1990. IEEE Computer Society Press order number 2036. IEEE catalog number 90CH2854-8.

Chazarain:1998:ASR

[CM98] Jacques Chazarain and Serge Muller. Automated synthesis of recursive programs from a ∀∃ logical specification. Journal of Automated Reasoning, 21(2): 233–275, October 1998. CO-DEN JAREEW. ISSN 0168-7433 (print), 1573-0670 (elec-

tronic). URL http://link.
springer.com/article/10.
1023/A%3A1005903504159.

Clark:1991:PDE

[CN91] G. C. Clark and M. D. Noakes.

Program development environment for high performance HF modems. IEE conference publication, 339:265–268, 1991. CO-DEN IECPB4. ISSN 0537-9989.

Cooperman:1995:SBP

[Coo95a] G. Cooperman. STAR/MPI:

binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Cooperman:1995:SMB

[Coo95b] Gene Cooperman. STAR/MPI:

binding a parallel library to interactive symbolic algebra systems. In Levelt [Lev95], pages 126–132. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Cornell:1986:KAT

[Cor86] Matthew Cornell. Knowledgebased automatic test equipment. In Anonymous [Ano86a], pages 97–104. ISBN 0-87664-966-5. LCCN TJ210.3.W66 1986.

Chandrasekar:1989:ICS

[CR89] R. Chandrasekar and S. Ramani. Interactive communication of sentential structure and content: An alternative approach to man-machine communication. International Journal of Man-Machine Studies, 30(2):121–148, February 1989. CODEN IJMMBC. ISSN 0020-7373.

Charness:1990:CSC

[CR90] Dave Charness and Lawrence A. Rowe. CLING/SQL: Common LISP to INGRES/SQL interface. Memorandum UCB/ERL M90/40, Electronics Research Laboratory, College of Engineering, University of California, Berkeley, Berkeley, CA, USA, May 8, 1990. 28 pp.

Cecchini:1989:SLA

[CT89] R. Cecchini and M. Tarlini. Symbolic Lie algebras manipulations using COM-MON LISP. Computer Physics Communications, 52 (2):283–289, January/February 1989. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
0010465589900106.

Cecchini:1990:SSM

[CT90] R. Cecchini and M. Tarlini. Symbolic superalgebra manipulations using COMMON LISP. Computer Physics Communications, 60(2):265–270, September 1990. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL

http://www.sciencedirect.
com/science/article/pii/
001046559090013Q.

Calmet:1991:MSH

[CTB91] J. Calmet, I. A. Tjandra, and G. Bittencourt. MANTRA: a shell for hybrid knowledge representation. In IEEE [IEE91d], pages 164–171. ISBN 0-8186-2300-4 (paperback), 0-8186-2302-0 (casebound), 0-8186-2301-2 (microfiche). LCCN Q334 .I563 1991. IEEE Computer Society Press Number 2300. IEEE Catalog Number 91CH3054-4.

Czech:1990:GOO

[Cze90] M. Czech. GKS in an objectoriented environment. Computers and Graphics, 14(3-4):373– 375, 1990. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Davis:1988:DDP

[Dav88] Stephen Davis. Dypas: a dynamic payload scheduler for Shuttle missions. NASA conference publication, 3013:205– ??, 1988. CODEN NACPDX. ISSN 0191-7811.

Davies:1997:PWC

[DB97] Byron Davies and Victoria Bryan Davies. Patching onto the Web: Common Lisp hypermedia for the Intranet. Communications of the ACM, 40(5):66–69, May 1997. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/cacm/
                                        1997-40-5/p66-davies/.

Danieli:1988:POP

[DCSS88] Irit Danieli, Shimon Cohen, Raymond Sackler, and Beverly Sackler. Parallel objectoriented programming in SYM-PAL. In IEEE [IEE88i], pages 133–142. ISBN 0-8186-0884-6, 0-8186-4884-8 (microfiche). LCCN QA76.758 .I872 1988.

Diamant:1990:PBI

[DD90] John R. Diamant and Gerald P. Duggan. A preprocessor based implementation of Common Lisp. Lisp and Symbolic Computation, 3(3):211– 228, September 1990. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Demichiel:1988:OCL

[Dem88] Linda G. Demichiel. Overview: The Common Lisp Object System. Lisp and Symbolic Computation, 1(2):227–244, September 1988. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Du:1998:ASN

[DH98] Ding-Zhu Du and Frank K.

Hwang, editors. Advances in switching networks: DIMACS workshop, July 7–9, 1997, volume 42 of DIMACS series in discrete mathematics and theoretical computer science. American Mathematical Society, Providence, RI, USA, 1998. ISBN 0-8218-0831-1. LCCN TK5105.5 .A34 1998.

De:1995:KSS

[DKL95] Suranjan De, Chaiho Kim, and Anita Lee. Knowledge-based scheduling of semiconductor testing operations. In IEEE [IEE95b], pages 247–251. CO-DEN PCAAEF. ISBN 0-8186-7070-3 (paperback), 0-7803-2522-2 (casebound) 0-7803-2523-0 (microfiche). LCCN Q334 .C65 1995.

De:1993:FAS

[DL93] Suranjan De and Anita Lee.

Flexible assembly scheduling using a knowledge-based approach. Expert Systems with Applications, 6(3):309–326, July–September 1993. CODEN ESAPEH. ISSN 0957-4174.

Dattolo:1996:CVC

[DL96] Antonina Dattolo and Vincenzo Loia. Collaborative version control in an agent-based hypertext environment. Information systems, 21(2):127–145, April 1996. CODEN INSYD6. ISSN 0306-4379 (print), 1873-6076 (electronic).

De:1998:TKB

[DL98] S. De and A. Lee. Towards a knowledge-based scheduling system for semiconductor testing. International Journal of Production Research, 36(4): 1045–1073, April 1998. CO-DEN IJPRB8. ISSN 0020-7543.

Domazet:1992:CDP

[DLK92] D. S. Domazet, S. C.-Y. Lu, and M. Kalajdzic. Concurrent design and process planning of rotational parts. CIRP Annals, 41(1):181–184, 1992. CODEN CIRAAT. ISSN 0007-8506.

Dollin:1990:IAC

[Dol90] C. Dollin. Implementing Axis in Common Lisp. In Steels [Ste90b], pages 195–208. ISBN ???? LCCN ????

DEC:1986:VLV

[DS86] Digital Equipment Corporation and Guy L. Steele, Jr. VAX LISP/VMS. Digital Equipment Corporation, Maynard, MA, USA, 1986. various pp. Three volumes.

Dandekar:1997:UIS

[DZB97] Abhay Dandekar, Ibrahim Zeid, and Theodore Bardasz. User interface for specification language for case-based mechanical design. Artificial Intelligence for Engineering Design, Analysis and Manufacturing: AIEDAM, 11(1):17– 31, January 1997. CODEN AIEMEG. ISSN 0890-0604.

Evans:1993:WSC

[E^{+}93] Gerald W. Evans et al., editors. 1993 Winter Simulation Conference proceedings: The Biltmore Hotel, Los Angeles, California, December 12– 15, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN WSCPDK. ISBN 0-7803-1381-X (casebound). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1993. IEEE catalog no. 93CH3338-1.

Eccles:1988:PCL

[Ecc88] Joseph Eccles. Porting from Common Lisp with Flavors to C++. In USENIX Association [USE88], pages 31–40. ISBN ???? LCCN ????

Emerick:2012:CP

[ECG12] Chas Emerick, Brian Carper, and Christophe Grand. Clojure programming. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2012. ISBN 1-4493-9470-1. xviii + 607 pp. LCCN QA76.73.C565 E538

2012. URL http://www.loc.
gov/catdir/enhancements/
fy1211/2012405367-b.html;
 http://www.loc.gov/catdir/
enhancements/fy1211/2012405367-
d.html;   http://www.loc.
gov/catdir/enhancements/
fy1211/2012405367-t.html.

Ehn:1989:CIE

[Ehn89a] Ludovit Ehn. Contribution to the increase of efficiency of on-the-fly garbage collection. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 8(1):83–91, 1989. CO-DEN CARIDY. ISSN 0232-0274.

Ehn:1989:PAF

[Ehn89b] L’udovit Ehn. Performance analysis of on-the-fly garbage collection system. Computers and Artificial Intelligence = Vychislitel’nye mashiny i iskusstvennyi intellekt, 8(2): 141–152, 1989. CODEN CARIDY. ISSN 0232-0274.

Ewerbring:1989:SVC

[EL89] L. Magnus Ewerbring and Franklin T. Luk. Singular value computations on a massively parallel machine. In IEEE [IEE89g], pages 348–351. ISBN ???? LCCN TK7874 .T44 1989. IEEE catalog order no. A89CH2766-4.

Esakov:1990:KKR

[Esa90] Jeffrey Esakov. KB: a knowledge representation package for Common Lisp. Technical report MS-CIS-90-03, University of Pennsylvania, School of Engineering and Applied Science, Dept. of Computer and Information Science, Philadelphia, PA, USA, January 1990. 7 + 33 pp.

Ewing:1987:SIM

[EWB87] Juanita J. Ewing and Rebecca Wirfs-Brock. Smalltalk isn’t meaningless chatter. Computer Design, 26(2):76–79, January 1987. CODEN CMPDAM. ISSN 0010-4566.

Falk:1988:ATE

[Fal88] Howard Falk. AI techniques enter the realm of conventional languages. Computer Design, 27(19):45–49, October 1988. CODEN CMPDAM. ISSN 0010-4566.

Fateman:1990:LLM

[Fat90] Richard J. Fateman. A Lisplanguage Mathematica-to-Lisp translator. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 24(2):19–21, April 1990. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Fateman:1991:CRL

[Fat91] Richard J. Fateman. Canonical representations in Lisp and applications to computer algebra systems. In Watt [Wat91c], pages 360–

369. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

URL http://www.acm.org:
80/pubs/citations/proceedings/
issac/120694/p360-fateman/

.

Fateman:1997:NSS

[Fat97] Richard J. Fateman. Network servers for symbolic mathematics. In Küchlin [Küc97], pages 249–256. ISBN ???? LCCN ????

Ford:1993:OOM

[FBB93] Steve Ford, José A. Blakeley, and Thomas J. Bannon. Open OODB: a modular object-oriented DBMS. SIG-MOD Record (ACM Special Interest Group on Management of Data), 22(2):552–553, June 1993. CODEN SRECD8. ISBN 0-89791-592-5. ISSN 0163-5808 (print), 1943-5835 (electronic).

Funk:1987:AIA

[FBHJ87] T. L. Funk, G. F. Braun, D. Heidenrich, and J. E. Jones. Artificial intelligence approach to the selection of welding electrodes. In Anonymous [Ano87a], pages 149–150. ISBN 0-87171-277-6. LCCN ????

Fateman:1995:FFP

[FBWR95] Richard J. Fateman, Kevin A. Broughan, Diane K. Willcock, and Duane Rettig. Fast floating point processing in Common Lisp. ACM Transactions on Mathematical Software, 21(1):26–62, March 1995. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

http://www.acm.org/pubs/
citations/journals/toms/
1995-21-1/p26-fateman/.

Frydman:1989:DDB

[FGGB89] C. Frydman, N. Giambiasi, M. Gatumel, and P. Bayle.

DeBuMA: description, building and management of applications. Proceedings — Design Automation Conference, pages 203–208, 1989. CO-DEN PDAWDJ. ISSN 0146-7123. IEEE catalog number 89CH2734-2.

Fogus:2011:JC

[FH11] Michael Fogus and Chris Houser. The joy of Clojure. Manning, Greenwich, CT, USA, 2011. ISBN 1-935182-64-1 (paperback). xxxi + 328 pp. LCCN QA76.62 .F64 2011.

Freeman:1988:DSM

[FIKY88] Timothy S. Freeman, Gregory M. Imirzian, Erich Kaltofen, and Lakshman Yagati. Dagwood: a system for manipulating polynomials given by straight-line programs. ACM Transactions on Mathematical Software, 14(3): 218–240, September 1988. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Fateman:1991:SDP

[FL91] R. J. Fateman and D. T.

Lai. A simple display package for polynomials and rational functions in Common Lisp. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):1–3, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Flores:1999:CFR

[Flo99] Juan Flores. Complex fans: a representation for vectors in polar form with interval attributes. ACM Transactions on Mathematical Software, 25(2):129–156, June 1999. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        1999-25-2/p129-flores/;
                                        http://www.acm.org/pubs/
                                        citations/journals/toms/
                                        cgi-bin/TOMSbibget?Flores:
                                        1999:CFR.

Flug:1991:KEC

[Flu91]  Christian Flug.  Konzep-

tion eines effizienten Common Lisp-Systems für Universal-Prozessoren. (German) [Conception of an efficient Common Lisp system for universal processors]. Thesis (ph.d.), Technische Universität Carolo-Wilhelmina zu Braunschweig, Braunschweig, Germany, 1991. 234 pp.

Fujita:1981:IMI

[FN81] Yoneharu Fujita and Fujio Nishida. Improved method of immediate garbage collection. Systems — Computers — Controls, 12(2):39–44, March–April 1981. CODEN SYCCBB. ISSN 0096-8765.

Follin:1986:MPL

[Fol86] John F. Follin. Multirobot process line using artificial intelligence. In Anonymous [Ano86b], pages 4.87–4.101. ISBN 0-87263-223-7. LCCN TS191.8.R637 1986.

FranzInc:1988:ACL

[Fra88a] Franz Inc. Allegro COMMON LISP user guide: Release 3.1. Franz Inc., Berkeley, CA, USA, November 1988. various pp.

FranzInc:1988:CLR

[Fra88b] Franz Inc. COMMON LISP:

the reference. Addison-Wes-ley, Reading, MA, USA, 1988. ISBN 0-201-11458-5. xx + 899 pp. LCCN QA76.73.L23 C651 1988. US$29.95.

FranzInc:1994:ACU

[Fra94] Franz Inc. Allegro CL user guide. Berkeley, CA, USA, January 1994. ???? pp.

FranzInc:1996:ACU

[Fra96] Franz Inc. Allegro CL user guide. Franz Inc., Berkeley, CA, USA, version 4.3. edition, March 1996. ISBN ???? LCCN ????

Freeman:1990:PIC

[Fre90] Herbert Freeman, editor. Proceedings / 10th International Conference on Pattern Recognition, 16–21 June 1990, Atlantic City, New Jersey, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN PICREG. ISBN 0-8186-2062-5 (v. 1, paperback), 0-8186-2063-3 (v.2, paperback), 0-8186-6062-7 (v. 1, microfiche), 0-8186-6063-5 (v.2, microfiche), 0-8186-9062-3 (v.1, hardback), 0-8186-9063-1 (v.2, hardback). LCCN Q 327 I615 1990; TK7882.P3 I58 1990. Two volumes. IEEE catalog number 90CH2898-5. IEEE Computer Society order number 2062-2063.

Futrelle:1990:SDU

[Fut90] Robert P. Futrelle. Strategies for diagram understanding: Generalized equivalence, spatial/object pyramids and animate vision. In Freeman [Fre90], pages 403–408. CODEN PICREG. ISBN 0-8186-2062-5 (v. 1, paperback), 0-8186-2063-3 (v.2, paperback), 0-8186-6062-7 (v. 1, microfiche), 0-8186-6063-5 (v.2, microfiche), 0-8186-9062-3 (v.1, hardback), 0-8186-9063-1 (v.2, hardback). LCCN Q 327 I615 1990; TK7882.P3 I58 1990. Two volumes. IEEE catalog number 90CH2898-5. IEEE Computer Society order number 2062-2063.

Gupta:1986:CEP

[GA^{+}86] G. D. (Gopal D.) Gupta, K. S. Ahluwalia, et al., editors. Computers in engineering, 1986: proceedings of the 1986 ASME International Computers in Engineering Conference and Exhibition, July 20–24, 1986, Chicago, Illinois. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1986. CODEN COENEF. ISBN ???? LCCN TA 345 A86 1986. Three volumes.

Gadol:1988:SCL

[Gad88] S. Gadol. SPE — a Common Lisp environment on workstations. In Teitell [Tei88], pages 399–404. ISBN ???? LCCN Q335.A7868 1988.

Gat:1997:ELS

[Gat97]  Erann Gat.  ESL: a lan-

guage for supporting robust plan execution in embedded autonomous agents. In IEEE [IEE97b], pages 319–324. ISBN ???? ISSN 1095-323X. LCCN

????  URL http://iel.
ihs.com/ (Go to IEL; then
Aerospace); then Conferences;
then Search.

Goldman:1989:QPPa

[GG89a] Ron Goldman and Richard P.

Gabriel. Qlisp: parallel processing in Lisp. In Hoevel et al. [HMS^{+}89], pages 751–

760. CODEN PHISD7. ISBN 0-8186-1911-2 (paperback: vol. 1), 0-8186-5911-4 (microfiche: vol. 1), 0-8186-9911-6 (casebound: vol. 1), 0-8186-1912-0 (paperback: vol. 2), 0-8186-5912-2 (microfiche: vol. 2), 0-8186-9912-4 (casebound: vol. 2), 0-8186-1913-9 (paperback: vol. 3), 0-8186-1914-7 (paperback: vol. 4), 0-8186-5914-9 (microfiche: vol. 4), 0-8186-9914-0 (casebound: vol. 4). ISSN 0073-1129. LCCN ???? Computer Society order number 1911. IEEE catalog number 89TH0242-8.

Goldman:1989:QPPb

[GG89b] Ron Goldman and Richard P. Gabriel. Qlisp: Parallel processing in Lisp. IEEE Software, 6(4):51–59, July 1989. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Groschwitz:1994:EST

[GH94] Hans Groschwitz and Joachim Holtz. Expert system on transformer design for switched mode power supplies. In IEEE [IEE94b], pages 1413– 1418. CODEN IEPREA. ISBN 0-7803-1329-1 (casebound), 0-7803-1330-5 (microfiche), 0-7803-1328-3 (softbound). LCCN TK7881 .I55 1994. Three volumes. IEEE catalog number 94CH3319-1.

Goblick:1990:MCS

[GHL^{+}90] T. J. Goblick, P. G. Harmon, J. I. Leivent, J. J. Olsen, D. S. Cogen, J. J. DiTuri, P. G. McHugh, and R. L. Walton. A multiprocessor computer system for integrated numeric and symbolic computation. In Meystel et al. [MHG90], pages 336–341. ISBN 0-8186-2108-7 (paperback), 0-8186-9108-5 (casebound), 0-8186-6108-9 (microfiche). LCCN TJ 217.5 I33 1990a. Two volumes. IEEE Computer Society Press Order Number 2108. IEEE Catalog Number 90TH0333-5.

Gilmore:1986:AAI

[Gil86] John F. Gilmore, editor. Applications of artificial intelligence IV: 15-16 April 1986, Innsbruck, Austria, volume 657 of SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1986. CO-DEN PSISDG. ISBN 0-89252-692-0. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510 .P632 v. 657.

Giunchiglia:1987:KCL

[Giu87] Fausto Giunchiglia. Kyoto Common Lisp: a high level language implementation of Common Lisp on general purpose machines. In IEEE [IEE87b], pages 124–127. CODEN DC-SIDU. ISBN 0-8186-0764-5. LCCN QA75.5.C58 1987.

Goldberg:1990:MCL

[GK90] David E. (David Edward)

Goldberg and Travis Kerzic. mGA1.0: a Common LISP implementation of a messy genetic algorithm. NASA contractor report NASA CR-187260, Research Institute for Computing and Information Systems University of Houston, Clear Lake, Houston, TX, USA, May 1990. 47 pp.

Gokhale:1990:LDG

[GKLM90] Maya B. Gokhale, Andrew Kopser, Sara P. Lucas, and Ronald G. Minnich. The logic description generator. In Kung et al. [K^{+}90], pages 111–120.

ISBN ???? LCCN TK7874.6 .I57 1990.

Galperin:1989:KBS

[GKS89] A. Galperin, S. Kimhi, and M. Segev. Knowledge-based system for optimization of fuel reload configurations. Nuclear Science and Engineering, 102 (1):43–53, May 1989. CODEN NSENAO. ISSN 0029-5639.

Gisolfi:1994:DCS

[GL94] Antonio Gisolfi and Vincenzo Loia. Designing complex systems within distributed architectures: An intelligent tutoring systems perspective. Applied Artificial Intelligence, 8 (3):393–411, July–September 1994. CODEN AAINEH. ISSN 0883-9514.

Glicksman:1986:SEA

[Gli86] Jay Glicksman. Simulator environment for an autonomous land vehicle. Simulation series, 17(1):53–57, January 1986. CODEN SISEDL. ISSN 0735-9276.

Gloess:1986:OFF

[GM86] Paul Y. Gloess and Jacques Marcovich. Oblogis, a flexible flavor implementation of Prolog logic and its application to the design of a broaching expert system. In Sriram and Adey [SA86], page various. ISBN 0-905451-47-3. LCCN TA174.I48 1986. Two volumes.

Goto:1991:GTS

[Got91] Kimio Goto. Generation of the three stages NAND gate circuits with the single rail inputs by using the inhibiting loop method. In IEEE [IEE91b], pages 2994–2998. CODEN PICSDI. ISBN 0-7803-0051-3 (casebound), 0-7803-0050-5 (softbound), 0-7803-0052-1 (microfiche). ISSN 0271-4310. LCCN TK7801 .I245 1991. Five volumes. IEEE catalog number 91CH3006-4.

Gat:1997:ARM

[GP97] Erann Gat and Barney Pell.

Abstract resource management in an unconstrained plan execution system. In IEEE [IEE97b], pages 383–389. ISBN ???? ISSN 1095-323X. LCCN

????  URL http://iel.
ihs.com/ (Go to IEL; then
 Aerospace); then Conferences;
then Search.

Graham:1994:LAT

[Gra94] Paul Graham. On Lisp: advanced techniques for Common Lisp. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1994. ISBN 0-13-030552-9. xiii + 413 pp. LCCN QA76.73.C28 G73 1994.

Graham:1996:ACL

[Gra96] Paul Graham. ANSI Common Lisp. Prentice Hall series in artificial intelligence. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1996. ISBN 0-13-370875-6 (paperback). xiii + 432 pp. LCCN QA76.73.C28 G69 1996. Also available in a German translation [Gra97].

Graham:1997:ACL

[Gra97] Paul Graham. ANSI Common Lisp. Prentice-Hall, Haar bei München, Germany, 1997. ISBN 3-8272-9543-2. 510 pp. LCCN ???? DM 69,95; S 511,00. German translation of [Gra96].

Graham:1999:ACL

[Gra99] Paul Graham. ANSI Common LISP. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, second edition, 1999. ISBN 0-13-096138-8. 432 pp. LCCN ????

Greenwood:1984:ADA

[Gre84] Leonard J. Greenwood. Apollo Domain: an architectural review. In IEE Colloquium (Digest), volume 1984/80, pages 8.1–8.4. IEE, London, UK, 1984. CODEN DCILDN. ISBN ???? LCCN ????

Guise:1988:LRP

[Gui88] Dario Guise. LISP as a rapid prototyping environment. the Chinese tutor. Lisp and Symbolic Computation, 1(2):165– 184, September 1988. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Gertner:1998:TOD

[GW98] Abigail S. Gertner and Bonnie L. Webber. TraumaTIQ: Online decision support for trauma management. IEEE expert: intelligent systems and their applications, 13(1):32–39, January–February 1998. CO-DEN IEEXE7. ISSN 0885-9000.

Gabriel:1991:CIO

[GWB91] Richard P. Gabriel, Jon L.

White, and Daniel G. Bobrow. CLOS: integrating object-oriented and functional programming. Communications of the ACM, 34(9):28– 38, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0001-
0782/114671.html.

Hadden:1986:MES

[Had86] George D. Hadden. Mentor, an expert system for preventive maintenance. Advances in Instrumentation, 41(pt 2):797– 807, 1986. CODEN AVINBP. ISSN 0065-2814.

Hadingham:1987:AIH

[Had87] Evan Hadingham. Artificial intelligence in the hospital environment. Physicians & Computers, 5(4):8–11, August 1987. CODEN PHCOEY.

Halstead:1990:DAE

[Hal90] Rodd Halstead. Develop advanced expert systems. Byte Magazine, 15(1), January 1990. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Hall:1992:CPP

[Hal92] Robert J. Hall. Call path profiling. Proceedings — International Conference on Software Engineering, 14:296–306, May 1992. CODEN PCSEDE. ISSN 0270-5257.

Halloway:2009:PC

[Hal09] Stuart Dabbs Halloway. Programming Clojure. The pragmatic programmer. Pragmatic Bookshelf, Raleigh, NC, USA, 2009. ISBN 1-934356-33-6 (paperback). xxi + 280 pp. LCCN QA76.62 .H35 2009.

Harris:1990:PBN

[Har90a] Nomi L. Harris. Probabilistic belief networks for genetic counseling. Computer Methods and Programs in Biomedicine, 32(1):37–44, May 1990. CO-DEN CMPBEK. ISSN 0169-2607 (print), 1872-7565 (electronic).

Harrison:1990:CLA

[Har90b] Patrick R. Harrison. Common Lisp and artificial intelligence. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1990. ISBN 0-13-155243-0. xi + 244 pp. LCCN QA 76.73 C28 H36 1990. US$22.50.

Hohl:1996:HAH

[HBG96] Hubertus Hohl, Heinz-Dieter Boecker, and Rul Gunzenhaeuser. Hypadapter: An adaptive hypertext system for exploratory learning and programming. User Modelling and User-Adapted Interaction, 6(2-3):131–156, July 1996. CODEN UMUIEQ.

Harten:1985:SLD

[HC85] Leo P. Harten and George J.

Carrette. Share library in DOE-MACSYMA. In Anonymous [Ano85a], pages 99–102. ISBN 0-87664-872-3. LCCN ????

Hasemer:1989:CLP

[HD89] Tony Hasemer and John Domingue. Common LISP programming for artificial intelligence. International computer science series. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17579-7 (paperback). xiii + 444 pp. LCCN QA76.73.C28 H37 1989. UK£15.95.

Hennessey:1989:CL

[Hen89] Wade L. Hennessey. COM-MON LISP. McGraw-Hill, New York, NY, USA, 1989. ISBN 0-07-028177-7. xv + 395 pp. LCCN QA76.73.L23H48 1989.

Henglein:1992:GTO

[Hen92a] Fritz Henglein. Global tagging optimization by type inference. In ACM [ACM92], pages 205–215. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Hennessey:1992:WDE

[Hen92b] Wade Hennessey. WCL: Delivering efficient Common Lisp applications under Unix. In ACM [ACM92], pages 260– 269. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Herczeg:1993:CAB

[Her93] J. Herczeg. A constructionkit approach to building tracers for the Common Lisp Object System. In Nierstrasz [Nie93], pages 55–66. ISBN 3-540-57120-5, 0-387-57120-5. LCCN QA76.64 .E95 1993.

Hidde:1991:AMD

[HG91] Axel R. Hidde and Andreas Gierse. An AI-based manufacturing design rule checker and path optimizer for PCB production preparation and manufacturing. In IEEE/CHMT European International Electronic Manufacturing Technology Symposium, pages 244– 249. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, January 1991. CODEN PIESET. ISBN ???? LCCN ????

Hewett:1993:LAE

[HH93] Micheal Hewett and Rattikorn Hewett. Language and architecture for efficient blackboard systems. In IEEE [IEE93c], pages 34–40. CO-DEN PCAAEF. ISBN 0-8186-3840-0 (paperback), 0-8186-3842-7 (casebound), 0-8186-3841-9 (microfiche). LCCN Q334 .C662 1993. IEEE catalog number 93CH3454-0.

Hammad:1994:BMI

[HI94] Amin Hammad and Yoshito Itoh. BRICAL; a multimedia intelligent educational system for bridge engineering. Computing in Civil Engineering (New York), 2:1671–1678, 1994. CODEN CCENEX.

Hwang:1986:PIC

[HJS86] Kai Hwang, Steven M. Jacobs, and Earl E. Swartzlander, editors. Proceedings of the 1986 International Conference on Parallel Processing, August 19–22, 1986. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN PCPADL. ISBN 0-8186-0724-6. ISSN 0190-3918. LCCN QA76.6 .I581 1986.

Heragu:1990:MLO

[HK90] Sunderesh S. Heragu and Andrew Kusiak. Machine layout: an optimization and knowledge-based approach. International Journal of Production Research, 28(4):615–635, April 1990. CODEN IJPRB8. ISSN 0020-7543.

Heger:1991:KAD

[HK91] A. Sharif Heger and Billy V.

Koen. KNOWBOT. an adaptive data base interface. Nuclear Science and Engineering, 107(2):142–157, February 1991. CODEN NSENAO. ISSN 0029-

Hesse:1993:PIB

[HKLH93] Jan Hesse, Rainer König, Filippo Logi, and Jens Herder. A prototype of an interface builder for the Common Lisp Interface Manager — CLIB. ACM SIGPLAN Notices, 28 (8):19–28, August 1993. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Hoevel:1989:PTS

[HMS^{+}89] Lee W. Hoevel, Veljko Milutinovic, Bruce D. Shriver, Robert Blanning, David King, and Ralph H. Sprague, Jr., editors. Proceedings of the Twentysecond annual Hawaii International Conference on System Sciences, January 3–6, 1989, Kailua Kona, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN PHISD7. ISBN 0-8186-1911-2 (paperback: vol. 1), 0-8186-5911-4 (microfiche: vol. 1), 0-8186-9911-6 (casebound: vol. 1), 0-8186-1912-0 (paperback: vol. 2), 0-8186-5912-2 (microfiche: vol. 2), 0-8186-9912-4 (casebound: vol. 2), 0-8186-1913-9 (paperback: vol. 3), 0-8186-1914-7 (paperback: vol. 4), 0-8186-5914-9 (microfiche: vol. 4), 0-8186-9914-0 (casebound: vol. 4). ISSN 0073-1129. LCCN ???? Computer Society order number 1911. IEEE catalog number 89TH0242-8.

Hoevel:1990:PTT

[HMS^{+}90] Lee W. Hoevel, Veljko Milutinovic, Bruce D. Shriver, Jay F. Nunamaker, Jr., and Ralph H. Sprague, Jr., editors. Proceedings of the Twenty-Third Annual Hawaii International Conference on System Sciences. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN PHISD7. ISBN 0-8186-2008-0 (paperback: vol.

1) 0-8186-2009-9 (paperback: vol. 2) 0-8186-2010-2 (paperback: vol. 3) 0-8186-2011-0 (paperback: vol. 4) 0-8186-9008-9 (hard: vol. 1) 0-8186-9009-7 (hard: vol. 2) 0-8186-9010-0 (hard: vol. 3) 0-8186-9011-9 (hard: vol. 4) 0-8186-6008-2 (microfiche: vol. 1) 0-8186-6009-0 (microfiche: vol. 2) 0-8186-6010-4 (microfiche: vol.

3) 0-8186-6011-2 (microfiche: vol. 4). ISSN 0073-1129. LCCN QA76.9.S88H38 1990. Four volumes. Computer Society order no.: 2008 (vol. 1), 2009 (vol. 2), 2010 (vol. 3), 2011 (vol. 4); IEEE catalog no.: 90TH0293-1 (vol. 1), 90TH0294-9 (vol. 2), 90TH0295- 6 (vol. 3), 90TH0296-4 (vol. 4).

Hollnagel:1992:DFT

[Hol92] Erik Hollnagel. Design of fault tolerant systems. prevention is better than cure. Reliability Engineering & System Safety, 36(3):231–237, 1992. CODEN RESSEP. ISSN 0951-8320.

Holden:2014:BYO

[Hol14] Daniel Holden. Build your own lisp: learn C and build your own programming language. ????, ????, 2014. ISBN 1-5010-0662-2. 210 pp. LCCN QA76.73.L23 H65 2014.

Habert:1996:OCC

[HP96] Benoit Habert and Jean-François Perrot. Objectif CLOS (Common Lisp Object System): en 10 leçons. Collection objectif; 1240-1013. Masson Editeur, Masson, France, 1996. ISBN 2-225-85120-4. 217 pp. LCCN ????

Hart:1989:HKE

[HR89] P. K. Hart and J. Rodriguez. Hybrid knowledge-based expert system for design of structural elements. In Riley et al. [RCG^{+}89], pages 287–292. CO-DEN COENEF. ISBN 0-7918-0357-0 (vol. 1) 0-7918-0358-9 (vol. 2). LCCN TA 345 I5488 1989. Two volumes.

Hoevel:1988:PTF

[HSKS88] Lee W. Hoevel, Bruce D. Shriver, Benn R. Konsynski, and Ralph H. Sprague, Jr., editors. Proceedings of the Twenty-First Annual Hawaii International Conference on System Sciences, Kailua Kona, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CODEN PHISD7. ISBN 0-8186-0841-2 (paperback: vol. 1), 0-8186-8841-6 (hardcover: vol. 1), 0-8186-4841-4 (microfiche: vol. 1). ISSN 0073-1129. LCCN QA76.9.S88H375 1988. Four volumes. Computer Society order number 841. IEEE catalog number 88TH0209-7.

Horn:1985:GCL

[HSW85] Berthold Horn, Guy L. Steele, Jr., and Patrick Henry Winston. Golden Common LISP. Gold Hill Computers, Cambridge, MA, USA, version 1.01. edition, 1985. ISBN 0-917589-05-X. LCCN ???? 6 program files on 1 computer disk data files on 4 computer disks + 1 manual + 1 reference manual + 1 text.

Horn:1987:GCL

[HSW87] Berthold Horn, Guy L. Steele, Jr., and Patrick Henry Winston. Golden Common LISP. Gold Hill Computers, Cambridge, MA, USA, version 1.10. edition, 1987. ISBN 0-201-08372-8. Program files on 10 computer disks manual, plus one text.

Hummel:1994:CLI

[Hum94] T. Hummel. A Common Lisp interface for dynamic patching with the IRCAM signal processing workstation. In Anonymous [Ano94a], pages 216–219. ISBN ???? LCCN ???? Also published as Stanford University, Center for Computer Research in Music and Acoustics, Report no. STAN-M-89.

Hurley:1991:OGA

[Hur91] Catherine Hurley. Objectoriented graphical analysis. In Nelson et al. [NKC91], pages 877–881. CODEN WSCPDK. ISBN 0-7803-0181-X (??invalid checksum??). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1991. IEEE catalog no. 91CH3050-2.

Herlihy:1987:ALS

[HW87] Maurice P. Herlihy and Jeannette M. Wing. Avalon: Language support for reliable distributed systems. In IEEE [IEE87d], pages 89–94. CO-DEN DPFTDL. ISBN 0-8186-0778-5. ISSN 0731-3071. LCCN QA 76.9 F38 I57 1987.

Hug:1995:UIH

[HWK95] Berthold Hug, Gerhard J. Walz, and Markus Kuhl. User interaction in HP PE/ SolidDesigner. Hewlett-Pack-ard Journal: technical information from the laboratories of Hewlett-Packard Company, 46 (5):14–23, October 1995. CO-DEN HPJOAX. ISSN 0018-

                                        1153. URL http://www.hp.
                                        com/hpj/toc-10-95.html.

IPS:1994:DPA

[IAC94] Information Processing Systems, American National Standards Institute. Committee X3, and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard Programming Language Common LISP. X3 Secretariat/CBEMA (Reproduced by Global Engineering Documents), Washington, DC, USA, January 10, 1994.

Ida:1990:WWD

[Ida90] M. Ida. What we have done with Common Lisp in Japan during 1984–1989. In Steels [Ste90b], pages 49–56. ISBN ???? LCCN ????

IEEE:1984:CPT

[IEE84a] IEEE, editor. 1984 conference proceedings / third annual International Phoenix Conference on Computers and Communications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN CPACE3. ISBN 0-8186-0555-3. LCCN TK 7885 A1 I567 1984.

IEEE:1984:ILD

[IEE84b]  IEEE, editor.  Intellectual

Leverage the Driving Technologies: Digest of Papers, Compcon Spring 84, February 27–March 1, Twenty-eighth IEEE Computer Society International Conference, Meridien Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN PCI-CDQ. ISBN 0-8186-0525-1 (paperback), 0-8186-8525-5 (hardcover). LCCN QA75.5 .C58 1984, TK7885.A1 C53 1984. IEEE catalog no. 84CH2017-2.

IEEE:1985:PIC

[IEE85a]  IEEE, editor.  Proceedings

/ 1st International Conference on Computer Workstations, November 11–14, 1985, San Jose, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0649-5 (paperback), 0-8186-8649-9 (hardcover), 0-8186-4649-7 (microfiche). LCCN QA76.5 .I544 1985.

IEEE:1985:PIN

[IEE85b]  IEEE, editor.  Proceedings

of the IEEE 1985 National Aerospace and Electronics Conference, NAECON 1985: held at the Dayton Convention Center, May 20–24, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. CODEN NASEA9. ISBN ???? LCCN TL 693 N37 1985. Two volumes. IEEE catalog no. 86CH2189-9.

IEEE:1985:SCA

[IEE85c] IEEE, editor. The Second Conference on Artificial Intelligence Applications: the engineering of knowledge-based systems / sponsored by IEE Computer Society, Fontainebleau Hilton, Miami Beach, Florida, December 11–13, 1985. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0688-6, 0-8186-8688-X (hardcover), 0-8186-4688-8 (microfiche). LCCN Q 334 C66 1985.

IEEE:1986:ATC

[IEE86a] IEEE, editor. Advancing test concepts and the user: proceedings / Autotestcon 86, San Antonio, TX, September 8– 11, 1986, IEEE International Automatic Testing Conference, Henry B. Gonzalez Convention Center, San Antonio, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN AUPRDX. ISBN ???? ISSN 0734-7510. LCCN TK 7895 A8 A98 1986.

IEEE:1986:CPA

[IEE86b] IEEE, editor. Conference proceedings / the 13th annual International Symposium on Computer Architecture, June 2–5, 1986, Tokyo, Japan. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN CPAADU. ISBN 0-8186-0719-X (pbk.), 0-8186-8719-3. ISSN 0149-7111. LCCN QA 76.9 A73 I56 1986.

IEEE:1986:ICS

[IEE86c] IEEE, editor. IEEE Computer Society Second International Conference on Ada Applications and Environments: April 8–10, 1986, Konover Hotel, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0718-1 (paperback), 0-8186-8718-5 (hard), 0-8186-4718-3 (microfiche). LCCN QA76.73.A35 I33 1986. IEEE catalog number 86CH2281-4. IEEE Computer Society order number 718.

IEEE:1986:PII

[IEE86d] IEEE, editor. Proceedings / 1986 IEEE International Conference on Robotics and Automation, April 7–10, 1986, the San Francisco Hilton and Tower, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-0695-9. LCCN TJ210.3 .I44 1986. Three volumes. IEEE Computer Society order number 695.

IEEE:1986:PIA

[IEE86e] IEEE, editor. Proceedings / IEEE/AIAA 7th Digital Avionics Systems Conference, October 13–16, 1986, Worthington Hotel, Fort Worth, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN ???? LCCN TL693.D55 1986. IEEE catalog number: 86CH2359-8.

IEEE:1987:IIS

[IEE87a] IEEE, editor. 1987 IEEE International Symposium on Circuits and Systems, Dunfey City Line Hotel, Philadelphia, PA, May 4–7, 1987. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN PICSDI. ISBN ???? LCCN TK 7801 I22 1987. Three volumes.

IEEE:1987:DPI

[IEE87b] IEEE, editor. Digest of papers:

intellectual leverage / Compcon spring 87, February 23–27, Thirty-Second IEEE Computer Society International Conference, Cathedral Hill Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN DCSIDU. ISBN 0-8186-0764-5. LCCN QA75.5.C58 1987.

IEEE:1987:DPT

[IEE87c] IEEE, editor. Digest of papers:

thirty-second IEEE Computer Society international conference, Cathedral Hill Hotel, San Francisco, California, February 23–27, Compcon, spring 87. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN DC-SIDU. ISBN 0-8186-0764-5 (paperback), 0-8186-8764-9 (casebound), 0-8186-4764-7 (microfiche). LCCN QA76 .I49 1987. IEEE catalog number 87CH2409-1. Computer Society Order number 764.

IEEE:1987:FDP

[IEE87d] IEEE, editor. FTCS 17: digest of papers / the Seventeenth International Symposium on Fault-Tolerant Computing, July 6–8, 1987, Sheraton Station Square, Pittsburgh, Pennsylvania, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CODEN DPFTDL. ISBN 0-8186-0778-5. ISSN 0731-3071. LCCN QA 76.9 F38 I57 1987.

IEEE:1987:PII

[IEE87e] IEEE, editor. Proceedings / 1987 IEEE International Conference on Computer Design, VLSI in Computers & Processors, ICCD ’87, Rye Town Hilton, Rye Brook, New York, October 5–October 8, 1987. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0802-1 (paperback), 0-8186-4802-3 (microfiche), 0-8186-8802-5 (casebound). LCCN TK 7888.4 I23 1987. IEEE catalog number 87CH2473-7. IEEE Computer Society order number 802.

IEEE:1987:PAC

[IEE87f]  IEEE, editor.  Proceedings

of the 1987 American Control Conference: Hyatt Regency Hotel, Minneapolis, MN, June 10–12, 1987. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. CO-DEN PRACEO. ISBN ???? ISSN 0743-1619. LCCN TJ 213 A43 1987. Three volumes. IEEE Catalog Number: 87CH2378-8.

IEEE:1988:ICC

[IEE88a] IEEE, editor. 1988 International Conference on Computer Languages: proceedings, October 9–13, 1988, Castle Hotel and Resort, Miami Beach, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0874-9. LCCN QA76.7 .I561 1988. IEEE catalog number 88CH2647-6. Computer Society order number 874.

IEEE:1988:AID

[IEE88b] IEEE, editor. 25th ACM/ IEEE Design Automation Conference: 25th DAC, June 12– June 15, 1988, Anaheim Convention Center: proceedings 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CODEN PDAWDJ. ISBN 0-8186-8864-5. ISSN 0146-7123. LCCN TA 174 D46p 1988. IEEE catalog number 88CH2540-3.

IEEE:1988:OPP

[IEE88c] IEEE, editor. Oceans 88:

proceedings, a partnership of marine interests / conference sponsored by Marine Technology Society, IEEE: Baltimore, Maryland, October 31–November 2, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988.

ISBN ???? LCCN GC2 .O32 1988. Four volumes. IEEE catalog number 88-CH2585-8.

IEEE:1988:PIIa

[IEE88d]  IEEE, editor.  Proceedings

— IEEE International Symposium on Circuits and Systems, June 7–9, Espoo, Finland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CODEN PICSDI. ISBN 951-721-240-2. ISSN 0271-4310. LCCN ???? Available from IEEE Service Cent (catalog number 88CH2458-8). Piscataway, NJ, USA.

IEEE:1988:PFC

[IEE88e] IEEE, editor. Proceedings / the Fourth Conference on Artificial Intelligence Applications. Sheraton Harbor Island Hotel, San Diego, California, March 14– 18, 1988. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0837-4. LCCN Q 334 C66 1988. IEEE catalog number 88CH2552-8.

IEEE:1988:PWF

[IEE88f]  IEEE, editor.  Proceedings

/ Workshop on the Future Trends of Distributed Computing Systems in the 1990s, 14– 16 September 1988, Royal Garden Hotel, Hong Kong. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0897-8. LCCN QA76.9.D5W67 1988. IEEE catalog number 88TH0228-7.

IEEE:1988:PIIb

[IEE88g]  IEEE, editor.  Proceedings

1988 IEEE International Conference on Robotics and Automation, April 24–29, 1988, Franklin Plaza Hotel, Philadelphia, Pennsylvania. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN ???? LCCN ???? Three volumes. IEEE catalog number 88CH2555-1. Computer Society order number 852.

IEEE:1988:PHI

[IEE88h]  IEEE, editor.  Proceedings

of the Twenty-First Annual Hawaii International Conference on System Science (Kailua-Kona, HI, USA). Volume II: Software Track. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. CODEN PHISD7. ISBN 0-8186-0842-0. ISSN 0073-1129. LCCN ????

IEEE:1988:TIC

[IEE88i] IEEE, editor. The Third Israel Conference on Computer Systems and Software Engineering: proceedings: June 6– 7, 1988, Hilton Hotel, Tel-Aviv. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0884-6, 0-8186-4884-8 (microfiche). LCCN QA76.758 .I872 1988.

IEEE:1989:IIC

[IEE89a] IEEE, editor. 1989 IEEE International Conference on Systems, Man, and Cybernetics: conference proceedings, November 14–17, 1989, Hyatt Regency, Cambridge, Cambridge, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN PICYE3. ISBN ???? ISSN 0884-3627. LCCN Q300 .I485 1989. Three volumes. IEEE catalog number 89CH2809-2.

IEEE:1989:IPR

[IEE89b] IEEE, editor. IEEE Pacific Rim Conference on Communications, Computers and Signal Processing, Victoria, BC, Canada, June 1st–2nd, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN ???? LCCN TK5102.5 .I4 1989. IEEE Cat No. 89CH2691-4.

IEEE:1989:PIC

[IEE89c] IEEE, editor. Proceedings: 11th International Conference on Software Engineering, Nice, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN PCSEDE. ISBN 0-89791-258-6, 0-8186-1941-4 (paperback), 0-8186-5941-6 (microfiche). ISSN 0270-5257. LCCN QA 76.6 I553 1989.

IEEE:1989:PCJ

[IEE89d] IEEE, editor. Proceedings, CVPR ’89: June 4–8, 1989, San Diego, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1952-X (paperback), 0-8186-8952-8 (hardcover), 0-8186-5952-1 (microfiche). LCCN TA1632 .I36 1989.

IEEE:1989:PAC

[IEE89e]  IEEE, editor.  Proceedings

of the 1989 American Control Conference: Pittsburgh Hilton Hotel and Towers, Pittsburgh, PA, June 21–23, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN ???? LCCN TJ 212.2 A44 1989. Three volumes. IEEE catalog number: 89CH2772-2.

IEEE:1989:PAI

[IEE89f] IEEE, editor. Proceedings:

the 16th Annual International Symposium on Computer Architecture, Jerusalem, 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN CPAADU. ISBN 0-8186-8948-X (casebound), 0-8186-1948-1 (paperback). ISSN 0149-7111. LCCN QA 76.9 A73 C65 v.17 no.3 1989. IEEE Computer Society order number 1948. IEEE catalog number 89CH2705-2. ACM order number 415890.

IEEE:1989:TFI

[IEE89g] IEEE, editor. TENCON ’89: fourth IEEE Region 10 International Conference: “Information technologies for the 90’s,” EC, energy, electrons, computers, communications: Oberoi Towers, Bombay India, 22, 23 & 24 November 1989. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN ???? LCCN TK7874 .T44 1989. IEEE catalog order no. A89CH2766-4.

IEEE:1990:IIS

[IEE90a] IEEE, editor. 1990 IEEE International Symposium on Circuits and Systems, Sheraton Hotel, New Orleans, LA, May 1–3, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN PICSDI. ISBN ???? ISSN 0271-4310. LCCN TK 7801 I22 1990. Four volumes.

IEEE:1990:ICC

[IEE90b] IEEE, editor. 1990 International Conference on Computer Languages, New Orleans, LA, USA, March 12–15, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2036-6. LCCN QA76.7 .I576 1990. IEEE Computer Society Press order number 2036. IEEE catalog number 90CH2854-8.

IEEE:1990:CPT

[IEE90c] IEEE, editor. The changing philosophy of test: International Test Conference, 1990 proceedings: September 10– 14, 1990, Sheraton Washington Hotel, Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN DITCDP. ISBN 0-8186-2064-1 (paperback), 0-8186-9064-X (hardcover), 0-8186-6064-3 (microfiche). ISSN 0743-1686. LCCN TK 7874 I593 1990. IEEE catalog number 90CH2910-6.

IEEE:1990:ITC

[IEE90d] IEEE, editor. IEEE TEN-CON’90: conference proceedings / 1990 IEEE Region 10 Conference on Computer and Communication Systems, 24– 27 September, 1990, Hong Kong Convention and Exhibition Centre, Hong Kong. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-87942-557-1, 0-87942-556-3 (paperback), 0-87942-558-X (microfiche). LCCN TK 5105.5 I34 1990. Two volumes. IEEE catalog number 90CH2866-2.

IEEE:1990:PSN

[IEE90e] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0. LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

IEEE:1991:IIC

[IEE91a] IEEE, editor. 1991 IEEE International Conference on Systems, Man, and Cybernetics: decision aiding for complex systems: October 13– 16, 1991, Omni Charlottesville Hotel and the University of Virginia, Charlottesville, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN PICYE3. ISBN 0-7803-0234-6 (set: casebound), 0-7803-0233-8 (set: softbound), 0-7803-0235-4 (microfiche). ISSN 0884-3627. LCCN TA 168 I19 1991; Q300 .I485 1991. Three volumes. IEEE catalog number 91CH3067-6.

IEEE:1991:IIS

[IEE91b] IEEE, editor. 1991 IEEE International Symposium on Circuits and Systems: the Westin Stamford and Westin Plaza, Singapore, 11–14 June 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN PICSDI. ISBN 0-7803-0051-3 (casebound), 0-7803-0050-5 (softbound), 0-7803-0052-1 (microfiche). ISSN 0271-4310. LCCN TK7801 .I245 1991. Five volumes. IEEE catalog number 91CH3006-4.

IEEE:1991:PIN

[IEE91c]  IEEE, editor.  Proceedings

of the IEEE 1991 National Aerospace and Electronics Conference, NAECON 1991, held at the Dayton Convention Center, May 20–24, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN NASEA9. ISBN 0-7803-0085-8. LCCN TL 693 N37 1991. Three volumes.

IEEE:1991:TIC

[IEE91d] IEEE, editor. Third International Conference on Tools for Artificial Intelligence: TAI ’91, November 5–8, 1991, San Jose, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2300-4 (paperback), 0-8186-2302-0 (casebound), 0-8186-2301-2 (microfiche). LCCN Q334 .I563 1991. IEEE Computer Society Press Number 2300. IEEE Catalog Number 91CH3054-4.

IEEE:1992:FIC

[IEE92] IEEE, editor. Fourth International Conference on Software Engineering and Knowledge Engineering: proceedings (SE KE 92, June 15–20, 1992, Europa Palace Hotel, Capri, Italy). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2830-8. LCCN QA76.758.I573 1992.

IEEE:1993:CPS

[IEE93a]  IEEE, editor.  COMPSAC

93: Proceedings of the Seventeenth Annual International Computer Software & Applications Conference, November 1–5, 1993, Phoenix, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN PSICD2. ISBN 0-8186-4440-0 (paperback), 0-8186-4442-7 (casebound), 0-8186-4441-9 (microfiche). ISSN 0730-6512. LCCN QA76.6.C6295 1993. IEEE catalog number 93CH3349-8.

IEEE:1993:FIC

[IEE93b] IEEE, editor. Fifth International Conference on Tools with Artificial Intelligence: TAI ’93: proceedings, November 8–11, 1993, Boston, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4202-5 (casebound), 0-8186-4200-9 (paperback). ISSN 1063-6730. LCCN Q334 .I5634 1993. IEEE Catalog Number 93CH3325-8. IEEE Computer Society Press Order Number 4200-02.

IEEE:1993:NCA

[IEE93c] IEEE, editor. The Ninth Conference on Artificial Intelligence for Applications: proceedings / March 1–5, 1993, Orlando, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. CODEN PCAAEF. ISBN 0-8186-3840-0 (paperback), 0-8186-3842-7 (casebound), 0-8186-3841-9 (microfiche). LCCN Q334 .C662 1993. IEEE catalog number 93CH3454-0.

IEEE:1993:SII

[IEE93d] IEEE, editor. Second IEEE International Conference on Fuzzy Systems: San Francisco, California, March 28–April 1, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-7803-0616-3 (microfiche), 0-7803-0615-5 (casebound), 0-7803-0614-7 (softbound). LCCN TJ212.2.I3249 1993. Two volumes. IEEE catalog number: 93CH3136-9.

IEEE:1994:CIC

[IEE94a] IEEE, editor. CVPR ’94:

1994 IEEE Computer Society Conference on Computer Vision and Pattern Recognition: proceedings, June 21– 23, 1994, Seattle, Washington. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5827-4 (casebound), 0-8186-5825-8 (paperback), 0-8186-5826-6 (microfiche). ISSN 1063-6919. LCCN TA1632 .I36 1994. IEEE Computer Society Press order number 5825-02. IEEE catalog number 94CH3405-8.

IEEE:1994:IIC

[IEE94b] IEEE, editor. IECON ’94, 20th International Conference on Industrial Electronics, Control, and Instrumentation, Sept. 5– 9, 1994, Universita di Bologna, Italy. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. CO-DEN IEPREA. ISBN 0-7803-1329-1 (casebound), 0-7803-1330-5 (microfiche), 0-7803-1328-3 (softbound). LCCN TK7881 .I55 1994. Three volumes. IEEE catalog number 94CH3319-1.

IEEE:1994:PIC

[IEE94c] IEEE, editor. Proceedings of the 1994 International Conference on Computer Languages: May 16–19, 1994, Toulouse, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5642-5. ISSN 1074-8970. LCCN QB76.7 .I34 1994.

IEEE:1995:IIC

[IEE95a] IEEE, editor. 1995 IEEE International Conference on Evolutionary Computation, the University of Western Australia, Perth, Western Australia, 29 November–1 December, 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-2759-4 (softbound), 0-7803-2760-8 (microfiche). LCCN QA76.87.I342 1995. Two volumes. IEEE catalog number: 95TH8099.

IEEE:1995:PCA

[IEE95b] IEEE, editor. Proceedings / the 11th Conference on Artificial Intelligence for Applications, February 20–23, 1995, Los Angeles, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN PCAAEF. ISBN 0-8186-7070-3 (paperback), 0-7803-2522-2 (casebound) 0-7803-2523-0 (microfiche). LCCN Q334 .C65 1995.

IEEE:1996:PEI

[IEE96] IEEE, editor. Proceedings.

Eighth IEEE International Conference on Tools with Artificial Intelligence. November 16–19, 1996, Toulouse, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. CODEN PCTIFX. ISBN 0-8186-7686-8. ISSN 1063-6730. LCCN Q 334 I564 1996.

IEEE:1997:IIC

[IEE97a] IEEE, editor. 1997 IEEE International Conference on Acoustics, Speech, and Signal Processing: April 21–24, 1997, Munich, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. CODEN IPRODJ. ISBN 0-8186-7920-4 (casebound), 0-8186-7919-0, 0-8186-7921-2 (microfiche). ISSN 0736-7791. LCCN TK 7882 S65 I16 1997. Five volumes. IEEE Computer Society Press order number PR07919. IEEE Order Plan catalog number 97CB36052.

IEEE:1997:IAA

[IEE97b] IEEE, editor. IEEE Aerospace Applications Conference. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN ???? ISSN 1095-323X.

                                        LCCN ???? URL http://iel.
                                        ihs.com/ (Go to IEL; then
                                         Aerospace); then Conferences;
                                        then Search.

Ito:1985:IAC

[IO85] Ikuo Ito and Takao Onozawa. Intelligent aspect of CAD for mechanical design: (the conceptual design of simple object). Nippon Kikai Gakkai Ronbunshu, C Hen, 51(465): 1158–1161, May 1985. CODEN NKCHDB. ISSN 0387-5024.

Jacobs:1994:UPC

[Jac94] James H. Jacobs. UCL+P: a persistent Common Lisp. Thesis (ph. d.), Department of Computer Science, University of Utah, Salt Lake City, UT, USA, 1994. ix + 114 pp.

Janota:1990:DIP

[Jan90] Paul Janota, editor. Digital image processing and visual communications technologies in the earth and atmospheric sciences: 18–19 April 1990, Orlando, Florida, volume 1301 of SPIE Proceedings. SPIE Optical Engineering Press, Bellingham, WA, USA, 1990. CODEN PSISDG. ISBN 0-8194-0352-0. ISSN 0277-786X (print), 1996-756X (electronic). LCCN QC871 .D54 1990.

Jellesed:1989:DES

[Jel89] R. H. Jellesed. Developing an expert system to link the piping designer to computer-aided design. American Society of Mechanical Engineers, Pressure Vessels and Piping Division (Publication) PVP, 177:49–54, 1989. CODEN AMPPD5. ISSN 0277-027X.

Jambunathan:1992:DIF

[JLHB92] K. Jambunathan, E. Lai, S. L.

Hartle, and B. L. Button. Development of an intelligent front end using LISP. Applications of Artificial Intelligence in Engineering, pages 228–243, 1992. CODEN AAIEEO.

Jacob:1990:BCP

[JR90] G. Jacob and R. Rorschach.

Building a commercial-quality parallel Common Lisp system. In Steels [Ste90b], pages 101– 108. ISBN ???? LCCN ????

Jacobs:1994:SSP

[JS94] J. H. Jacobs and M. R. Swanson. Syntax and semantics of a persistent Common Lisp. In ACM [ACM94b], pages 103– 112. ISBN 0-89791-643-3. LCCN QA76.73.L23A26 1994. ACM order no. 522940.

Jacobs:1997:UDI

[JS97] J. H. Jacobs and M. R. Swanson. UCL+P — defining and implementing persistent Common Lisp. Lisp and Symbolic Computation, 10(1):5– 38, June 1997. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

                                        URL http://www.wkap.nl/
                                        oasis.htm/134677.

Jaaksoo:1991:ACW

[JU91] U. Jaaksoo and V. I. Utkin, editors. Automatic control: World Congress 1990: in the service of mankind: proceedings of the 11th Triennial World Congress of the International Federation of Automatic Control, Tallinn, Estonia, USSR, 13–17 August 1990. Pergamon, New York, NY, USA, 1991. CODEN ISYSEK. ISBN 0-08-040174-0. LCCN TJ212.2 .I58 1990. Six volumes.

Kung:1990:PIC

[K^{+}90] Sun-Yuan Kung et al., editors. Proceedings of the International Conference on Application Specific Array Processors: September 5–7, 1990, Princeton, New Jersey. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN ???? LCCN TK7874.6 .I57 1990.

Katz:1995:HFC

[K^{+}95] Irvin R. Katz et al., editors. Human factors in computing systems: CHI ’95 conference proceedings, “Mosaic of Creativity,” Denver, Colorado, May 7–11, 1995. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-201-84705-1. LCCN QA76.9.H85 C44 1995.

Kilian:1995:CLE

[KA95] Jens Kilian and Heinz-Peter Arndt. Common Lisp as an embedded extension language. Hewlett-Packard Journal: technical information from the laboratories of Hewlett-Packard Company, 46(5):69–73, October 1995. CODEN HPJOAX. ISSN 0018-1153. URL http://

www.hp.com/hpj/toc-10-95.
html.

Kaisler:1986:LAP

[Kai86] Stephen H. Kaisler. Lisp: an AI programming language. IEEE Journal of Oceanic Engineering, OE-11(4):468–473, October 1986. CODEN IJOEDY. ISSN 0364-9059.

Kantrowitz:1991:PUC

[Kan91] Mark Kantrowitz. Portable utilities for Common Lisp user guide and implementation notes. Research paper CMU-CS-91-143, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, May 1991. ii + 74 pp.

Karjalainen:1989:LBH

[Kar89] Matti Karjalainen. Lisp-based high-level programming environment for the TMS320C30. Proceedings of the International Conference on Acoustics, Speech, and Signal Processing, 2:1150–1153, 1989. CO-DEN IPRODJ. ISSN 0736-7791. IEEE catalog number 89CH2673-2.

Krueger:1987:HLP

[KB87] Steven D. Krueger and Patrick W. Bosshart. High Lisp performance on a high-level language processor. In IEEE [IEE87c], pages 120–123. CODEN DC-SIDU. ISBN 0-8186-0764-5 (paperback), 0-8186-8764-9 (casebound), 0-8186-4764-7 (microfiche). LCCN QA76 .I49 1987. IEEE catalog number 87CH2409-1. Computer Society Order number 764.

Kawamura:1987:DCE

[KBS^{+}87] K. Kawamura, G. Beale, J. Schaffer, B.-J. Hsieh, S. Padalkar, J. Rodriguez-Moscoso, F. Vinz, and K. Fernandez. Development of a coupled expert system for the spacecraft attitude control problem. NASA conference publication, 2491:125– 132, 1987. CODEN NACPDX. ISSN 0191-7811.

Kiczales:1991:AMP

[KdR91] Gregor Kiczales and Jim des Rivieres. The art of the metaobject protocol. MIT Press, Cambridge, MA, USA, 1991. ISBN 0-262-11158-6 (hardcover), 0-262-61074-4 (paperback). viii + 335 pp. LCCN QA76.73.C28 K53 1991.

Kempe:1991:CTAa

[Kem91] Johann Kempe. CL/TB: an Allegro Common Lisp programming interface for transbase. Technical report TUM-I9106, Mathematisches Institut und Institut für Informatik der Technischen Universitãt München, Munich, Germany, April 1991. 56 pp.

Keene:1989:OOP

[KG89] Sonya E. Keene and Dan Gerson. Object-oriented programming in Common LISP: a programmer’s guide to CLOS. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-17589-4 (paperback). xix + 266 pp. LCCN QA76.73.C28 K441 1989.

Karjalainen:1988:BDC

[KH88] Matti Karjalainen and Seppo Helle. Block diagram compilation and graphical editing of DSP algorithms in the Quicksig system. In IEEE [IEE88d], pages 1057–1060. CODEN PICSDI. ISBN 951-721-240-2. ISSN 0271-4310. LCCN ???? Available from IEEE Service Cent (catalog number 88CH2458-8). Piscataway, NJ, USA.

Kempf:1987:EC

[KHDS87] James Kempf, Warren Harris, Roy D’Souza, and Alan Snyder. Experience with Common-Loops. ACM SIGPLAN Notices, 22(12):214–226, December 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Kimbrough:1990:PCC

[Kim90] Kerry Kimbrough. Programming CLUE: A Common Lisp X toolkit. Integrated Computer Solutions, Inc., Cambridge, MA, USA, 1990. ISBN ???? 29 + 4 + 13 pp. LCCN ???? Tutorial notes. “Xhibition 90, X and beyond, May 21-24, San Jose Convention Center, San Jose, California.

Kingsland:1989:PTA

[Kin89] Lawrence C. Kingsland III, editor. Proceedings: the thirteenth annual Symposium on Computer Applications in Medical Care, November 5–8, 1989, Washington, DC. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN PCMCDC. ISBN 0-8186-1985-6. ISSN 0195-4210. IEEE Catalog Number 89TH0286-5. IEEE Computer Society order number 1985.

Kersten:1993:TLO

[KK93] P. R. Kersten and A. C.

Kak. Tutorial on LISP object-oriented programming for blackboard computation (solving the radar tracking problem). International Journal of Intelligent Systems, 8(5): 617–669, June 1993. CODEN IJISED. ISSN 0884-8173.

Kunze:2022:HFL

[KK22] Fritz Kunze and Lauren Kunze.

The history of Franz and Lisp. IEEE Annals of the History of Computing, 44(1):103–108, January/March 2022. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Kota:1989:CMC

[KL89] S. Kota and Chun-Liang Lee.

Computational model for the conceptual design synthesis of hydraulic circuits. In Riley et al. [RCG^{+}89], pages 161– 170. CODEN COENEF. ISBN 0-7918-0357-0 (vol. 1) 0-7918-0358-9 (vol. 2). LCCN TA 345 I5488 1989. Two volumes.

Kim:1995:BRE

[KL95] Bonn-Oh Kim and Sang M.

Lee. Bond rating expert system for industrial companies. Expert Systems with Applications, 9(1):63–70, 1995. CO-DEN ESAPEH. ISSN 0957-4174.

Klassner:2004:ELI

[Kla04] Frank Klassner. Enhancing Lisp instruction with RCXLisp and robotics. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 36(1):214–218, March 2004. CODEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

Kempe:1991:CTAb

[KLF^{+}91] J. Kempe, T. Lenz, B. Freitag, H. Shütz, and G. Specht. CL/TB an allegro common Lisp programming interface for TransBase. ACM SIGPLAN Notices, 26(8):60–69, August 1991. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karp:1994:GCG

[KLSW94] Peter D. Karp, John D. Lowrance, Thomas M. Strat, and David E. Wilkins. Grasper-CL graph management system. Lisp and Symbolic Computation, 7(4):251–290, 1994. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Kwak:1990:RBM

[KM90] S. H. Kwak and R. B. McGhee. Rule-based motion coordination for a hexapod walking machine. Advanced Robotics, 4 (3):263–282, 1990. CODEN ADROEI. ISSN 0169-1864.

Kohl:1994:DIU

[KM94] Charles Kohl and Joe Mundy.

Development of the image understanding environment. In IEEE [IEE94a], pages 443– 447. ISBN 0-8186-5827-4 (casebound), 0-8186-5825-8 (paperback), 0-8186-5826-6 (microfiche). ISSN 1063-6919. LCCN TA1632 .I36 1994. IEEE Computer Society Press order number 5825-02. IEEE catalog number 94CH3405-8.

Kaufmann:1996:AIS

[KM96] Matt Kaufmann and J. Strother Moore. ACL2: An industrial strength version of Nqthm. COMPASS — Proceedings of the Annual Conference on Computer Assurance, pages 23–34, 1996. CODEN CCOAEA. IEEE catalog number 96CH35960.

Kaufmann:1997:IST

[KM97] Matt Kaufmann and J. S.

Moore. Industrial strength theorem prover for a logic based on Common Lisp. IEEE Transactions on Software Engineering, 23(4):203–213, April 1997. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Knafl:1989:PSC

[Kna89] George J. Knafl, editor. Proceedings 13: September 20-22, 1989: Compsac 89 / the Thirteenth Annual International Computer Software and Applications Conference, Hyatt Orlando Hotel, Orlando, Florida. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. CODEN PSICD2. ISBN 0-8186-1964-3. ISSN 0730-6512. LCCN QA76.6 .C62951 1989.

Koomen:1987:TTR

[Koo87] Johannes A. G. M. Koomen. The TIMELOGIC temporal reasoning system in Common Lisp. Technical report TR 231, University of Rochester, Dept. of Computer Science, Rochester, NY, USA, 1987. i + 33 pp.

Koschmann:1990:CLC

[Kos90] Timothy D. Koschmann. The Common Lisp companion. John Wiley and Sons, New York, NY, USA; London, UK; Sydney, Australia, 1990. ISBN 0-471-50308-8. xvi + 459 pp. LCCN QA76.73.L23 K67 1990.

Kempf:1988:LLP

[KPB^{+}88] James Kempf, Andreas Paepcke, Brian Beach, Joseph Mohan, Brom Mahbod, and Alan Snyder. Language level persistence for an object-oriented application programming platform. In Hoevel et al. [HSKS88], pages 424–433. CODEN PHISD7. ISBN 0-8186-0841-2 (paperback: vol. 1), 0-8186-8841-6 (hardcover: vol. 1), 0-8186-4841-4 (microfiche: vol. 1). ISSN 0073-1129. LCCN QA76.9.S88H375 1988. Four volumes. Computer Society order number 841. IEEE catalog number 88TH0209-7.

Krueger:1985:MVC

[Kru85] Steven D. Krueger. Microprogrammed VLSI computer for Lisp. WESCON Conference Record, 1985. CODEN WCREDI. ISSN 1044-6036, 0083-8837.

Kaiser:1988:DIM

[KSv88] F. Kaiser, L. Stok, and R. van den Born. Design and implementation of a module library to support the structural synthesis. Eut report, Eindhoven University of Technology, Department of Electrical Engineering, Eindhoven, The Netherlands, January 1988. 53 pp.

Kamio:1988:ICL

[KT88] Minoru Kamio and Ikuo Takeuchi. On the implementation of Common Lisp and compiler on Elis. Denki Tsushin Kenkyusho Kenkyu Jitsuyoka Hokoku, 37(2):153–157, 1988. CODEN DTKKAA. ISSN 0415-3200.

Kuchlin:1997:IPI

[Küc97] Wolfgang W. Küchlin, editor. ISSAC ’97. Proceedings of the 1997 International Symposium on Symbolic and Algebraic Computation, July 21– 23, 1997, Maui, Hawaii. ACM Press, New York, NY 10036, USA, 1997. ISBN ???? LCCN ????

Kalet:1997:EPR

[KUJP97] I. J. Kalet, J. M. Unger, J. P. Jacky, and M. H. Phillips. Experience programming radiotherapy applications in Common Lisp. In Leavitt and Starkschall [LS97], pages 445– 448. ISBN 0-944838-77-4. LCCN ????

Klaassen:1989:TPL

[Kv89] Arno J. Klaassen and Anton M. van Wezenbeek. A three-processor Lisp machine architecture based on statistical analysis of Common Lisp programs. ACM SIGPLAN Notices, 24(9):86–91, September 1989. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Karjalainen:1990:PCG

[KVJ90] Matti Karjalainen, Matti Vartiainen, and Kari Jarvinen. A programming and code generation environment for nonhomogeneous DSP multiprocessors. In IEEE [IEE90a], pages 2881–2884. CODEN PICSDI. ISBN ???? ISSN 0271-4310. LCCN TK 7801 I22 1990. Four volumes.

Lamer:1990:DOS

[Lam90] Daniel L. Lamer. Distributed, operating system based, blackboard architecture for real-time control. In ACM [ACM90a], pages 99–108. ISBN 0-89791-372-8. LCCN Q334.I556 1990. Two volumes.

Lamkins:2005:SLH

[Lam05] David B. Lamkins. Successful Lisp: How to Understand and Use Common Lisp. bookfix.com, Frankfurt am Main, Germany, 2005. ISBN 3-937526-00-5. xx + 360 pp. LCCN PPPP. Foreword by Richard P. Gabriel.

Langemyr:1986:CSC

[Lan86] Lars Langemyr. Converting SAC-2 code to Lisp. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 20(4): 11–13, December 1986. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lane:1988:PPAb

[Lan88a] Darrin J. Lane. Porting Pan I to Allegro Common Lisp: research project. Master of science, plan ii, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, 1988. 35 pp.

Lane:1988:PPAa

[Lan88b]  Darrin J. Lane.  Porting

Pan I to Allegro COMMON pLisp. Report UCB/CSD 88/453, University of California, Berkeley, Computer Science Division, Berkeley, CA, USA, 1988. 35 pp.

Lane:1989:MAH

[Lan89a] A. Lane. MicroExplorer in action! (hybrid Macintosh/Lisp technology). Byte Magazine, 14 (12):247–249, November 1989. CODEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Langa:1989:DES

[Lan89b] J. M. Langa. Diagnostic expert system for Bayer digestion ratio control. In Light Metals: Proceedings of Sessions, AIME Annual Meeting (Warrendale, Pennsylvania), pages 9–19. ????, ????, 1989. CO-DEN LMPMDF. ISBN ???? ISSN 0147-0809. LCCN ????

Landi:1992:USA

[Lan92] William Landi. Undecidability of static analysis. ACM Letters on Programming Languages and Systems, 1(4):323– 337, December 1992. CO-DEN ALPSE8. ISSN 1057-4514 (print), 1557-7384 (electronic).

Larsson:1996:DBE

[Lar96] Jan Eric Larsson. Diagnosis based on explicit meansend models. Artificial Intelligence, 80(1):29–93, January 1996. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Lee:1995:KBA

[LC95] Anita Lee and Chun Hung Cheng. Knowledge-based approach for task allocation in heterogeneous distributed computer systems. Expert Systems, 12(4):303–312, November 1995. CODEN EXSYEX. ISSN 0266-4720.

Leitao:2008:UGO

[Lei08] A. Menezes Leitão. UCL-GLORP — an ORM for Common Lisp. J.UCS: Journal of Universal Computer Science, 14(20):3333–??, ???? 2008. CODEN ???? ISSN 0948-

6968. URL http://www.jucs.
org/jucs_14_20/ucl_glorp_
an_orm.

Levelt:1995:IPI

[Lev95] A. H. M. (Antonius Henricus Maria) Levelt, editor. IS-SAC’95: proceedings of the 1995 International Symposium on Symbolic and Algebraic Computation, July 10– 12, 1995, Montréal, Canada, ISSAC — Proceedings. ACM Press, New York, NY 10036, USA, 1995. ISBN 0-89791-699-9. LCCN A 76.95 I59 1995.

Lim:1987:LMA

[Lim87] Raymond S. Lim. Lisp Machine architecture issues. In IEEE [IEE87c], pages 116–119. CO-DEN DCSIDU. ISBN 0-8186-0764-5 (paperback), 0-8186-8764-9 (casebound), 0-8186-4764-7 (microfiche). LCCN QA76 .I49 1987. IEEE catalog number 87CH2409-1. Computer Society Order number 764.

Lin:1991:RBE

[Lin91] Chaung Lin. Rule-based expert system for automatic control rod pattern generation for boiling water reactors. Nuclear Technology, 95(1):1–8, July 1991. CODEN NUTYBB. ISSN 0029-5450 (print), 1943-7471 (electronic).

Lieslehto:1987:ESI

[LK87] J. Lieslehto and H. N. Koivo. Expert system for interaction analysis of multivariable systems. In IEEE [IEE87f], pages 961–962. CODEN PRACEO. ISBN ???? ISSN 0743-1619. LCCN TJ 213 A43 1987. Three volumes. IEEE Catalog Number: 87CH2378-8.

Leung:1989:LPA

[LLC89] K. S. Leung, K. H. Lee, and S. M. Cheang. Lists processing for artificial intelligence applications. Microprocessing and Microprogramming, 26(4):271– 287, December 1989. CODEN MMICDT. ISSN 0165-6074 (print), 1878-7061 (electronic).

Lue:1989:ESL

[LM89] Wen-Jeng Lue and L. P. Mc-Namee. Extracting schematiclike information from CMOS circuit net-lists. Proceedings — Design Automation Conference, pages 690–693, 1989. CO-DEN PDAWDJ. ISSN 0146-7123. IEEE catalog number 89CH2734-2.

Lawless:1991:UCC

[LM91] Jo A. Lawless and Molly M. Miller. Understanding CLOS: the COMMON LISP OBJECT SYSTEM. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1991. ISBN 1-55558-064-5 (Digital Press), 0-13-735283-2 (Prentice Hall). ix + 192 pp. LCCN QA76.64.L38 1991.

Ling:1989:ESC

[LMM89a] N. Ling, M. E. Malowany, and A. S. Malowany. Expert system for circuit board repair. In IEEE [IEE89b], pages 253– 256. ISBN ???? LCCN TK5102.5 .I4 1989. IEEE Cat No. 89CH2691-4.

Ling:1989:ESS

[LMM89b] N. Ling, M. E. Malowany, and A. S. Malowany. Expert system scheduler for circuit-board repair task in a robotic workcell. In Riley et al. [RCG^{+}89], pages 523–529. CODEN COENEF. ISBN 0-7918-0357-0 (vol. 1) 0-7918-0358-9 (vol. 2). LCCN TA 345 I5488 1989. Two volumes.

Loui:1989:UMC

[Lou89] Ronald P. Loui. User’s manual for CCRC: (Common Lisp Version) computing reference classes, statistical reasoning shell v. 2.5. Technical report WUCS-89-08, Washington University, Dept. of Computer Science, Saint Louis, MO, USA, June 1989. 39 pp.

Loia:1994:IOO

[LQ94] Vincenzo Loia and Michel Quaggetto. Integrating objectoriented paradigms and logic programming: the OPLA language. Knowledge — Based Software Engineering — Conference Proceedings, KBSE, pages 158–164, 1994. ISSN 1068-3062. IEEE catalog number 94TH0664-3.

Loia:1996:OSD

[LQ96] Vincenzo Loia and Michel Quaggetto. OPLA system: designing complex systems in an object-oriented logic programming framework. The Computer Journal, 39(1):20– 35, 1996. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic).

Li:1992:SBS

[LS92a] Xiaofeng Li and Osamu Shigo. Simulation-based SDL support system. In IEEE [IEE92], pages 284–291. ISBN 0-8186-2830-8. LCCN QA76.758.I573 1992.

Li:1992:SSS

[LS92b] Xiaofeng Li and Osamu Shigo. Simulation-based SDL support system. In IEEE [IEE92], pages 284–291. ISBN 0-8186-2830-8. LCCN QA76.758.I573 1992.

Leavitt:1997:UCR

[LS97] D. D. Leavitt and G. Starkschall, editors. Use of computers in radiation therapy. Proceedings of the 12th international conference on the use of computers in radiation therapy: International conference; 12th — May 1997, Salt Lake City, UT, Proceedings of the International Conference on the Use of Computers in Radiation Therapy. Medical Physics Publishing, Madison, WI, USA, 1997.

ISBN 0-944838-77-4. LCCN ????

Lubars:1991:RDD

[Lub91] Mitchell D. Lubars. Representing design dependencies in an issue-based style. IEEE Software, 8(4):81–89, July 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Lucid:1992:CLI

[Luc92a] Lucid, Inc. Common Lisp interface manager, user’s guide. Lucid, Inc., Menlo Park, CA, USA, 1992. ISBN ???? various pp. LCCN ????

Lucid:1992:XUG

[Luc92b] Lucid, Inc. XLT User’s Guide:

an X-based programming environment for Lucid Common Lisp. Lucid, Inc., Menlo Park, CA, USA, first edition (december 1992). edition, 1992. ISBN ???? various pp. LCCN ????

Li:1995:CPP

[LW95] Liwei Li and Paul S. Wang. The CL–PVM package. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 29(3–4): 2–8, December 1995. CO-DEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Lee:1994:UMA

[LZ94] Arthur H. Lee and Joseph L.

Zachary. Using metaprogramming to add persistence to CLOS. In IEEE [IEE94c], pages 136–147. ISBN 0-8186-5642-5. ISSN 1074-8970. LCCN QB76.7 .I34 1994.

Lee:1995:RM

[LZ95] Arthur H. Lee and Joseph L. Zachary. Reflections on metaprogramming. IEEE Transactions on Software Engineering, 21(11):883–893, November 1995. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Lee:1997:ASP

[LZ97] Arthur H. Lee and Joseph L. Zachary. Adding support for persistence to CLOS via its metaobject protocol. Lisp and Symbolic Computation, 10 (1):39–60, June 1997. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (elec-

tronic). URL http://www.
wkap.nl/oasis.htm/134678.

Ma:1989:TTI

[Ma89] Kwan-Lin Ma. TICL: a type inference system of Common Lisp. Thesis (m.s.), Department of Computer Science, University of Utah, Salt Lake City, UT, USA, 1989. xi + 70 pp.

Mattsson:1991:KSR

[MA91] S. E. Mattsson and M. Andersson. Kernel for system representation. In Jaaksoo and Utkin [JU91], pages 373–378. CODEN ISYSEK. ISBN 0-08-040174-0. LCCN TJ212.2 .I58 1990. Six volumes.

MacLachlan:1991:CCL

[Mac91] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-91-108, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, February 1991. vi + 168 pp. This is a revised version of Technical Report CMU-CS-87-156.

MacLachlan:1992:CCL

[Mac92a] Rob MacLachlan. CMU Common Lisp user’s manual. Research paper CMU-CS-92-161, School of Computer Science, Carnegie Mellon University, Pittsburgh, PA, USA, July 1992. v + 142 pp. Supersedes Technical Reports CMU-CS-87-156 and CMU-CS-91-108.

MacLachlan:1992:PCC

[Mac92b] Robert A. MacLachlan. Python compiler for CMU Common Lisp. In ACM [ACM92], pages 235–246. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Maida:1991:MMM

[Mai91] Anthony S. Maida. Maintaining mental models of agents who have existential misconceptions. Artificial Intelligence, 50(3):331–383, August 1991. CODEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Malec:1991:DSD

[Mal91] J. Malec. Dynamic scene description system. ODYS. In Jaaksoo and Utkin [JU91], pages 159–164. CODEN ISY-SEK. ISBN 0-08-040174-0. LCCN TJ212.2 .I58 1990. Six volumes.

Margolies:2004:ACL

[Mar04]  D. Margolies.  The ANSI

Common Lisp reference book. Apress, Berkeley, CA, USA, 2004. ISBN 1-59059-220-4. 1000 (est.) pp. LCCN ????

Matsui:1990:OOI

[Mat90] T. Matsui. Object-oriented implementation of EusLisp. Denshi Gijutsu Sogo Kenkyusho Iho/Bulletin of the Electrotechnical Laboratory, 54(8):18–40, 1990. CODEN DGSKAR. ISSN 0366-9092.

Medina-Bulo:2010:VCL

[MBPLRR10] Inmaculada Medina-Bulo, Francisco Palomo-Lozano, and José-Luis Ruiz-Reina. A verified Common Lisp. Journal of Symbolic Computation, 45(1): 96–123, January 2010. CO-DEN JSYCEH. ISSN 0747-7171 (print), 1095-855X (electronic).

Manivannan:1990:RBA

[MC90] S. Manivannan and Richard Cotton. Rule-based approach to robotic collision avoidance in a flexible assembly cell. American Society of Mechanical Engineers, Production Engineering Division (Publication) PED, 47:235–244, 1990. CODEN ASMDDU.

McDonald:1987:CCL

[McD87a] David B. McDonald. CMU Common Lisp user’s manual: Mash/IBM RT PC edition. Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. iii + 94 pp.

McDonald:1987:ECL

[McD87b] David B. McDonald. An efficient Common Lisp for the IBM RT PC. Research paper. Carnegie-Mellon University. Computer Science Department; 87-134 Carnegie-Mellon University. Computer Science Dept. Research paper; 87-134. CMU-CS-87-134, Carnegie Mellon University, Computer Science Department, Pittsburgh, PA, USA, 1987. 18 pp.

McDonald:1989:CCL

[McD89a] David B. McDonald. CMU Common Lisp user’s manual Mach/IBM RT edition. Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1989. v + 120 pp. Revised version of technical report CMU-CS-87-156.

McDonald:1989:OOP

[McD89b] J. A. McDonald. Objectoriented programming for linear algebra. ACM SIGPLAN Notices, 24(10):175–184, October 1989. CODEN SINODQ.

ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Miller:1994:FPI

[MCd^{+}94] Philip Miller, Michael Clancy, Andrea diSessa, Jeremy Roschelle, Marcia Linn, Michael Eisenberg, Mark Guzdial, Mitchell Resnick, and Elliot Soloway. Future of programming instruction. SIGCSE Bulletin (ACM Special Interest Group on Computer Science Education), 26(1), March 1994. CO-DEN SIGSD3. ISSN 0097-8418 (print), 2331-3927 (electronic).

McLean:1991:GEC

[MCH91] Francis G. McLean, De-Wayne A. Campbell, and David W. Harris, editors. Geotechnical Engineering Congress 1991: proceedings of the Congress / sponsored by the Geotechnical Engineering Division of the American Society of Civil Engineers in cooperation with the University of Colorado and U.S. Bureau of Reclamation; hosted by the Colorado Section, ASCE, Boulder, Colorado, June 10–12, 1991. American Society of Civil Engineers, Dallas, Texas, 1991. CO-DEN GSPUER. ISBN 0-87262-806-x. ISSN 0895-0563. LCCN TA703.5.G47 1991. Two volumes.

McKay:1991:CCL

[McK91] Scott McKay. CLIM: the Common Lisp interface manager. Communications of the ACM, 34(9):58–59, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0001-0782/
114675.html.

Mingins:1997:TPC

[MDM97] Christine Mingins, Roger Duke, and Bertrand Meyer, editors. TOOLS 25: Proceedings of the 1997 Conference on Technology of Object-Oriented Languages and Systems Nov 24–28 1997 1997 Melbourne, Australia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-8485-2, 0-8186-8487-9 (microfiche). LCCN QA76.64.T66 1997b.

Morrison:1995:CDP

[ME95] Ernest E. Morrison, Jr. and Robert M. Ebeling. Computation of dynamic passive earth pressure using LISP. Journal of Computing in Civil Engineering, 9(4):285–291, October 1995. CODEN JCCEE5. ISSN 0887-3801.

McDonald:1987:IDC

[MFW87] David B. McDonald, Scott E.

Fahlman, and Skef Wholey. Internal design of CMU Common Lisp on the IBM RT PC. Research paper CMU-CS-87-157, Carnegie Mellon University, Computer Science Dept., Pittsburgh, PA, USA, 1987. ii + 64 pp.

McDonald:1988:IDC

[MFW88] David B. McDonald, Scott E. Fahlman, and Skef Wholey. Internal design of CMU Common Lisp on the IBM RT PC. Technical report CMU-CS-87-157, Computer Science Dept., Carnegie Mellon University, Pittsburgh, PA, USA, April 1988. ii + 64 pp.

Meystel:1990:PII

[MHG90] Alex Meystel, Jayantha Herath, and Steve Gray, editors. Proceedings / 5th IEEE International Symposium on Intelligent Control 1990, September 5–7, 1990, Philadelphia, Pennsylvania. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2108-7 (paperback), 0-8186-9108-5 (casebound), 0-8186-6108-9 (microfiche). LCCN TJ 217.5 I33 1990a. Two volumes. IEEE Computer Society Press Order Number 2108. IEEE Catalog Number 90TH0333-5.

Messimer:1996:CDM

[MHMR96] Sherri L. Messimer, John M. Henshaw, John Montgomery, and John Rogers. Composite design and manufacturing critiquing system. Artificial Intelligence for Engineering Design, Analysis and Manufacturing: AIEDAM, 10(1):65– 79, January 1996. CODEN AIEMEG. ISSN 0890-0604.

Milner:1988:CLT

[Mil88] Wendy L. Milner. Common Lisp: a tutorial. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1988. ISBN 0-13-152844-0. xx + 521 pp. LCCN QA76.73.L23 M55 1988.

McGehearty:1986:PPE

[MK86] Patrick F. McGehearty and Edward J. Krall. Potentials for parallel execution of Common Lisp programs. In Hwang et al. [HJS86], pages 696–702. CO-DEN PCPADL. ISBN 0-8186-0724-6. ISSN 0190-3918. LCCN QA76.6 .I581 1986. IEEE Service Cent. Piscataway, NJ, USA.

Ma:1990:TTI

[MK90] Kwan-Liu Ma and Robert R.

Kessler. TICL – A type inference system for Common Lisp. Software—Practice and Experience, 20(6):593–625 (or 593–623??), June 1990. CO-DEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Mayfield:1991:PMC

[MN91] Blayne E. Mayfield and Jin Cheon Na. Prolog methods for Common LISP + Flavors. OSU-CS-TR; 91-17. Oklahoma State University, Dept. of Computer Science, Stillwater, OK, USA, 1991. ISBN ???? 12 pp. LCCN ????

Moon:1986:OOP

[Moo86]  David A. Moon.  Object-

oriented programming with Flavors. ACM SIGPLAN Notices, 21(11):1–8, November 1986. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Moore:1999:MCP

[Moo99] J. Strother Moore. Mechanically checked proof of a multiprocessor result via a uniprocessor view. Formal Methods in System Design, 14(2):213– 228, March 1999. CODEN FMSDE6. ISSN 0925-9856.

Moore:2012:MLF

[Moo12] J. Strother Moore. Metalevel features in an industrialstrength theorem prover. ACM SIGPLAN Notices, 47(1):425– 426, January 2012. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Mayaram:1987:COM

[MP87] Kartikeya Mayaram and Donald O. Pederson. Codecs: an object-oriented mixed-level circuit and device simulator. In IEEE [IEE87a], pages 604– 607. CODEN PICSDI. ISBN ???? LCCN TK 7801 I22 1987. Three volumes.

Monnier:2020:EEL

[MS20] Stefan Monnier and Michael Sperber. Evolution of Emacs Lisp. Proceedings of the ACM on Programming Languages (PACMPL), 4(HOPL):74:1– 74:55, June 2020. URL https:

//dl.acm.org/doi/abs/10.
1145/3386324.

Morrey:1993:USC

[MSBH93] Ian Morrey, Jawed Siddiqi, Graham Buckberry, and Richard Hibberd. Use of a specification construction and animation tool to teach formal methods. In IEEE [IEE93a], pages 327–333. CODEN PSICD2. ISBN 0-8186-4440-0 (paperback), 0-8186-4442-7 (casebound), 0-8186-4441-9 (microfiche). ISSN 0730-6512. LCCN QA76.6.C6295 1993. IEEE catalog number 93CH3349-8.

Maegawa:1985:FLM

[MSD^{+}85] Hirotoshi Maegawa, Toshifumi Saito, Toshio Doi, Takeshi Nishikawa, and Hiroshi Yasui. Fast Lisp machine and listevaluation processor Eval II — processor architecture and hardware configuration. Journal of Information Processing, 8(2):121–126, 1985. CODEN JIPRDE. ISSN 0387-6101.

Mizoguchi:1995:TAI

[MTI95] Riichiro Mizoguchi, Yuri Tijerino, and Mitsuru Ikeda. Task analysis interview based on task ontology. Expert Systems with Applications, 9(1):15–25, 1995. CODEN ESAPEH. ISSN 0957-4174.

MuQqoz:1995:MTW

[Mu 95] J. Mu Qq oz. Manual TIG welding and control chart advisor: a prototype program in Common Lisp using pattern matching techniques. Quality Engineering, 7(2):277–??, 1995. CODEN QUENE7. ISSN 0898-2112.

Muller:1991:LLT

[Mul91] Hans Muller. LispView: leverage through integration. Communications of the ACM, 34(9): 59–61, September 1991. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/toc/Abstracts/
                                        0001-0782/126333.html.

Mundy:1995:IUE

[Mun95a] Joseph L. Mundy. Image understanding environment program. IEEE expert: intelligent systems and their applications, 10(6):64–73, December 1995. CODEN IEEXE7. ISSN 0885-9000.

Munoz:1994:MTW

[Mun95b] J. Munoz. Manual TIG welding and control chart advisor. A prototype program in Common Lisp using pattern matching techniques. Quality Engineering, 7(2):277–??, 1994/ 1995. CODEN QUENE7. ISSN 0898-2112.

Murray:1990:FCL

[Mur90] K. E. Murray. The future of Common Lisp: Higher performance through parallelism. In Steels [Ste90b], pages 35–42. ISBN ???? LCCN ????

MuThoz:1995:MTW

[MuT95] J. MuThoz. Manual TIG welding and control chart advisor: a prototype program in Common Lisp using pattern matching techniques. Quality engineering, 7(2):277–??, ???? 1995. ISSN 0898-2112.

Myers:1995:GAU

[Mye95] Brad A. Myers. Garnet and amulet user interface development environments. In Katz et al. [K^{+}95], pages 334–?? ISBN 0-201-84705-1. LCCN QA76.9.H85 C44 1995.

Ng:1990:IFI

[N^{+}90] Peter A. Ng et al., editors. ICSI ’90, the First International Conference on Systems Integration: April 23–26, 1990, Headquarters Plaza Hotel, Morristown, New Jersey, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-9027-5. LCCN TS176.I645 1990. IEEE catalog number 90TH0309-5.

Naef:1984:AIE

[Nae84] F. E. Naef. Artificial intelligence expert program for a restricted memory microcomputer. In IEEE [IEE84b], pages 4–10. CODEN PCI-CDQ. ISBN 0-8186-0525-1 (paperback), 0-8186-8525-5 (hardcover). LCCN QA75.5 .C58 1984, TK7885.A1 C53 1984. IEEE catalog no. 84CH2017-2.

Narayanaswamy:1988:SAA

[Nar88a] K. Narayanaswamy. Static analysis: an aid to program maintenance and development. In IEEE [IEE88h], pages 492– 499. CODEN PHISD7. ISBN 0-8186-0842-0. ISSN 0073-1129. LCCN ????

Narayanaswamy:1988:SAB

[Nar88b] K. Narayanaswamy. Static analysis-based program evolution support in the Common Lisp framework. Proceedings — International Conference on Software Engineering, pages 222–230, 1988. CODEN PCSEDE. ISBN 0-89791-258-6. ISSN 0270-5257. IEEE catalog number 88CH2554-4.

Nachtsheim:1988:SFE

[NC88] Philip R. Nachtsheim and Ames Research Center. Simulating futures in extended Common LISP. NASA technical memorandum A-88176, National Aeronautics and Space Administration, Ames Research Center, Moffett Field, CA, USA, 1988.

Nishida:1990:PPE

[ND90] Toyoaki Nishida and Shuji Doshita. PSX. A program that explores phase portraits of twodimensional piecewise linear differential equations. Memoirs of the Faculty of Engineering, Kyoto University, 52(pt 4):

311–355, October 1990. CO-DEN MEKYAC. ISSN 0023-6063.

Nishida:1995:QAB

[ND95] Toyoaki Nishida and Shuji Doshita. Qualitative analysis of behavior of systems of piecewise linear differential equations with two state variables. Artificial Intelligence, 75(1):3–29, May 1995. CO-DEN AINTBB. ISSN 0004-3702 (print), 1872-7921 (electronic).

Neusss:2002:UCL

[Neu02]  Nicolas Neusß.  On using

Common Lisp for scientific computing. Sonderforschungsbereich 359, Ruprecht-Karls-Universität Heidelberg, Interdisziplinäres Zentrum für Wissenschaftliches Rechnen, Heidelberg, Germany, 2002. 10 pp.

Neuss:2003:UCL

[Neu03] Nicolas Neuss. On using common Lisp for scientific computing. In Bänsch [Bän03], pages 237–245. CODEN LNCSA6. ISBN 3-642-62406-5 (print), 3-642-19014-6 (e-book). ISSN 1439-7358. LCCN ???? URL

http://link.springer.com/
content/pdf/10.1007/978-
3-642-19014-8_11.

Nierstrasz:1993:EOO

[Nie93] Oscar M. Nierstrasz, editor. ECOOP ’93, objectoriented programming, 7th European conference, Kaiserslautern, Germany, July 26– 30, 1993: proceedings, number 788 in Arbeitspapiere GMD 1993. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-57120-5, 0-387-57120-5. LCCN QA76.64 .E95 1993.

Nishida:1994:QRA

[Nis94]  Toyoaki Nishida.  Qualita-

tive reasoning for automated exploration for chaos. In Anonymous [Ano94b], pages 1211–1216. CODEN PNAIEE. ISBN 0-262-61102-3. LCCN Q334.N36 1994. Two volumes.

Nelson:1989:CDT

[NJB89] Gordon R. Nelson, Michael A. Johnston, and John L. Blue. CODER. A design tool for aerospace applications. Simulation series, 20(4):165–169, March 1989. CODEN SISEDL. ISSN 0735-9276.

Nelson:1991:WSC

[NKC91] Barry L. Nelson, W. David Kelton, and Gordon M. Clark, editors. 1991 Winter Simulation Conference proceedings: December 8–11, 1991, the Arizona Biltmore, Phoenix, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. CODEN WSCPDK. ISBN 0-7803-0181-X (??invalid checksum??). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1991. IEEE catalog no. 91CH3050-2.

Norvig:1991:PAI

[Nor91] Peter Norvig. Paradigms of artificial intelligence programming: case studies in Common LISP. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1991. ISBN 1-55860-191-0. xxviii + 946 pp. LCCN QA76.6.N687 1992.

Norvig:1992:PAI

[Nor92] Peter Norvig. Paradigms of artificial intelligence programming: case studies in Common LISP. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 1992. ISBN ???? xxviii + 946 pp. LCCN ????

Norvig:2004:PAI

[Nor04] Peter Norvig. Paradigms of artificial intelligence programming: case studies in Common Lisp. Morgan Kaufmann Publishers, Los Altos, CA 94022, USA, 2004. ISBN 1-55860-191-0. xxviii + 946 pp. LCCN ????

Noyes:1992:AIC

[Noy92] James L. Noyes. Artificial intelligence with Common Lisp: fundamentals of symbolic and numeric processing. D. C. Heath and Company, Lexington, MA, USA, 1992. ISBN 0-669-19473-5. xvii + 542 + A80 pp. LCCN QA76.73.C62 N68 1992 Bar.

Nord:1988:EAS

[NP88] Robert L. Nord and Frank Pfenning. The Ergo Attribute System. ACM SIGSOFT Software Engineering Notes, 13(5): 110–120, November 1988. CO-DEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Nann:1989:DSS

[NRK89] Steven R. Nann, Asok Ray, and Soundar Kumara. Decision support system for realtime control and monitoring of dynamical processes. In IEEE [IEE89e], pages 361–365. ISBN ???? LCCN TJ 212.2 A44 1989. Three volumes. IEEE catalog number: 89CH2772-2.

Nann:1991:DSS

[NRK91] Steven R. Nann, Asok Ray, and Soundar Kumara. Decision support system for realtime monitoring and control of dynamical processes. International Journal of Intelligent Systems, 6(7):739–758, December 1991. CODEN IJISED. ISSN 0884-8173.

Nagasaka:1992:TCL

[NSI^{+}92] Atsushi Nagasaka, Yoshihiro Shintani, Tanji Ito, Hiroshi Gomi, and Junichi Takahashi. Tachyon Common Lisp: An efficient and portable implementation of CLtL2. In ACM [ACM92], pages 270–277. ISBN 0-89791-481-3, 0-89791-483-X. LCCN QA76.73.L23A26 1992. LISP pointers, volume V, number 1, January-March 1992. ACM order no. 552920.

Narikawa:1994:IID

[NSS94] N. Narikawa, T. Sato, and N. Sasaki. Integrated and intelligent database systems for plant engineering framework. In Bocks et al. [BPR^{+}94], pages 13–20. ISBN 0-7918-1281-2. LCCN TA345 .A85 1994.

Nayama:1990:DES

[NTS90] Michisuke Nayama, Masahiko Toyoda, and Susumu Shono. Development of an expert system for vacuum brazing. Yosetsu Gakkai Ronbunshu/Quarterly Journal of the Japan Welding Society, 8(3): 359–364, August 1990. CO-DEN YGRODU. ISSN 0288-4771.

Nyman:1987:TCL

[Nym87] Lars Nyman. Transliteration of Common Lisp into executable C code. Thesis (m.s.), Department of Computer Engineering and Science, Case Western Reserve University, Cleveland, OH, USA, 1987. vi + 105 pp.

Olthoff:1989:ASM

[OK89] Walter Olthoff and James Kempf. An algebraic specification of method combination for the Common Lisp Object System. Lisp and Symbolic Computation, 2(1):115–152, February 1989. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Okuno:1987:FAF

[OOT87] Hiroshi G. Okuno, Nobuyasu Osato, and Ikuo Takeuchi. Firmware approach to fast Lisp interpreter. In ACM [ACM87], pages 1–11. CODEN MAWMDB. ISBN 0-89791-250-0. ISSN 0361-2163. LCCN ???? ACM order no. 520870. IEEE catalog no. 87TH0208-9. IEEE Computer Society order no. 831.

Ong:1986:KPE

[OSD86] K. K. Ong, R. E. Seviora, and P. Dasiewicz. Knowledge-based position estimation for a multisensor house robot. In Sriram and Adey [SA86], pages 119–130. ISBN 0-905451-47-3. LCCN TA174.I48 1986. Two volumes.

Paeke:1993:OOP

[Pae93] Andreas Paeke, editor. Objectoriented programming: the CLOS perspective. MIT Press, Cambridge, MA, USA, 1993. ISBN 0-262-16136-2. viii + 352 pp. LCCN QA76.64 .O27 1993.

Praehofer:1993:EDB

[PAR93] H. Praehofer, F. Auernig, and G. Reisinger. Environment for DEVS-based multiformalism simulation in Common Lisp/CLOS. Discrete Event Dynamic Systems: Theory and Applications, 3(2-3): 119–??, July 1993. CODEN DEDAEE. ISSN 0924-6703.

Pavelle:1985:PPC

[Pav85] Richard Pavelle. The power of present computer algebra systems: MACSYMA on a LISP-Machine. In Caviness [Cav85], page ?? CODEN LNCSD9. ISBN 0-387-15984-

3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA155.7.E4 E85 1985. URL

http://link.springer-ny.
com/link/service/series/
0558/tocs/t0204.htm;   http:
//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-
9743&volume=204.

Pfenning:1988:HOA

[PE88] F. Pfenning and C. Elliot.

Higher-order abstract syntax. ACM SIGPLAN Notices, 23 (7):199–208, July 1988. CO-DEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Perret-Gallix:1992:NCT

[PG92] D. Perret-Gallix, editor. New computing techniques in physics research II: proceedings of the Second International Workshop on Software Engineering, Artificial Intelligence and Expert Systems in High Energy and Nuclear Physics, January 13– 18, 1992, L’Agelonde France-Telecom La Londe-les-Maures (France). World Scientific Publishing Co. Pte. Ltd., P. O. Box 128, Farrer Road, Singapore 9128, 1992. ISBN 981-02-1122-8. LCCN QC793.47.E4 I58 1992; Q334.I595 1992.

Pinkowski:1985:LSG

[Pin85] Ben Pinkowski. Lisp-based system for generating diagnostic keys. In IEEE [IEE85c], pages 592–596. ISBN 0-8186-0688-6, 0-8186-8688-X (hardcover), 0-8186-4688-8 (microfiche). LCCN Q 334 C66 1985.

Pirolli:1986:CMC

[Pir86] Peter Pirolli. Cognitive model and computer tutor for programming recursion. Human-Computer Interaction, 2(4): 319–355, 1986. CODEN HCOIEL. ISSN 0737-0024.

Parikh:1991:ESC

[PK91] Sanjay A. Parikh and N. S. V. Kameswara Rao. Expert system for civil engineering applications. In McLean et al. [MCH91], pages 413–421. CO-DEN GSPUER. ISBN 0-87262-806-x. ISSN 0895-0563. LCCN TA703.5.G47 1991. Two volumes.

Pase:1989:PME

[PKSA89] Bill Pase, Sentot Kromodimoeljo, Karen. Summerskill, and I. P. Sharp Associates. Porting m-EVES to Common Lisp: final report. I.P. Sharp Associates Ltd., Ottawa, ON, Canada, 1989. ISBN ???? 14 pp. LCCN ????

Prakash:1993:STA

[PM93] G. Ravi Prakash and H. N. Mahabala. SVEPOA: a tool to aid verification and validation of OPS5-based AI applications. International Journal of Expert Systems, 6(2):193–236, 1993. CODEN IJSYED. ISSN 0894-9077.

Pedersen:1990:BEP

[PO90] Peder C. Pedersen and Banu Onaral, editors. Biomedical engineering perspectives: health care technologies for the 1990’s and beyond: proceedings of the Twelfth Annual International Conference of the IEEE Engineering in Medicine and Biology Society, Philadelphia, Pennsylvania, USA, November 1–4, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. CODEN CEMBAD. ISBN 0-87942-559-8 (paperback), 0-87942-560-1 (casebound). ISSN 0589-1019. LCCN R856.A2I344 1990. Five volumes. IEEE catalog number 90CH2936-3.

Poggi:1996:IAO

[Pog96] Agostino Poggi. Integrating agents and objects to develop distributed AI systems. In IEEE [IEE96], pages 443–444. CODEN PCTIFX. ISBN 0-8186-7686-8. ISSN 1063-6730. LCCN Q 334 I564 1996.

Pierce:1988:IKA

[PRD88] Jonathan A. Pierce, Linda L.

Rodi, and Robert E. Dalton. Interactive knowledge acquisition for fault isolation and repair. In IEEE [IEE88e], page 421. ISBN 0-8186-0837-4. LCCN Q 334 C66 1988. IEEE catalog number 88CH2552-8.

Puppe:1987:RCE

[Pup87] Frank Puppe. Requirements for a classification expert system shell and their realization in Med2. Applied Artificial Intelligence, 1(2):163–171, 1987. CODEN AAINEH.

PurdueUniversity:1990:EECa

[Pur90a] Purdue University. ECN: engineering computer network: Sun COMMON LISP 4.0. Purdue University, West Lafayette, IN, USA, 1990. ISBN ???? various pp. LCCN ???? Two volumes in one.

PurdueUniversity:1990:EECb

[Pur90b] Purdue University. ECN: engineering computer network: Sun COMMON LISP: the Window tool kit/programmers’ manual. Purdue University, West Lafayette, IN, USA, 1990. ISBN ???? various pp. LCCN ????

Quednau:1976:LLB

[Que76] H. D. Quednau. Libaform: a Lisp-based computer language for symbolic formula manipulation. Angewandte Informatik/Applied Informatics, 18 (4):168–174, April 1976. CO-DEN AWIFA7.

Queinnec:1996:LSP

[Que96]  Christian Queinnec.  LISP

in small pieces. Cambridge University Press, Cambridge, UK, 1996. ISBN 0-521-56247-3. xx + 514 pp. LCCN QA76.73.L23Q4613 1996.

Rathke:1993:OPF

[Rat93] Christian Rathke. Objectoriented programming and frame-based knowledge representation. In IEEE [IEE93b], pages 95–98. ISBN 0-8186-4202-5 (casebound), 0-8186-4200-9 (paperback). ISSN 1063-6730. LCCN Q334 .I5634 1993. IEEE Catalog Number 93CH3325-8. IEEE Computer Society Press Order Number 4200-02.

Rathore:2012:CA

[Rat12] Amit Rathore. Clojure in action. Manning, Greenwich, CT, USA, 2012. ISBN 1-935182-59-5 (paperback). xxii + 410 pp. LCCN QA76.7 .R375 2012.

Riva:1998:DEK

[RBLS98] A. Riva, R. Bellazzi, G. Lanzola, and M. Stefanelli. Development environment for knowledge-based medical applications on the World-Wide Web. Artificial Intelligence in Medicine, 14(3):279–293, November 1998. CODEN AIMEEW. ISSN 0933-3657.

Riley:1989:CEP

[RCG^{+}89] Donald R. Riley, T. J. Cokonis, Gary A. Gabriele, et al., editors. Computers in engineering 1989: proceedings of the 1989 ASME International Computers in Engineering Conference and Exposition, July 30–August 3, 1989, Anaheim, California. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1989. CODEN COENEF. ISBN 0-7918-0357-0 (vol. 1) 0-7918-0358-9 (vol. 2). LCCN TA 345 I5488 1989. Two volumes.

Reid:1996:RFF

[Rei96] J. K. Reid. Remark on “Fast Floating-Point Processing in Common Lisp”. ACM Transactions on Mathematical Software, 22(4):496–497, December 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Renkes:1986:SCP

[Ren86] Gordon D. Renkes. Symbolic computer programs applied to group theory. In ACS Symposium Series, pages 176– 185. American Chemical Society, Washington, DC, USA, 1986. CODEN ACSMC8. ISBN ???? ISSN 0097-6156. LCCN ????

Ressler:1992:RLP

[Res92] Eugene K. Ressler. Random list permutations in place. Information Processing Letters, 43(5):271–275, October 1992. CODEN IFPLAT. ISSN 0020-0190 (print), 1872-6119 (electronic).

Ruggiero:1990:UFC

[RGG90] C. Ruggiero, M. Giacomini, and S. Gaglio. A user friendly cognitive modelling environment. In Pedersen and Onaral [PO90], pages 1381–1383. CO-DEN CEMBAD. ISBN 0-87942-559-8 (paperback), 0-87942-560-1 (casebound). ISSN 0589-1019. LCCN R856.A2I344 1990. Five volumes. IEEE catalog number 90CH2936-3.

Richardson:1990:LCL

[Ric90] C. Richardson. LispWorks:

a Common Lisp programming environment for Unix workstations. In Steels [Ste90b], pages 127–134. ISBN ???? LCCN ????

Richards:1994:PC

[Ric94] Barry Richards. Planning with constraints. IEE Colloquium (Digest), 038:1/1–1/7, February 1994. CODEN DCILDN. ISSN 0963-3308.

Riley:1991:CLW

[Ril91] John H. Riley. A Common Lisp workbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1991. ISBN 0-13-155797-1 (paperback). 192 pp. LCCN QA76.73.C62R55 1991.

Riley:1992:CLW

[Ril92] John H. Riley. A Common LISP workbook. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1992. ISBN 0-13-155797-1. xiv + 242 pp. LCCN QA76.73.C62R55 1991.

Rahkila:1997:IDT

[RK97] Martti Rahkila and Matti Karjalainen. Interactive DSP tutorial on the Web. In IEEE [IEE97a], pages 2253–2256. CODEN IPRODJ. ISBN 0-8186-7920-4 (casebound), 0-8186-7919-0, 0-8186-7921-2 (microfiche). ISSN 0736-7791. LCCN TK 7882 S65 I16 1997. Five volumes. IEEE Computer Society Press order number PR07919. IEEE Order Plan catalog number 97CB36052.

Richardson:1993:CLI

[RMM93] Chris Richardson, Colin Meldrum, and Scott Mckay. The Common LISP Interface Manager. PC AI, 7(5):38–41, September 1993. ISSN 0894-0711.

Richardson:1994:CLI

[RMM94] Chris Richardson, Colin Meldrum, and Scott McKay. Common LISP interface manager architecture. PC AI Intelligent Solutions for Desktop Computers, 8(2):34–35, March–April 1994. CODEN PCAIE5. ISSN 0894-0711.

Ruokangas:1988:RTC

[Ruo88] Corinne C. Ruokangas. Realtime control for manufacturing Space Shuttle main engines: Work in progress. NASA conference publication, 3013:5– 18, 1988. CODEN NACPDX. ISSN 0191-7811.

Reeker:1986:PPA

[RW86] Larry H. Reeker and Kenneth Wauchope. Pattern-directed processing in Ada. In IEEE [IEE86c], pages 49–56. ISBN 0-8186-0718-1 (paperback), 0-8186-8718-5 (hard), 0-8186-4718-3 (microfiche). LCCN QA76.73.A35 I33 1986. IEEE catalog number 86CH2281-4. IEEE Computer Society order number 718.

Rich:1990:RPD

[RW90] Charles Rich and Linda M.

Wills. Recognizing a program’s design: a graph-parsing approach. IEEE Software, 7(1): 82–89, January 1990. CO-DEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Roberts:1996:UO

[RZ96] Jonathan Roberts and Dan Zigmond. Using OODCE. Dr. Dobb’s Journal of Software Tools, 21(3):98–101, March 1996. CODEN DDJTEQ. ISSN 1044-789X.

Sriram:1986:PIC

[SA86] D. Sriram and R. Adey, editors. Proceedings of the 1st International Conference on the Application of Artificial Intelligence in Engineering Problems, Southampton University, U.K., April 1986. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-905451-47-3. LCCN TA174.I48 1986. Two volumes.

Salem:1990:ABC

[Sal90] Nabeel H. Salem. AutoCAD based computer-aided mechanism design system. American Society of Mechanical Engineers, Design Engineering Division (Publication) DE, 26:331–339, 1990. CODEN AMEDEH.

Samet:1977:MDF

[Sam77] Hanan Samet. Machine description facility for compiler testing. IEEE Transactions on Software Engineering, SE-3(5): 343–351, September 1977. CO-DEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Smith:1989:UKS

[SB89] Barry R. Smith and Carolyn P. Banda. Use of a knowledgebased system to assess aircrew training requirements as part of conceptual design. In IEEE [IEE89a], pages 105– 110. CODEN PICYE3. ISBN ???? ISSN 0884-3627. LCCN Q300 .I485 1989. Three volumes. IEEE catalog number 89CH2809-2.

Schryver:1992:OOQ

[Sch92] Jack C. Schryver. Objectoriented qualitative simulation of human mental models of complex systems. IEEE Transactions on Systems, Man, and Cybernetics, 22(3):526– 541, May–June 1992. CODEN ISYMAW. ISSN 0018-9472, 1083-4427.

Schottstaedt:1994:MTX

[Sch94] Bill Schottstaedt. Machine tongues XVII: CLM: Music V meets Common Lisp. Computer Music Journal, 18(2):30– ??, Summer 1994. CODEN CMUJDY. ISSN 0148-9267 (print), 1531-5169 (electronic).

Sansonnet:1985:RTO

[SCP^{+}85] J. P. Sansonnet, P. Clere, E. Papon, S. Bourgault, and B. Grandjean. Real time oriented machine for A.I. applications. In IEEE [IEE85a], pages 99–106. ISBN 0-8186-0649-5 (paperback), 0-8186-8649-9 (hardcover), 0-8186-4649-7 (microfiche). LCCN QA76.5 .I544 1985.

Seibel:2005:PCL

[Sei05]  Peter Seibel.  Practical

Common Lisp. Apress, Berkeley, CA, USA, 2005. ISBN 1-59059-239-5 (hardcover). xxv + 499 pp. LCCN QA76.73.L23 S45 2005. URL

ftp://uiarchive.cso.uiuc.
edu/pub/etext/gutenberg/;
 http://www.gigamonkeys.
com/book/;   http://www.
loc.gov/catdir/toc/ecip058/
2005005859.html.

Sevinc:1991:ECL

[Sev91] Suleyman Sevinc. Extending Common LISP Object System for discrete event modeling and simulation. In Nelson et al. [NKC91], pages 204–206. CO-DEN WSCPDK. ISBN 0-7803-0181-X (??invalid checksum??). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1991. IEEE catalog no. 91CH3050-2.

Steele:1993:EL

[SG93] Guy L. Steele, Jr. and Richard P. Gabriel. The evolution of Lisp. ACM SIGPLAN Notices, 28(3):231–270, March 1993. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

                                        URL http://www.acm.org:
                                        80/pubs/citations/proceedings/
                                        plan/154766/p231-steele/.

Silbert:1986:TDA

[SGBL86] Mark Silbert, Steven Gabai, William Bremner, and David Loewenstern. Tool for development of AI hybrid systems. In IEEE [IEE86e], pages 601–607. ISBN ???? LCCN TL693.D55 1986. IEEE catalog number: 86CH2359-8.

Sansen:1989:SSA

[SGW89] Willy Sansen, Georges Glelen, and Herman Walscharts. Symbolic simulator for analog circuits. In Wuorinen [Wuo89], pages 204–205. CODEN DT-PCDE. ISBN ???? LCCN TK7870 .I58 1989. IEEE catalog no. 89CH2684-9.

Skrien:1996:CTL

[SH96] D. Skrien and J. Hallstrom. Composers toolbox, a Lispbased precomposition environment: general system overview and an introduction to the package’s new Krayola MIDI paint module. Computers and Mathematics with Applications, 32(1):49–55, July 1996. CO-DEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Shapiro:1986:LIA

[Sha86] Stuart Charles Shapiro. LISP:

an interactive approach. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1986. ISBN 0-88175-069-7 (paperback). x + 150 pp. LCCN QA76.73.L23 S53 1986. See also revised edition [Sha92].

Shankar:1991:ZIT

[Sha91] Renuka Shankar. Z-scheduler:

Integrating theories on scheduling behavior into a computational model. In IEEE [IEE91a], pages 1367–1371. CODEN PICYE3. ISBN 0-7803-0234-6 (set: casebound), 0-7803-0233-8 (set: softbound), 0-7803-0235-4 (microfiche). ISSN 0884-3627. LCCN TA 168 I19 1991; Q300 .I485 1991. Three volumes. IEEE catalog number 91CH3067-6.

Shapiro:1992:CLI

[Sha92] Stuart Charles Shapiro. Common LISP: an interactive approach. Principles of computer science series. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1992. ISBN 0-7167-8218-9. xxi + 358 pp. LCCN QA 76.73 L23 S53 1992. Revised edition of [Sha86].

Shivers:1988:CFA

[Shi88] O. Shivers. Control flow analysis in Scheme. ACM SIGPLAN Notices, 23(7):164–174, July 1988. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/53990/p164-shivers/.

Sigal:1985:KAS

[Sig85] Burton C. Sigal. Knowledgebased advisory system for casualty procedures. In IEEE [IEE85b], pages 1364–1368. CODEN NASEA9. ISBN ???? LCCN TL 693 N37 1985. Two volumes. IEEE catalog no. 86CH2189-9.

Sato:1989:RCL

[SIG89] Mitsuhisa Sato, Shuichi Ichikawa, and Eiichi Goto. Run-time checking in Lisp by integrating memory addressing and range checking. In IEEE [IEE89f], pages 290–297. CO-DEN CPAADU. ISBN 0-8186-8948-X (casebound), 0-8186-1948-1 (paperback). ISSN 0149-7111. LCCN QA 76.9 A73 C65 v.17 no.3 1989. IEEE Computer Society order number 1948. IEEE catalog number 89CH2705-2. ACM order number 415890.

Simpson:1987:CLI

[Sim87] Rosemary Simpson. Common Lisp: the index: a crossreferenced index to Common Lisp: the language. Coral Software Corp., Cambridge, MA, USA, 1987. ISBN ???? 71 pp. LCCN QA76.73.C28 S55 1987.

Sitaram:1993:HC

[Sit93] Dorai Sitaram. Handling control. In ACM SIGPLAN PLDI ’93 [ACM93], pages 147–155. ISBN 0-89791-598-4. LCCN QA76.7.A152 v. 28, no. 6.

URL http://www.acm.org:
80/pubs/citations/proceedings/
pldi/155090/p147-sitaram/.

Smith:1990:RSA

[SK90a] Barry Smith and Joseph Katz.

Range scheduling aid. In ACM [ACM90a], pages 275–

280. ISBN 0-89791-372-8. LCCN Q334.I556 1990. Two volumes.

Su:1990:CDP

[SK90b] Chau-Chin Su and Charles R.

Kime. Computer-aided design of pseudoexhaustive BIST for semiregular circuits. In IEEE [IEE90c], pages 680–

689. CODEN DITCDP. ISBN 0-8186-2064-1 (paperback), 0-8186-9064-X (hardcover), 0-8186-6064-3 (microfiche). ISSN 0743-1686. LCCN TK 7874 I593 1990. IEEE catalog number 90CH2910-6.

Shapiro:1996:SHB

[SK96] Bruce A. Shapiro and Wojciech Kasprzak. STRUCTURELAB: a heterogeneous bioinformatics system for RNA structure analysis. Journal of Molecular Graphics, 14(4):194–205, August 1996. CODEN JMGRDV. ISSN 0263-7855.

Slade:1998:OOC

[Sla98] Stephen Slade. Object-oriented Common LISP. Prentice-Hall PTR, Upper Saddle River, NJ 07458, USA, 1998. ISBN 0-13-605940-6. xviii + 774 pp. LCCN QA76.64 .S576 1998.

Sly:1993:ABI

[Sly93] David Sly. AutoCAD based industrial layout planning and material flow analysis in factory flow and plan. In Evans et al. [E^{+}93], pages 281–284. CODEN WSCPDK. ISBN 0-7803-1381-X (casebound). ISSN 0275-0708, 0743-1902. LCCN QA 76.9 C65 W56 1993. IEEE catalog no. 93CH3338-1.

Shevenell:1988:LER

[SM88] Michael P. Shevenell and Charles Millett. LISP environment for real-time ocean systems. In IEEE [IEE88c], pages 1276–1279. ISBN ???? LCCN GC2 .O32 1988. Four volumes. IEEE catalog number 88-CH2585-8.

Siskind:1993:NLS

[SM93] Jeffrey Mark Siskind and David Allen McAllester. Nondeterministic Lisp as a substrate for constraint logic programming. In AAAI [AAA93], pages 133–136. CODEN PNAIEE. ISBN 0-262-51071-5. LCCN Q334.N36 1993.

Springfield:1988:GST

[SMK^{+}88] J. Springfield, A. Mutammara, G. Karsai, G. E. Cook, J. Sztipanovits, and K. Fernandez. Graphic simulation test bed for robotics applications in a workstation environment. NASA conference publication, 3013:303–312, 1988. CODEN NACPDX. ISSN 0191-7811.

Snyder:1986:COC

[Sny86a] Alan Snyder. CommonObjects: an overview (Common Lisp extension). ACM SIG-PLAN Notices, 21(10):19–28, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Snyder:1986:COO

[Sny86b] Alan Snyder. CommonObjects: object-oriented programming in Common Lisp. ACM SIGPLAN Notices, 21(10):189, October 1986. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Schmidt:1991:CES

[SO91] Heinz W. Schmidt and Stephen M.

Omohundro. CLOS, Eiffel, and Sather: A comparison. Technical Report TR-91-047, International Computer Science Institute, Berkeley, CA, September 1991.

Spenke:1992:GDA

[Spe92a] Michael Spenke. GINA demo applications version 2.1 for Common Lisp. Arbeitspapiere der GMD 613, Gesellschaft für Mathematik und Datenverarbeitung mbH, Sankt Augustin, Germany, February 1992. 69 pp.

Spenke:1992:GUM

[Spe92b] Michael Spenke. GINA user manual version 2.1 for Common Lisp. Arbeitspapiere der GMD 614, Gesellschaft für Mathematik und Datenverarbeitung mbH, Sankt Augustin, Germany, February 1992. 64 pp.

Swartout:1991:EKS

[SPM91] William Swartout, Cecile Paris, and Johanna Moore. Explanations in knowledge systems: Design for explainable expert systems. IEEE expert: intelligent systems and their applications, 6(3):58–64, June 1991. CODEN IEEXE7. ISSN 0885-9000.

Sugimoto:1989:CLB

[SST89] S. Sugimoto, T. Sakaguchi, and K. Tabata. Concurrent LISP based on lexical scope. In Knafl [Kna89], pages 700–707. CO-DEN PSICD2. ISBN 0-8186-1964-3. ISSN 0730-6512. LCCN QA76.6 .C62951 1989.

Steele:1984:CLL

[Ste84] Guy L. Steele Jr. Common Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1984. ISBN 0-13-152851-3, 0-932376-41-X (paperback). xii + 465 pp. LCCN QA76.73.L23 S73 1984. US$22.00. With contributions by Scott E. Fahlman and Richard P. Gabriel and David A. Moon and Daniel L. Weinreb. See also [Tat87].

Steele:1986:CLC

[Ste86]  Guy Lewis Steele.  Com-

mon LISP, and the connection machine. Computer Language Magazine, 11(8):??, August 1986. CODEN COMLEF. ISSN 0749-2839.

Steele:1989:GCL

[Ste89] David J. (David Jay) Steele.

Golden Common LISP: a hands-on approach. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-41653-0 (paperback). xxxii + 543 pp. LCCN QA76.73.C28 S74 1989 Bar.

Steele:1990:CLL

[Ste90a]  Guy L. Steele Jr.  Com-

mon Lisp: The Language. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, second edition, 1990. ISBN 1-55558-041-6 (paperback), 1-55558-042-4 (hardcover), 0-08-050226-1 (e-book), 0-13-152414-3 (Prentice-Hall). xxiii + 1029 pp. LCCN QA76.73.L23 S73 1990. With contributions by Scott E. Fahlman and others, and with contributions to the second edition by Daniel G. Bobrow and others. See also [?, ?].

Steels:1990:EPF

[Ste90b] Luc Steels, editor. Europal 90: Proceedings of the First European Conference on the Practical Application of Lisp, Churchill College, Cambridge. Europal, Dorking, UK, 1990. ISBN ???? LCCN ????

Stewart:1991:BNS

[Ste91a] Rocky L. Stewart. Biosphere 2 nerve system. Communications of the ACM, 34(9):69– 71, September 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.acm.org/
                                        pubs/toc/Abstracts/0001-
                                        0782/114680.html.

Stewart:1991:LRG

[Ste91b] William P. (William Paschal) Stewart. LINK — a representation of gene linkage in Common LISP. Thesis (b.s.), Massachusetts Institute of Technology, Department of Electrical Engineering and Computer Science, Cambridge, MA, USA, 1991. 35 pp.

Stoyenko:1994:SRS

[Sto94] Alexander D. Stoyenko. SUPRA-RPC: SUbprogram PaRAmeters in remote procedure calls. Software—Practice and Experience, 24(1):27–49, January 1994. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

SunMicrosystems:1988:SCL

[Sun88] Sun Microsystems. Sun Common Lisp 3.0 manual set. Sun Microsystems, 2550 Garcia Avenue, Mountain View, CA 94043, USA, August 1988. Nine parts.

Saaty:1997:INF

[SV97] T. L. Saaty and L. G. Vargas. Implementing neural firing: Towards a new technology. Mathematical and Computer Modelling (Oxford), 26 (4):113–124, August 1997. CO-DEN MCMOEG. ISSN 0895-7177 (print), 1872-9479 (electronic).

Swaine:1990:PPG

[Swa90a]  M. Swaine.  Programming

paradigms: Getting CLOS (Lisp extension). Dr. Dobb’s Journal of Software Tools, 15 (3):122–125, March 1990. CO-DEN DDJOEB. ISSN 1044-789X.

Swaine:1990:PPb

[Swa90b] Michael Swaine. Programming paradigms. Dr. Dobb’s Journal of Software Tools, 15(2):118– ??, February 1990. CODEN DDJOEB. ISSN 1044-789X.

Steele:1985:GCL

[SWH85] Guy Steele, Patrick Henry Winston, and Berthold Horn. Golden Common LISP. Gold Hill Computers, Cambridge, Mass., version 1.01 edition, 1985. ISBN 0-09-175805-X. LCCN QA76.73.L23 G64 1985 Type EXPLAIN UCBS Tolman Microcomputer Facility. Title from title screen. Title on reference manual: COMMON LISP. Title on text: LISP.

Reference manual by Guy L. Steele, Jr.; text by Patrick Henry Winston and Berthold Klaus Paul Horn. Computer program language for creating artificial intelligence programs. Includes tutorial programs.

Symbolics:1986:SCL

[Sym86]  Symbolics, Inc.  Symbolics

Common Lisp: language dictionary: 2b. Symbolics, Inc., Concord, MA, USA, 1986.

Symbolics:1988:SCL

[Sym88]  Symbolics, Inc.  Symbolics

Common Lisp: language concepts, 2A. Symbolics, Cambridge, MA, USA, February 1988. ISBN ???? xvi + 828 pp. LCCN ????

Sugihara:1992:ALD

[SYM92] Kazuo Sugihara, Kazunari Yamamoto, and Isao Miyamoto. Automatic layout of diagrams for software specification. In IEEE [IEE92], pages 245–252. ISBN 0-8186-2830-8. LCCN QA76.758.I573 1992.

Tichy:1989:NNL

[TAH89] Walter F. Tichy, Rolf L. Adams, and Lars Holter. NLH/E: a natural language help system. In IEEE [IEE89c], pages 364–374. CODEN PCSEDE. ISBN 0-89791-258-6, 0-8186-1941-4 (paperback), 0-8186-5941-6 (microfiche). ISSN 0270-5257. LCCN QA 76.6 I553

Tanimoto:1990:EAI

[Tan90] S. (Steven) Tanimoto. The elements of artificial intelligence using Common LISP. Principles of computer science series. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, 1990. ISBN 0-7167-8230-8. xxiv + 562 pp. LCCN Q336.T37 1990.

Tanimoto:1995:EAI

[Tan95] S. (Steven) Tanimoto. The elements of artificial intelligence using Common Lisp. Computer Science Press, Inc., 11 Taft Court, Rockville, MD 20850, USA, second edition, 1995. ISBN 0-7167-8269-3. xxix + 823 pp. LCCN Q336.T37.

Tatar:1987:PGC

[Tat87] Deborah G. Tatar. A programmer’s guide to COMMON LISP. Digital Press, 12 Crosby Drive, Bedford, MA 01730, USA, 1987. ISBN 0-932376-87-8 (paperback). x + 327 pp. LCCN QA76.73.C28T38 1987.

Tate:2010:SLS

[Tat10] Bruce Tate. Seven languages in seven weeks: a pragmatic guide to learning programming languages. Pragmatic programmers. Pragmatic Bookshelf, Raleigh, NC, USA, 2010. ISBN 1-934356-59-X (paperback). xvii + 317 pp. LCCN QA76.7 .T38 2010.

Taube:1990:CMM

[Tau90] H. Taube. Common Music: a music composition language in Common Lisp and CLOS. In Steels [Ste90b], pages 235–244. ISBN ???? LCCN ????

Taube:1991:CMM

[Tau91]  Heinrich Taube.  Common

Music: a music composition language in Common Lisp and CLOS. Computer Music Journal, 15(2):21–32, Summer 1991. CODEN CMUJDY. ISSN 0148-9267 (print), 1531-5169 (electronic).

Tou:1994:PSG

[TBE^{+}94] Ivan Tou, Steven Berson, Gerald Estrin, Yadran Eterovic, and Elsie Wu. Prototyping synchronous group applications. Computer, 27(5):48–56, May 1994. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Teitelman:1976:CCL

[Tei76] W. Teitelman. Clisp: Conversational Lisp. IEEE Transactions on Computers, C-25 (4):354–357, April 1976. CO-DEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (electronic).

                                        URL http://ieeexplore.
                                        ieee.org/stamp/stamp.jsp?
                                        tp=&arnumber=1674617.

Teitell:1988:PFA

[Tei88] Murray Teitell, editor. Proceedings of the fourth annual artificial intelligence & advanced computer technology conference: held as part of the artificial intelligence & advanced computer technology conference/exhibition, Long Beach Convention Center, Long Beach, California, May 4–6, 1988. Tower Conference Management, Glen Ellyn, IL, USA, 1988. ISBN ???? LCCN Q335.A7868 1988.

Taylor:1986:ESL

[THL^{+}86] George S. Taylor, Paul N.

Hilfinger, James R. Larus, David A. Patterson, and Benjamin G. Zorn. Evaluation of the Spur Lisp architecture. In IEEE [IEE86b], pages 444– 452. CODEN CPAADU. ISBN 0-8186-0719-X (pbk.), 0-8186-8719-3. ISSN 0149-7111. LCCN QA 76.9 A73 I56 1986.

Thrift:1988:CLR

[Thr88] Philip Thrift. Common LISP relations: An extension of LISP for logic programming. In IEEE [IEE88a], pages 332–337. ISBN 0-8186-0874-9. LCCN QA76.7 .I561 1988. IEEE catalog number 88CH2647-6. Computer Society order number 874.

Tibbitts:1993:FSC

[Tib93] B. R. Tibbitts. Flexible simulation of a complex semiconductor manufacturing line using a rule-based system. IBM Journal of Research and Development, 37(4):507–521, July 1993. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Touretzky:1990:BRC

[Tou90a] David S. Touretzky. Book review: Common Lisp: the Reference, by Franz, Inc. Addison-Wesley Publishing Company, 1988. softcover; 919 pages. Lisp and Symbolic Computation, 3(1):101–102, January 1990. CODEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Touretzky:1990:CLG

[Tou90b] David S. Touretzky. Common LISP: a gentle introduction to symbolic computation. Benjamin/Cummings Pub. Co., Menlo Park, CA, USA, 1990. ISBN 0-8053-0492-4. xxiii + 442 + 122 pp. LCCN QA76.73.C28T68 1990.

Tipnis:1988:CEP

[TPB^{+}88] V. A. Tipnis, E. M. Patton, D. W. Bennett, et al., editors. Computers in engineering, 1988: proceedings of the 1988 ASME International Computers in Engineering Conference and Exhibition, July 31–August 4, 1988, San Francisco, California. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1988. LCCN TA 345 A86 1988. Three volumes.

Tsang:1990:DMS

[TS90] C. P. Tsang and S. E. Smith. Designing a microcode synthesis system. In IEEE [IEE90d], pages 437–441. ISBN 0-87942-557-1, 0-87942-556-3 (paperback), 0-87942-558-X (microfiche). LCCN TK 5105.5 I34 1990. Two volumes. IEEE catalog number 90CH2866-2.

Torrance:1988:DFI

[TSL^{+}88] Randy R. Torrance, Steve P.

Sadler, J. Pierre Lamoureux, Jean Lamarche, David J. Frank, and François Leveille. Design of a fully integrated graphics system. IEEE Journal of Solid-State Circuits, 23 (2):368–376, April 1988. CO-DEN IJSCBC. ISSN 0018-9200 (print), 1558-173X (electronic).

Tana:1987:VGS

[TW87] David Tana and Neil Weste.

Virtual Grid Symbolic Layout 1987. In IEEE [IEE87e], pages 192–196. ISBN 0-8186-0802-1 (paperback), 0-8186-4802-3 (microfiche), 0-8186-8802-5 (casebound). LCCN TK 7888.4 I23 1987. IEEE catalog number 87CH2473-7. IEEE Computer Society order number 802.

Umaretiya:1992:IES

[UJ92] J. R. Umaretiya and S. P.

Joshi. Insight into the expertseisd: a knowledge based system for structural design. Engineering with Computers, 8(3): 151–161, Summer 1992. CO-DEN ENGCE7. ISSN 0177-0667.

USENIX:1988:UPC

[USE88] USENIX Association, editor. USENIX proceedings: C++ Conference, Denver, CO, October 17–21, 1988. USENIX, San Francisco, CA, USA, 1988. ISBN ???? LCCN ????

Vantaggiato:1998:AEM

[Van98]  Antonio Vantaggiato.  Au-

tomatic exams management with the Common Lisp HTTP Server. Computer Networks and ISDN Systems, 30(1–7):697–698, April 1, 1998. CODEN CNISE9. ISSN 0169-7552 (print), 1879-2324 (electronic). URL

                                        http://www.elsevier.com/
                                        cas/tree/store/comnet/sub/
                                        1998/30/1-7/1847.pdf.

Vashisht:1997:CPI

[Vas97] Rohini Vashisht. Is CLIM 2.0 a platform independent window manager library? Thesis (m.s.), University of Kansas, Electrical Engineering and Computer Science, Lawrence, KS, USA, 1997. 83 pp.

Vaucher:1985:VMC

[Vau85] Jean G. Vaucher. Views of modelling: Comparing the simulation and AI approaches. In Birtwistle [Bir85], pages 3–7. ISBN ???? LCCN Q334 .A47 1985.

Valente:1995:OAM

[VB95] Andre Valente and Joost Breuker. On-line: an architecture for modelling legal information. In ACM [ACM95], pages 307–315. ISBN 0-89791-758-8. LCCN ???? URL

http://www.acm.org/pubs/
contents/proceedings/ai/
222092/index.html.

Verna:2019:QCL

[Ver19] Didier Verna. Quickref: Common Lisp reference documentation as a stress test for Texinfo. TUGboat, 40(2):119–125, ???? 2019. CODEN ???? ISSN 0896-3207. URL https:

//tug.org/TUGboat/tb40-2/
tb125verna-quickref.pdf.

Vogt:1998:FPP

[Vog98] Christopher J. Vogt. Floating point performance of Common Lisp. ACM SIGPLAN Notices, 33(9):103–107, September 1998. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Voss:1994:PC

[Vos94] Greg Voss. Polymorphic C.

Dr. Dobb’s Journal of Software Tools, 19(8):36–42, August 1994. CODEN DDJTEQ. ISSN 1044-789X.

Vick:1990:TTD

[VS90] Shon Vick and Jeff Sponsler. Tools and techniques for the development of Common Lisp applications across heterogeneous machines. In Ng et al. [N^{+}90], pages 76– 85. ISBN 0-8186-9027-5. LCCN TS176.I645 1990. IEEE catalog number 90TH0309-5.

VanderHart:2010:PC

[VS10] Luke VanderHart and Stuart Sierra. Practical Clojure. The expert’s voice in open source. Apress, Berkeley, CA, USA, 2010. ISBN 1-4302-7231-7, 1-4302-7230-9 (e-book). xvi + 210 pp. LCCN ????

Watanabe:1990:ILS

[WAF90] Satoshi Watanabe, Takashi Arano, and Tsuneo Furuyama. Implementation of a LISP system. NTT Review, 2(3):97–102, May 1990. CODEN NTTREK. ISSN 0915-2334.

Wang:1986:ELM

[Wan86]  Lun Wang.  Experimental

Lisp machine. In Gilmore [Gil86], pages 122–124. CO-DEN PSISDG. ISBN 0-89252-692-0. ISSN 0277-786X (print), 1996-756X (electronic). LCCN TS510 .P632 v. 657.

Wang:1997:TPD

[Wan97] Paul S. Wang. Tools for parallel/distributed mathematical computation. In ACM [ACM97], pages 188–195. ISBN ???? LCCN ????

Warn:1986:LVA

[War86] Keith Warn. Lisp vs. Ada: Implications in diagnostics oriented expert systems. In IEEE [IEE86a], pages 409–415. CO-DEN AUPRDX. ISBN ???? ISSN 0734-7510. LCCN TK 7895 A8 A98 1986.

Waters:1987:EIS

[Wat87] R. C. Waters. Efficient interpretation of synchronizable series expressions. ACM SIG-PLAN Notices, 22(7):74–85, July 1987. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
proceedings/plan/29650/p74-
waters/.

Waters:1989:XCL

[Wat89] Richard C. Waters. XP, a Common Lisp pretty printing system. A.I. memo 1102, Massachusetts Institute of Technology, Artificial Intelligence Laboratory, Cambridge, MA, USA, 1989. 38 (or 40??) pp.

Watt:1990:EPM

[Wat90] S. Watt. Experiences programming the Macintosh Toolbox with Common Lisp and the CLOS. In Steels [Ste90b], pages 275–284. ISBN ???? LCCN ????

Waters:1991:ATS

[Wat91a] Richard C. Waters. Automatic transformation of series expressions into loops. ACM Transactions on Programming Languages and Systems, 13(1):52– 98, January 1991. CODEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

URL http://www.acm.org/
pubs/toc/Abstracts/0164-
0925/102806.html.

Watson:1991:CLM

[Wat91b] Mark Watson. Common LISP modules: artificial intelligence in the era of neural networks and chaos theory. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1991. ISBN 0-387-97614-0, 3-540-97614-0. 207 pp. LCCN QA76.73.C28W38 1991.

Watt:1991:IPI

[Wat91c] Stephen M. Watt, editor. IS-SAC ’91: proceedings of the 1991 International Symposium on Symbolic and Algebraic Computation, July 15–17, 1991, Bonn, Germany. ACM Press, New York, NY 10036, USA, 1991. ISBN 0-89791-437-6. LCCN QA 76.95 I59 1991.

Webber:1995:OFP

[Web95]  Adam Webber.  Optimiza-

tion of functional programs by grammar thinning. ACM Transactions on Programming Languages and Systems, 17(2): 293–330, March 1995. CO-DEN ATPSDT. ISSN 0164-0925 (print), 1558-4593 (electronic).

Wertz:1989:CLI

[Wer89] H. (Harald) Wertz. (common) LISP: une introduction a la programmation. Manuels informatiques Masson. Masson Editeur, Masson, France, deuxième edition, 1989. ISBN 2-225-81751-0. 331 pp. LCCN ????

Wholey:1984:DIS

[WF84] Skef Wholey and Scott E.

Fahlman. Design of an instruction set for Common Lisp. In ACM [ACM84], pages 150–158. ISBN 0-89791-142-3. LCCN QA76.73.L23 A26 1984.

Williams:1996:DKB

[WGS96] M. Douglas Williams, Eddy C.

Gorman, and Sajjan G. Shiva. Design of a knowledge-based simulation environment. Simulation, 67(2):121–135, August 1996. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

White:1991:BRK

[Whi91] Jon L. White. Book review:

A “Kinder, Gentler LISP?” COMMON LISP: A Gentle Introduction to Symbolic Computation, by David S. Touretzky. Benjamin/Cummings Publishing Company 1990. Lisp and Symbolic Computation, 4(2): 165–177, April 1991. CO-DEN LSCOEX. ISSN 0892-4635 (print), 1573-0557 (electronic).

Wilensky:1986:CL

[Wil86] Robert Wilensky. Common LISPcraft. W. W. Norton & Co., New York, NY, USA, 1986. ISBN 0-393-95544-3. xvii + 500 pp. LCCN QA76.73.C28 W55 1986.

Wong:1995:CGP

[WL95] Man Leung Wong and Kwong Sak Leung. Combining genetic programming and inductive logic programming using logic grammars. In IEEE [IEE95a], pages 733–736. ISBN 0-7803-2759-4 (softbound), 0-7803-2760-8 (microfiche). LCCN QA76.87.I342 1995. Two volumes. IEEE catalog number: 95TH8099.

Welling:1990:PLF

[WNA90] Joel Welling, Chris Nuuja, and Phil Andrews. P3D: a Lispbased format for representing general 3D models. In IEEE [IEE90e], pages 766–774. ISBN 0-8186-2056-0. LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Winston:1984:GCL

[WS84] Patrick Henry Winston and Guy L. Steele, Jr. Golden Common LISP. Gold Hill Computers, Cambridge, MA, USA, version 1.00. edition, 1984. ISBN 0-917589-05-X. LCCN ???? 5 computer disks plus manual.

Winston:1985:GCL

[WS85] Patrick Henry Winston and Guy L. Steele, Jr. Golden COMMON LISP. Gold Hill Computers, Cambridge, MA, USA, 1985. ISBN 0-917589-02-5. various pp. LCCN ???? Three volumes.

Wholey:1987:CML

[WS87] Skef Wholey and Guy L. Steele, Jr. Connection Machine Lisp: a dialect of Common Lisp for data parallel programming. Technical report TMC-149, Thinking Machines Corp., Cambridge, MA, USA, April 1987. 23 pp.

Wesenberg:1991:LCK

[WT91] David P. Wesenberg and C. Mark Turner. A low cost knowledge based software assistant. In IEEE [IEE91c], pages 640–646. CODEN NASEA9. ISBN 0-7803-0085-8. LCCN TL 693 N37 1991. Three volumes.

Wuorinen:1989:DTP

[Wuo89] J. H. Wuorinen, editor. Digest of technical papers: 1989 IEEE International Solid-State Circuits Conference, volume 32. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 1989. CODEN DT-PCDE. ISBN ???? LCCN TK7870 .I58 1989. IEEE catalog no. 89CH2684-9.

Weerawarana:1992:PCG

[WW92] Sanjiva Weerawarana and Paul S. Wang. Portable code generator for CRAY FOR-TRAN. ACM Transactions on Mathematical Software, 18(3): 241–255, September 1992. CO-DEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Xu:1996:SEU

[XRS96] Chuanbo Xu, Aamir Razzak, and Peter N. Saurugger. Some experience using OODB in the interoperation of heterogeneous genome database systems. American Society of Mechanical Engineers, Petroleum Division (Publication) PD, 74 (2):53–61, 1996. CODEN ASMPEX.

Xu:1992:SES

[XXL92] Lu Xu, Yonghui Xiao, and Deqian Li. Studies on expert system for extraction of rare earths. Journal of the Chinese Rare Earth Society (English Edition), 10(4):247–252, December 1992. CODEN JCE-SEN. ISSN 1001-6503.

Yan:1991:RFA

[Yan91] Tak W. Yan. A rational function arithmetic and simplification system in Common Lisp. SIGSAM Bulletin (ACM Special Interest Group on Symbolic and Algebraic Manipulation), 25(4):4–6, October 1991. CODEN SIGSBZ. ISSN 0163-5824 (print), 1557-9492 (electronic).

Yuasa:1987:ICL

[YH87] T. (Taiichi) Yuasa and Masami Hagiya. Introduction to COM-MON LISP. Academic Press, New York, USA, 1987. ISBN 0-12-774860-1. ix + 293 pp. LCCN QA76.73.C28Y83 1987. Translated from the original Japanese edition by Richard Weyhrauch and Yasuko Kitajima.

Yen:1991:UPI

[YJM91] John Yen, Hsiao-Lei Juang, and Robert MacGregor. Using polymorphism to improve expert system maintainability. IEEE expert: intelligent systems and their applications, 6 (2):48–55, April 1991. CODEN IEEXE7. ISSN 0885-9000.

Young:1992:TMO

[You92] Norman Young. Two models of object-oriented programming and the Common Lisp Object System. ACM SIG-PLAN Notices, 27(4):27–36, April 1992. CODEN SINODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Young:1997:CVS

[You97] William D. Young. Comparing verification systems: Interactive consistency in ACL2. IEEE Transactions on Software Engineering, 23(4):214– 223, April 1997. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic).

Yuasa:1988:CLD

[Yua88] T. (Taiichi) Yuasa. Common Lisp drill. Academic Press, New York, USA, 1988. ISBN 0-12-774861-X (paperback). x + 232 pp. LCCN QA76.73.C28Y82 1988. Originally published in Japanese by Iwanami Shoten, Publishers, Tokyo, 1987.

Yuasa:1990:DIK

[Yua90] Taiichi Yuasa. Design and implementation of Kyoto Common Lisp. Journal of Information Processing, 13(3):284–295, 1990. CODEN JIPRDE. ISSN 0387-6101.

Yuasa:1992:MMG

[Yua92] T. Yuasa. Memory management and garbage collection of an extended Common Lisp system for massively parallel SIMD architecture. Lecture Notes in Computer Science, ?? (637):490–506, ???? 1992. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Yasumoto:1994:JCS

[YY94] Taichi Yasumoto and Taiichi Yuasa. Japanese character set handling in Kyoto Common Lisp. Advances in software science and technology, 5(??):165– 182, ???? 1994. ISSN 1044-7997.

Yasumura:1990:HCL

[YYK^{+}90] Michiaki Yasumura, Katsuhiko Yuura, Masaaki Kurosu, Toshihisa Aoshima, Nobuyuki Takeichi, and Yoshimitsu Oshima. HiLISP — a Common Lisp with an optimizing compiler on a mainframe. Journal of Information Processing, 13(3):296– 303, ???? 1990. CODEN JIPRDE. ISSN 0387-6101.

Yuura:1990:HHC

[YYN90] K. Yuura, M. Yasumura, and M. Nagata. HiLISP: Highperformance Common Lisp and its optimizing tool. In Steels [Ste90b], pages 91–100. ISBN ???? LCCN ????

Yang:1995:FSE

[YYS95] Chun-Ling Yang, Akihiko Yokoyama, and Yasuji Sekine. Fault section estimation of power system using colored and timed Petri nets. Electrical Engineering in Japan (English translation of Denki Gakkai Ronbunshi), 115(2):89– 101, April 1995. CODEN EEN-JAU. ISSN 0424-7760.

Zhang:1991:HFH

[ZB91] C. Zhang and D. A. Bell.

HECODES: a framework for HEterogeneous COoperative Distributed Expert Systems. Data and Knowledge Engineering, 6(3):251–273, May 1991. CODEN DKENEW. ISSN 0169-023X (print), 1872-6933 (electronic).

Zorn:1989:LEM

[ZHL^{+}89a] Benjamin Zorn, Kinson Ho, James Larus, Luigi Semenzato, and Paul Hilfinger. Lisp extensions for multiprocessing. In Hoevel et al. [HMS^{+}89], pages 761–770. CODEN PHISD7. ISBN 0-8186-1911-2 (paperback: vol. 1), 0-8186-5911-4 (microfiche: vol. 1), 0-8186-9911-6 (casebound: vol. 1), 0-8186-1912-0 (paperback: vol.

2), 0-8186-5912-2 (microfiche: vol. 2), 0-8186-9912-4 (casebound: vol. 2), 0-8186-1913-9 (paperback: vol. 3), 0-8186-1914-7 (paperback: vol. 4), 0-8186-5914-9 (microfiche: vol. 4), 0-8186-9914-0 (casebound: vol. 4). ISSN 0073-1129. LCCN ???? Computer Society order number 1911. IEEE catalog number 89TH0242-8.

Zorn:1989:MES

[ZHL^{+}89b] Benjamin Zorn, Kinson Ho, James Larus, Luigi Semenzato, and Paul Hilfinger. Multiprocessing extensions in Spur Lisp. IEEE Software, 6(4): 41–49, July 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Zorn:1990:CMS

[Zor90] Benjamin Zorn. Comparing mark-and-sweep and stop-and-copy garbage collection. In ACM [ACM90b], pages 87–98. ISBN 0-89791-368-X. LCCN QA 76.73 L23 A24 1990. ACM order no. 552900.

Zimmermann:1993:OFP

[ZS93] H.-J. Zimmermann and H.-J. Sebastian. Optimization and fuzziness in problems of design and configuration. In IEEE [IEE93d], pages 1237– 1240. ISBN 0-7803-0616-3 (microfiche), 0-7803-0615-5 (casebound), 0-7803-0614-7 (softbound). LCCN TJ212.2.I3249 1993. Two volumes. IEEE catalog number: 93CH3136-9.